A Complete Bibliography of Publications in the Journal of Computer Security

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: =beebe@math.utah.edu=, =beebe@acm.org=, =beebe@computer.org= (Internet) WWW URL: =https://www.math.utah.edu/ beebe/=

23 September 2019 Version 1.05

Title word cross-reference

2 [QJ97]. k [CYC18, LWZZ15]. Ω [RFLW96]. π [SV03].

-jump [LWZZ15]. -Skyband [CYC18].

5PM [BEM^{+}13].

9798 [BCM13].

ABAC [BAVS19]. Abadi [MW04]. Abelian [MS05a, MS05b]. abstract [BPT19, GM10, GTZ04, MMHS13, PG09]. Abstractions [NSC18, CDF^{+}08]. Accelerated [Elb08]. access [AJS12, AHB08, ASL96, ACDS08, ACK^{+}10, AR12, BBFR00, BDJP10, BS02, BGS06, Bro02, DFP^{+}13, DFP^{+}18, HSR08, HTS00, HPS03, KNS08, LKAJ16, LCL^{+}15, MMRS19, RW97, SAsC11, TR11, TL07, UAV^{+}14, VAGL09, YGSY15, YSD13, ZRG08]. Accessing [DLL^{+}18]. Accountable [BBS^{+}15, BLL02]. Achieving [AJB93, HH09]. Action [Sin97]. active [BPS08, KGA03]. activity [TC16]. Ad [ZSXJ06, AJS12, RCBM07, RH07, ZDM07]. Ad-Hoc [ZSXJ06, AJS12]. adaptation [Bis12]. adaptive [KB11]. Address [XC09, PSJ^{+}13]. Address-space [XC09]. Adjoining [GM10]. administration [SB99, YGSY15]. administrative [AR12, RTV18, VAGL09, YGSY15]. administrators [AWJS19]. Advances [Vis13]. adversarial [CJP^{+}18]. AES [Elb08]. against [BCFK15, HSR08, LHM^{+}10, MCL^{+}19, PV05, RCBM07, SPS^{+}14]. agenda [Pri06]. agent [DB11]. agents [BR04]. Aggregating [HZO^{+}13]. Aggregation [Fol92, CPPS07]. agreement [PQ03, PQ06]. ahead [SM05]. aid [BS04]. aided [NTU11]. algebra [RS01]. algebraic [ABG04, CDL06]. Algebras [FG95, BCLM05]. Algorithm [AJJ95, HDC95]. algorithms [AGHP14, LWZZ15, LMG^{+}18, MM96, SAE^{+}05]. Aligning [BBK14]. Allocation [Mil93b]. among [MTJ^{+}19, ZHZY18]. Amongst [CGM95]. analyses [Pau01]. Analysing [KL11, Low04a]. Analysis [HDC95, Mea92, NS06, Syv92, Wra92, ABG04, AS15, AR12, AH00b, BBDLM14, BBI15, BCLM05, BDGS16, BR05b, BHČ04, CFL13, CDL^{+}05, CYC17, CHM07, CS05, CEHM07, CS13, CW17, CDWM19, DKS10, DMV15, ES06, GBL18, GEL98, GTZ04, JKK10, KKN^{+}17, KM10, KK16, LOS16, LT17, LŁA15, Low98, Mea01, MSC04, MS05a, MS05b, MVB10, PV05, RS02, RTV18, RTWH11, Shm04, SBP01, SKEG14, TR11, UC07, UAV^{+}14, VVB^{+}09, VIS96, WYSJ08, War05, YGSY15, ZZW^{+}11, BDP05]. Analytic [MM96]. analyzed [TvdRO03]. Analyzer [MSC04]. Analyzing [CAFL10, CCP^{+}17, UTHP19, Weh07, ABK^{+}11]. Android [CLD^{+}17, JAH^{+}16, MMF15]. announced [CGLZ03]. Anomaly [SAE^{+}05, LWWJ02, VVB^{+}09]. anomaly-based [VVB^{+}09]. Anonymity [BMS95, HO05, CMS97, HS04, LS02, PSJ^{+}13, Shm04]. anonymity-revoking [CMS97]. anonymization [DRRW11, SVA11]. Anonymizing [ACM15, PS14]. anonymous [SBS05]. answering [CYC18]. answers [BH09]. Antigone [MP06]. APIs [HBS16]. applets [BCG^{+}02]. application [BGE^{+}18, BHČ04, NZS05, QJ97, VC05]. Applications [GG92, BHSV14, BLL02, CZ16, DB11, DT97, GLS^{+}17, GEL98, JKK10, KTV12, SR19]. applied [BS18, DKR10]. Applying [BFGS08, BFGS09, Mea92]. Approach [BY95, WL93, ABG04, ALP11, BBFR00, BFM05, BFO05, CLD^{+}17, CDF^{+}12, FF11, HS04, KSS13, MS03, Pau98, SBP01]. approaches [BDG^{+}10]. Approximate [DHW04, NG08]. APTs [RRAZ19]. Architecture [AXR12, TS93, ABFK03, DPV09, HTS00, MM96, dACD^{+}16]. architectures [LŁA15]. area [AFHS09]. argument [LZ13]. ARM [GNDB16]. aspects [CMPP14]. assessment [BT06, Mal10, SSBW12]. assignment [SB99]. assignments [AWJS19, VAGL09]. associating [SR19]. association [VC05]. associations [DFJ^{+}15]. Assurance [BMBS10, EMO^{+}93, CT09]. Assuring [HKOS18]. asymmetric [GJ04]. Asynchronous [Yah93]. Athena [SBP01]. ATM [Chu96]. attack [EVK02, GSAV18, HZO^{+}13, Kre13, NG08, OMSH04, SR19, WYSJ08]. attackers [RCBM07]. Attacking [CS13]. attacks [BPS08, BBDLM14, BWA10, BR04, BCFK15, CS05, DPV09, DS99, DLL^{+}18, Han11, HLS03, HNS^{+}14, KB11, LHM^{+}10, Low04a, Möl05, OMSH04, PQ03, PV05, SPS^{+}14, SSM18, SVSM07, VVB^{+}09, YTLD19, ZGDS13]. attempt [Mat98]. attribute [LCL^{+}15, PTMW10, TLY^{+}19]. attribute-based [LCL^{+}15, PTMW10, TLY^{+}19]. attribution [TC16]. auction [DDL15]. audio [SKEG14]. audit [DS99]. Auditable [FM98]. auditing [MTJ^{+}19]. augmented [SZP^{+}12]. Authentic [DGMS03]. authenticated [PQ03, PQ06]. Authentication [CGM95, GLP93, Gut04b, HP00, NR11, QJ97, TA92, ATI^{+}10, BGE^{+}18, BCM13, BCL97, BSSM^{+}07, BFM07, CDWM19, DGK^{+}04, DMV15, GTY08, GTZ04, HSH11, HCM11, JHS96, MvO11, SKEG14]. Authenticity [GJ03]. Author [Ano92a, Ano93, Ano95a, Ano96a, Ano97, Ano98, Ano01, Ano02, Ano03, Ano04, Ano05, Ano06, Ano07a, Ano08, Ano09, Ano10, Ano11, Ano12, Ano13, Ano14, Ano15a, Ano16, Ano17]. authorities [MG08]. Authorization [BOS95, DS97, DFJ^{+}11, Kar00, WL93, AH00b, BBDLM14, BBK14, BFG10, BBFS00, CCF98, GP12, MP06, SWC07]. authorizations [CDF97, DFP^{+}18]. authors [Ano96b]. Automated [BHSV14, BS18, CDF97, KK16, RTV18, AR12, CDE^{+}10, SHM02, ZGD04]. Automatic [Bla09, RTWH11, UC07, SBP01, ZF12]. Automatically [KB11, dSRCP17]. Automating [Gue09]. autoregressive [NG08]. availability [BW08]. aware [ACDS08, BDGS16, HZL^{+}18, RS16]. awareness [MSas13].

ballot [CS13]. bandwidth [DMM04, NR11]. barriers [BS18]. based [AJS12, AHB08, ATI^{+}10, AR12, AH00b, AHB00, BGJ03, BBI15, BBFR00, BY95, BFO05, BGS06, CG06, CMMV07, CFL13, CWT14, CLK04, CCP^{+}17, DHRS11, DR16, DPV09, DGMT19, DLYZ11, DLL^{+}18, Dug04, ES06, EVK02, GMR^{+}11, GGS^{+}09, HTS00, HSH11, KR03, KNS08, KTV12, LS02, LGGO18, LHM^{+}10, LCL^{+}15, LT17, LKWB06, LSMR16, MMHS13, Mea01, MMF15, MVB10, NR11, NTU11, PTMW10, PGK14, PG09, RS02, RAJ98, RVV17, RW97, SB99, SVA11, ST05, SKEG14, TLY^{+}19, TR11, UAV^{+}14, VAGL09, VPZ16, VK99, VVB^{+}09, WWJ04, WLJW07, WD08, WMF^{+}17, YGSY15, YSD13, vOT11, NZS05]. bases [BF99, BBFS00]. Basic [Jac92]. batch [AGHP14]. Bayes [CPP08]. Beacons [JHS96]. been [GLZ11]. Behavior [YSD13, RTWH11, WDDN00]. Behavior-based [YSD13]. Belief [Syv92]. beliefs [CMS09]. Bell [Mil96]. BetterTimes [HKOS18]. between [CDL^{+}05, Pau01]. BGV [GHPS13]. BGV-style [GHPS13]. bi [CGKW17]. bi-objective [CGKW17]. bilinear [KM10]. Biometric [ATI^{+}10]. biometrics [BSSM^{+}07]. bisimulation [DKR10]. Bitcoin [YTLD19]. Black [MSas13]. Black-box [MSas13]. Blind [ABO06, BFPV13]. block [MCL^{+}19]. boots [MS96]. bounded [DLMS04, ŢBEB08]. bounding [ABK^{+}11, BMV15, KKN^{+}17, YT11]. bounds [KB11]. box [MSas13, SV03]. Brandt [DDL15]. breaches [CGLZ03, GLZ11]. Broadcast [BMS95]. browser [ABR13, BCFK15]. browsers [BCL97, GBL18]. browsing [SRG97]. BRSIM [BPS08]. BRSIM/UC [BPS08]. BRSIM/UC-soundness [BPS08]. buffer [ZZW^{+}11]. Building [LAGM17, PQ06]. built [GSAV18]. built-in [GSAV18]. business [BBK14].

C [DGJN14]. CA [BVC^{+}14]. cache [MCL^{+}19]. calculus [AACP13, BS18, BGS06, DKR10, HM10, MS96]. Calibrating [HM10]. Call [Ano92b, Ano95b]. calls [HFS98]. Can [KNTU13, BBS^{+}15]. capability [MG08]. capable [TEML08]. CAPTCHA [SKEG14]. CAPTCHA-based [SKEG14]. CAPTCHAs [LUV^{+}18, MGS^{+}17]. capture [SST08]. card [Bel03, BCG^{+}02]. Cardinality [WWJ04, VC05]. Cardinality-based [WWJ04]. cards [SRS^{+}02]. Cascade [HCH^{+}93, BFO05, DPV09]. case [GP12]. cases [ABHS09]. Casper [Low98]. catalogue [SR19]. causal [dC96]. causality [SV03]. causes [CRB19]. CCA1 [MSas13]. cellular [HCM11, TEML08]. centers [CDE^{+}10]. centric [RW97]. centricity [BSCGS07]. Certificate [CEE^{+}01, HTML09, BLL02, CAB10, LHM^{+}10, WLM01, WLM02]. certificate-based [LHM^{+}10]. Certificates [TA92, GBL18, LKWB06]. certified [NZS05]. chain [CEE^{+}01, LWM03]. challenge [TC16]. change [HJT^{+}96]. channel [GSAV18, KSWH00, KB11, SCB^{+}19]. Channels [BMS95, Hu92, Wra92, DMV15, GT17, MCL^{+}19, NR11]. Characterizing [PM18]. check [ST15, ŢBEB08]. checked [MCB13]. checkers [RB99]. Checking [BCG^{+}02, BDP05, BJLT01, CC17, DHRS11, DR16, JR04, Low99, ZRG08]. checks [DFJ^{+}16]. checksum [BGJ03]. checksum-based [BGJ03]. Chinese [ACM04]. ciphers [KSWH00, MCL^{+}19]. ciphertexts [NMP^{+}13]. circuit [CMTB16]. Classification [FG95, BFGS08, BFGS09]. classified [GM10]. click [vOT11]. click-based [vOT11]. climate [ABM^{+}19]. Cliques [PQ06]. Cliques-type [PQ06]. cloud [ABR13, CWT14, DFJ^{+}16, HZL^{+}18, LT17, ST15, ZHZY18, ZGDS13]. cloud-based [CWT14]. cloud1 [YY15]. clouds [AWJS19, MTJ^{+}19]. code [PG09, SV03, WS02, XC09]. codes [ATI^{+}10, SVSM07]. coercion [KTV12]. cognitive [MGS^{+}17]. collaboration [AJS12]. Collaborative [FJ95, SZP^{+}12]. collection [BF99]. collective [CCBE13]. colluding [RCBM07]. Collusion [YTLD19]. combined [VVB^{+}09]. combiners [Her09]. Commerce [ABO06, BKA^{+}97]. Communication [Vis13, JAH^{+}16, LSMR16, SBS05, Xu07]. communications [DMM10]. community [CYC17]. comparative [IBS03, NR11, SAE^{+}05]. comparing [TL07]. comparison [CDL^{+}05, MM96]. compiled [WS02]. compiler [CDF^{+}08, Low98, LSMR16, ZZW^{+}11]. complete [ŢBEB08]. Completeness [MW04, ABHS09, Low99]. complex [TC16]. complexity [DLMS04, KKN^{+}17, OMSH04]. Composability [ZL95, BU16]. composition [BDF09, CDM12, NF18, SV03]. Compositional [BPR07, DDMP05, DMP03]. compression [Weh07]. computation [BA12]. Computational [BU10, BWA10, KKN^{+}17, RDDM10, War05]. Computationally [KM10, CB15, LZ13]. Computer [Ano92b, BR09, FJ95, LBF^{+}93, BS04, HSR08, Mat98, SVSM07]. computers [MvO11]. computing [ABR13, SV15, ZGDS13]. concept [RRAZ19]. concrete [BBDLM14]. Concurrency [AJJ95, DFP^{+}13]. concurrent [BN07, BPR07, VS99]. conditional [Hal17]. confidentiality [CDF^{+}12, DFP^{+}18, NSH14]. configuration [FF11, RS02]. configurations [UC07]. Configuring [AFHS09]. confinement [LJM00]. connections [SCB^{+}19]. Conspiracy [Bis96]. constant [BPT19, YY15]. constant-cost [YY15]. constant-time [BPT19]. constrained [DMV15]. Constraint [AHB08, MVB10, BFO05]. constraint-based [BFO05, MVB10]. constraints [CCF98, CDF^{+}12]. construction [HLVA11, LAGM17, MCB13, MSas13]. Constructions [LHM^{+}10, CGKO11]. constructs [Mal10]. consumption [LMG^{+}18, MMF15]. contactless [Sin11]. Content [DTG16]. Context [BDGS16, BGE^{+}18, RS05]. Context-aware [BDGS16]. context-explicit [RS05]. contexts [BMPR05, FR06]. contingency [HLVA11]. contract [NS06]. Control [AJJ95, AHB08, ASL96, ACDS08, ACK^{+}10, AR12, BBFR00, BJLT01, BDJP10, Bis12, BGS06, BvDS17, Bro02, HTS00, HPS03, HH09, KNS08, LGGO18, LCL^{+}15, MMRS19, RW97, SPS^{+}14, SAsC11, TR11, TL07, UAV^{+}14, VAGL09, WWJ04, WLJW07, YGSY15, YSD13, ZRG08, BDP05]. controlled [BW08]. controller [ZV19]. Controlling [dC96]. CookiExt [BCFK15]. cooperative [BFM05, LKAJ16]. Coordinated [OMSH04, ZGDS13]. Coprocessor [WD08]. Coprocessor-based [WD08]. CORBA [Bro02, Kar00]. core [AACP13]. correct [DJLP10, FHG99]. Correctability [Mil95b]. correcting [ATI^{+}10]. Correctness [AJB93, McL92]. correspondences [Bla09]. corruption [BGJ03]. Cost [IBS04, ALP11, CGLZ03, LKAJ16, LFM^{+}02, MBK^{+}15, Mea01, SSBW12, YY15]. cost-based [Mea01]. cost-sensitive [LFM^{+}02, SSBW12]. costs [GLZ11]. counterevidence [BLL02]. coupon [AS15]. coverage [JBH13]. Covert [Wra92]. CPU [LMG^{+}18]. creation [ASL96]. credential [LWM03]. credentials [WCJS97]. crime [HSR08]. cryptanalysis [KSWH00]. Crypto [SSM18]. Cryptographic [CFG^{+}17, Syv92, BR05b, CFL13, Coh03, CDL06, Dug04, DGJN14, GJ04, GEL98, KAM08, Pau98, YSM14]. cryptography [ACK^{+}10, DGMT19, Lop06, Mat98]. cubes [WWJ04]. curves [BGH^{+}13]. CWASAR [BKA^{+}97]. Cyberspace [GOvdR99]. cycle [LKWB06]. cycles [ABHS09, BPS08].

DABKE [TLY^{+}19]. Data [NBM95, Yah93, AXR12, ABM^{+}19, BDF^{+}12, BT06, BLB^{+}09, CDE^{+}10, CC17, CDF^{+}11, CDF^{+}12, Cli00, DS99, DFP^{+}13, DFJ^{+}15, DGMS03, DRRW11, DRD11, GLS^{+}17, HLVA11, LWZZ15, MRR^{+}18, RB99, SPS^{+}14, ST15, SM05, VPZ16, WWJ04, ZMHT07]. Data-Exchange [Yah93]. Database [AJJ95, AJB93, HDC95, CDF97, DS97, DCMP16, LKAJ16, MM96, VPZ16, YWW^{+}09]. Databases [BOS95, KK95, Mot92, TS93, CT09, SC00, TvdRO03, WYSJ08]. datasets [ACM15]. DDoS [PSJ^{+}13, SM05]. decentralized [ACM04, BFG10]. Decidability [RS05]. Deciphering [GSAV18]. decision [HS05]. Declassification [MSZ06, SS09, MB09, RS16]. declassification-aware [RS16]. decoys [BKP^{+}12]. decremental [ZV19]. Defeating [SSM18]. defense [CJP^{+}18, PSJ^{+}13]. Defining [Low04b]. definition [KTV12]. definitions [CGKO11, HMQU09]. definitive [DT97]. Delay [MCL^{+}19]. delegatable [MG08]. Delegated [ST15]. deletion [BF99]. delivery [NZS05, ZMHT07]. deniable [TLY^{+}19]. Denial [Mil93b, Mea01, Möl05]. Dependencies [BC92, MTJ^{+}19, SST08, dC96]. dependent [BPS08]. Depender [WLM01, WLM02]. deploy [dSRCP17]. Deploying [BAVS19]. deposits [YTLD19]. derivation [CDF97, DDMP05]. deriving [KB11]. Design [BFG10, LŁA15, ABFK03, Gut04b, KSS13, LWZZ15, RW97, SR19, ZAF08]. Designing [WMF^{+}17, LUV^{+}18]. detect [BR04, DS99]. Detecting [BBDG10, CYC17, HDC95, RCBM07, SG96, JKK10, WDDN00]. Detection [Mil99, BGJ03, BHSV14, DPV09, EVK02, HFS98, IBS03, IBS04, KPS16, KSZ02, LFM^{+}02, LWWJ02, MMF15, NG08, SM05, SVSM07, SG02, SHM02, SCB^{+}19, UC07, VK99, VVB^{+}09, XVW^{+}06, Yas02, SAE^{+}05]. detectors [KSZ02]. determinism [RWW96]. deterministic [BBG^{+}18]. develop [Lot97]. devices [BDJP10, CMTB16, CFL13, MvO11]. Diet [MGK^{+}17]. Diet-ESP [MGK^{+}17]. differential [CYC18]. differentially [AAC^{+}15]. differentially-private [AAC^{+}15]. differentiation [MVB10]. Diffie [MS05a, MS05b]. Digital [CMS97, AGHP14, BSSB06, WD08, WCJS97]. dimensional [CT09, WLJW07]. Dimensions [SS09]. directories [LKWB06]. disclosure [BH09, LWZZ15, MB09]. Discovering [BBDLM14]. discovery [CYC17, CEE^{+}01, LWM03]. discretionary [AH97]. disjoint [Gut04b]. disjointness [BHM14]. display [KNTU13]. display-equipped [KNTU13]. dissemination [BLB^{+}09]. distance [ABK^{+}11, BMV15, KKN^{+}17]. distance-bounding [BMV15, KKN^{+}17]. Distributed [GLP93, LWM03, WL93, BFM05, BGS06, DFJ^{+}11, HLVA11, HVL12, JHS96, LAGM17, MS03, MS96, SWC07, YSD13, YLZ05, ZDM07]. Distributing [CGM95]. Distribution [BMS95, CCD06, SPD^{+}10, WLM01, WLM02]. distributions [BDF^{+}12]. diversity [PS10b]. DNP3 [CDWM19]. DNSSEC [Gue09]. document [KPS16]. documents [BFM05, DGK^{+}04]. Dolev [BPS08]. Domain [MSC04, CS05, KM98, SG02]. domain-specific [SG02]. downward [GLZ11]. driven [AJS12, BGT15, PM18]. DRM [DTG16]. Duty [NBM95]. Dynamic [Bis12, BFM07, MMHS13, NBM95, BM99, FR06, MGS^{+}17, PSJ^{+}13, SST08, ZV19]. dynamically [LOS16].

e-mail [NZS05]. e-Passports [LG10]. E-voting [CW17]. eavesdropping [Han11]. economic [CGLZ03]. ecosystems [RRAZ19]. Edge [VAGL09]. Edge-RMP [VAGL09]. editor [BEN96, BMK97, DM00, Fol98, Fol99, Lin99, Lin00, Mer97, SS97, AH00a, Cup02, Foc05, Fri02, Gon95, Gor05, Gri11, Gut04a, JG02, LZ11, Li12, Mil93a, Mil95a, Mil96, San92a, Sch03, Sch04, Syv01, Syv03, TvdR03]. Editorial [Dam08, Foc10, JM10]. Editors [Ano07b, BZ14, BR05a, JM92a, JM92b, JM93, JM95a, JM95b, JM96a, JM96b, JM97, LM92, MK93, YGH08, YRY08]. Effective [NSH14, XVW^{+}06, BFGS08, BFGS09, IBS04]. effects [GJ04, HH09]. Efficient [ASV08, BDG14, DFJ^{+}16, MCB13, ZSXJ06, CGKO11, HVL12, KSS13, LYW^{+}10, LZ13, LŁA15, RTV18, SBP01, Zúq05]. Efficiently [SAsC11]. elastic [ZHZY18]. Electronic [ABO06, BKA^{+}97, DKR09, Sin11]. Eliminating [BLL02]. elliptic [BGH^{+}13]. email [BDG^{+}10]. embedded [KSZ02]. Embedding [BR04]. empirical [CGLZ03]. eMRTD [BB14]. encrypted [DRD11, MW04, MRR^{+}18]. Encryption [NZS05, ABHS09, BPS08, CG06, CGKO11, GHPS13, Gut04b, LCL^{+}15, MSas13, SP03, WMF^{+}17]. energy [MMF15]. energy-based [MMF15]. enforce [CDF^{+}12]. enforcement [ASV08, BBK14, BM12, CDE^{+}10, CFG^{+}17, DS97, DFJ^{+}11, LKAJ16, MBK^{+}15, MMHS13, MT08, ZAF08]. Enforcing [AH97, CCF98, DFP^{+}18, MP06, MSZ06, BBFR00, CDF^{+}11, SAsC11, WS02]. engineering [GEL98]. enhanced [ACK^{+}10, BDP05, DTG16, GHRS05]. Enhancing [LWWJ02]. enterprise [HZO^{+}13]. enterprises [HSR08, MT08]. entity [BCM13]. environment [ASV08, FBB^{+}18, IBS03, Yas02]. equipped [KNTU13]. equivalence [BWA10]. equivalences [BS18]. Eradicating [BMP18]. error [ATI^{+}10]. errors [VVB^{+}09]. ESP [MGK^{+}17]. EsPRESSO [BDG14]. Establishing [BSSB06, Gut14]. establishment [GT17, ZMHT07]. estimating [MMF15]. EU [CLM^{+}10, Cam10]. EU-funded [CLM^{+}10, Cam10]. European [Ano92b, BKA^{+}97]. evaluable [CZ16]. Evaluating [PS14, YWW^{+}09]. evaluation [BW08, BDG14, CMTB16, CMMV07, DFJ^{+}11, IBS03, KSS13, SAE^{+}05]. Event [BGT15]. Event-driven [BGT15]. events [JPSS16, MTJ^{+}19]. evidence [CGLZ03]. exact [GEL98]. Exchange [Yah93, BHČ04, KR03, TLY^{+}19]. execution [AHB00, DDNP14, RS16, dSRCP17]. existence [HS05]. Expected [DMV15]. Experimental [PV05, IBS03]. experts [BFGS08, BFGS09]. explicit [RS05]. Exploiting [ACK^{+}10, TEML08, vOT11]. Exploring [DRRW11, SA16]. exponentiation [MS05a, MS05b]. exposure [Cli00]. expressions [MW04]. Expressive [San92b, ASL96, TL07]. Extended [AS92, BOS95, BCG^{+}02]. extension [LYW^{+}10]. extensions [DKS10, Elb08].

Fabric [LAGM17]. factor [BSSM^{+}07]. fair [KR03, YTLD19]. faithfulness [GTZ04]. Fast [HDC95, ZF12, YTLD19]. fast-payment [YTLD19]. Fault [HL01, WLM01, WLM02]. Fault-preserving [HL01]. fault-tolerant [WLM01, WLM02]. feasible [CLD^{+}17]. feature [BFGS08, BFGS09]. features [SCB^{+}19]. federated [BMBS10, DS97, GTY08]. Federation [SS10, BSSB06]. federations [CDF97]. Field [GHPS13]. filtering [BDG^{+}10]. Finding [PS10a, dSRCP17]. Fine [RS16]. Fine-grained [RS16]. Finite [GL10]. firewall [NF18, UC07]. First [Coh03, Hal17]. First-order [Coh03, Hal17]. Fixed [WDDN00]. Fixed- [WDDN00]. fixing [CS13]. flaw [HLS03]. flaws [BBDG10]. Flexible [DGK^{+}04]. Flow [BDP05, Gra92, Kre13, AAP12, BN07, Bec12, BJLT01, Bis96, BMPR05, BPR07, BvDS17, CHM07, CMS09, CFG^{+}17, DHRS11, DR16, FR06, GHRS05, HH09, HBS16, LOS16, Low04b, MSB17, SST08, VIS96, YT11, dACD^{+}16]. Folklore [Her09]. Forcing [HM13]. forensic [BS04]. Foreword [Cam10, LaP96]. Formal [CCBE13, DKS10, Mea92, MSC04, NRW14, ABHS09, AS15, BBDLM14, BHČ04, CW17, CDWM19, GL10, Pau01, SRS^{+}02]. formalization [TR11]. formally [Lot97]. Forward [Mil95b]. forwarded [BGSW11]. Foundation [Gra92]. Foundational [CMPP14]. frame [LSMR16]. framework [AHB08, AS15, ABK^{+}11, BS02, DLYZ11, JAH^{+}16, KNS08, LVA14, Mea01, PS10b, TLY^{+}19, ZMHT07]. frameworks [IBS04]. free [BBK14, DTX09]. frequency [CS05, Han11]. freshness [ŢBEB08]. fully [DDL15]. function [HS05, KSS13, QJ97]. Functional [McL92]. functionality [TEML08]. functions [CZ16]. funded [CLM^{+}10, Cam10]. fusion [BT06, SM05]. future [HP00]. Fuzzy [Hu92, Tro93].

game [KR03, KTV12, LT17, MGS^{+}17, SKEG14]. game-based [KR03, KTV12]. game-theoretic [SKEG14]. garbage [BF99]. garbled [CMTB16]. general [DGJN14, KSS13]. general-purpose [DGJN14]. generalized [Elb08, MSAV15]. generated [AGHP14]. generating [BKP^{+}12]. generation [CB15, ZF12]. generator [Zúq05]. Generic [Wan06]. gigabit [IBS03]. GKMPAN [ZSXJ06]. Global [DLRS01, TA92, CDF97, KK16]. Globally [LG10]. goals [GHRS05, Gut14]. GPU [LŁA15]. grained [RS16]. Grant [Bis95, Bis96]. granularities [LWWJ02]. Graph [FF11]. graphical [vOT11]. graphs [BBI15, BJLT01, CAFL10, HZO^{+}13, WYSJ08, WLM01, WLM02]. grid [CDWM19]. Group [MMRS19, ZSXJ06, CLK04, CCBE13, DTX09, MS05a, MS05b, PQ03, PQ06, Xu07, MSC04]. groups [CYC17]. Guessing [BWA10, Low04a]. Guest [BEN96, BMK97, Dam08, DM00, Fol98, Fol99, Lin99, Lin00, Mer97, SS97, Ano07b, AH00a, BZ14, BR05a, Cup02, Foc05, Fri02, Gon95, Gor05, Gri11, Gut04a, JG02, LZ11, Li12, LM92, MK93, San92a, Sch03, Sch04, Syv01, Syv03, TvdR03, YGH08, YRY08]. Guest-editor [BEN96]. Guiding [DGJN14].

habits [LGGO18]. Hails [GLS^{+}17]. hardening [BWJS18]. HardIDX [FBB^{+}18]. hash [QJ97]. hashing [BGH^{+}13, LŁA15]. haystack [PS10a]. HB [HSH11]. healthcare [YSD13]. hear [BBS^{+}15]. Helios [CS13]. Hellman [MS05a, MS05b]. Hermes [ZMHT07]. heterogeneous [BSS97, BWJS18]. hiding [CPP08, HO05, HS04]. hierarchical [BFGS08, BFGS09, KAM08, ST15, SB18b, WD08]. Hierarchies [DMM10, RTV18]. hierarchy [BBI15, CLK04]. High [EMO^{+}93, DMM04, Han11, MR97, SV15, Zúq05]. high-frequency [Han11]. high-throughput [MR97]. hijacking [BCFK15, CRB19]. history [KNS08]. history-based [KNS08]. Hoc [ZSXJ06, AJS12, RCBM07, RH07, ZDM07]. homomorphic [GHPS13, WMF^{+}17]. honeypots [CJP^{+}18]. Hordes [LS02]. host [DS99]. Human [Mat98]. humans [SSM18]. Hybrid [CJP^{+}18]. Hyperproperties [CS10].

ICT [CLM^{+}10, Cam10]. ID [GTY08]. Identification [DS99]. Identifying [LMG^{+}18]. Identity [PSJ^{+}13, SS10, BMBS10, BSSB06, CG06, MT08, WD08]. identity-based [CG06]. IDS [BT06]. IEC [BCM13]. IEEE [RCBM07]. II [LB96]. illicit [SG96]. IME [DLL^{+}18]. imitation [SSM18]. impact [BGE^{+}18, GLZ11]. imperative [CHM07]. Implementation [CS05, AXR12, SB99]. implementations [BHM14, BPT19, CB15, Elb08]. Implementing [WYSJ08, GOvdR99]. implications [ABM^{+}19]. impossibility [PQ06]. Improved [CGKO11, Smi06]. in-browser [ABR13]. incidents [BT06]. increase [DFJ^{+}15]. incremental [BLB^{+}09]. indeed [vdM15]. Independence [PS10b, RB99]. Index [Ano92a, Ano93, Ano95a, Ano96a, Ano97, Ano98, Ano01, Ano02, Ano03, Ano04, Ano05, Ano06, Ano07a, Ano08, Ano09, Ano10, Ano11, Ano12, Ano13, Ano14, Ano15a, Ano16, Ano17, FBB^{+}18]. indexes [DFP^{+}13]. indifferentiable [BGH^{+}13]. indistinguishability [YWW^{+}09]. indistinguishable [BKP^{+}12]. individual [RCBM07]. Inductive [Bel03, RDDM10, Pau98]. industrial [RRAZ19]. Inference [BGSW11, HDC95, Bis12, WWJ04, WLJW07]. Inference-proof [BGSW11]. inferences [BDF^{+}12]. infinite [DR16]. infinite-state [DR16]. Information [Ano96b, Bec12, Bis95, BMPR05, BH09, FR06, Gra92, HBS16, HS04, LOS16, SM95, ABHS09, AAP12, AAC^{+}15, BCC^{+}19, BN07, Bis96, Bis12, BS02, BPR07, BvDS17, BS04, CGLZ03, CPPS07, CPP08, Cho12, CHM07, CMS09, CCP^{+}17, CFG^{+}17, DHRS11, DR16, DCMP16, GM10, GLZ11, GHRS05, HO05, HH09, KB11, LJM00, Low04b, MSB17, SG96, SAsC11, SST08, SVSM07, UTHP19, YT11, dACD^{+}16]. Information-flow [HBS16, dACD^{+}16]. information-hiding [CPP08]. information-theoretic [KB11]. infrastructure [BKA^{+}97, DLRS01, LG10]. Infrastructures [CMMV07, Pri06]. injecting [BKP^{+}12]. injection [KPS16]. inline [DJLP10]. insider [AWJS19]. Instant [CAB10]. institutions [ZHZY18]. instruction [Elb08]. integer [ZZW^{+}11]. integer-overflow-to-buffer-overflow [ZZW^{+}11]. integrating [MCL^{+}19]. integration [BBFS00]. Integrity [Mot92, NBM95, SM95, CC17, DFJ^{+}16, ST15, WD08]. Inter [GLP93, JAH^{+}16, LSMR16]. inter-frame [LSMR16]. inter-process [JAH^{+}16]. Inter-Realm [GLP93]. interactions [BCG^{+}02]. interactive [AAP12, BDJP10, LZ13, WYSJ08]. Interference [ZL95, BR05b, DHW04, RWW96, RS01]. internal [KSZ02]. Internet [GMR^{+}11, DGMS03, OMSH04, PS10a, SKEG14]. Interoperable [LG10]. interpolation [RVV17]. interpolation-based [RVV17]. Interpretation [MSC04, BPT19, GM10, MMHS13, PG09]. intersection [BHM14, VC05]. interval [WMF^{+}17]. intradomain [PV05]. intransitive [vdM15]. Introduction [Gut09, VG11, YGH08, YRY08]. intruder [BR04]. Intrusion [HFS98, LJM00, Mil99, SG02, EVK02, IBS03, IBS04, KSZ02, LFM^{+}02, MMF15, SSBW12, UC07, VK99, Yas02, YLZ05, ZGD04]. Invalid [GBL18]. investigations [BS04]. IoT [MGK^{+}17]. IP [DS99, MGK^{+}17]. iris [BA12]. islands [XC09]. ISO [BCM13]. ISO/IEC [BCM13]. isolation [GNDB16, LJM00]. Issue [Vis13, BSR97, CLM^{+}10, Cam10, MSB17, SV15]. issues [BMP18, BSCGS07]. Iterative [BM12].

Jannie [KNTU13]. Java [DJLP10]. Java-like [DJLP10]. JavaScript [HBS16]. JCS [CLM^{+}10, Cam10]. Johnny [HM13]. join [DFJ^{+}16]. Journal [BR09]. jump [LWZZ15].

Keccak [LŁA15]. Kernelized [TS93]. Key [BPS08, BMS95, CMMV07, DMM04, LG10, Mea92, Pri06, TA92, ABHS09, ABR13, CFL13, CLK04, CCD06, GT17, Gue09, GSAV18, HJT^{+}96, KAM08, KM98, LHM^{+}10, LYW^{+}10, Lop06, PQ03, PQ06, RFLW96, SPD^{+}10, TLY^{+}19]. Key-dependent [BPS08]. keys [DMM10, KAM08]. keystroke [BGE^{+}18]. know [BBS^{+}15]. Knowledge [Syv92, BU10, DLYZ11, LZ13, MMHS13]. knowledge-based [MMHS13]. known [LWZZ15].

Language [ES06, BN07, BFG10, CHM07, EVK02, LOS16, MW04, Smi06, SG02, VS99]. LaPadula [Mil96]. Large [GLP93, SA16, ZDM07]. large-scale [ZDM07]. Last [ZV19]. layer [JAH^{+}16, MGK^{+}17, RCBM07]. layered [GRKL15]. layout [XC09]. LDAP [LKWB06]. Leak [DTX09]. Leak-free [DTX09]. leakage [ABHS09, AAC^{+}15, BCC^{+}19, SG96, UTHP19]. leakages [VPZ16]. Learning [MTJ^{+}19, LMG^{+}18, RTWH11]. length [PB13, WDDN00]. LESS [SA16]. level [DS99, ZAF08]. Leveraging [HCM11, MvO11]. Life [LKWB06]. Life-cycle [LKWB06]. lightweight [FM98]. like [DJLP10]. limit [Cli00]. limited [JPSS16]. line [Wan06]. linear [HVL12]. Linguistic [BDGS16]. linked [Aba98, HvdM01]. Linux [GNDB16, GHRS05]. lists [PB13]. lived [ZDM07]. liveness [BPWS04]. local [Aba98, HvdM01]. location [SAsC11, SVA11]. location-based [SVA11]. locked [YTLD19]. logic [BBFR00, DDMP05, DMP03, GP12, HvdM01, Hal17, JH09, JPSS16]. logic-based [BBFR00]. Logical [BC92, SP03, CLK04, HvdM03, KNS08]. login [HM13]. logs [BT06]. long [ZDM07]. long-lived [ZDM07]. looping [Mal10]. Loose [DFJ^{+}15]. loss [DMV15]. Low [BBG^{+}18, ALP11, DS99, NR11]. low-bandwidth [NR11]. low-cost [ALP11]. Low-deterministic [BBG^{+}18]. low-level [DS99]. low-nondeterministic [BBG^{+}18]. Lowe [War05].

MAC [RCBM07]. Machine [AGHP14, HZL^{+}18, MCB13, RTWH11]. machine-checked [MCB13]. Machine-generated [AGHP14]. machines [SSM18]. mail [NZS05]. Maintaining [BMS95]. maintenance [AWJS19]. malicious [FBB^{+}18]. malleable [MSas13]. malware [LMG^{+}18, RTWH11, SCB^{+}19, UTHP19]. man [SSM18, ZV19]. man-in-the-middle [SSM18]. Manageable [Bro02]. Management [AJB93, ES06, FF11, KK95, Mea92, NBM95, AJS12, AH97, ACM04, BMBS10, Bec12, BVC^{+}14, BLL02, CG06, CFL13, CDM04, CCBE13, DMM04, GTY08, IBS04, KAM08, LWM03, LT17, LKWB06, MT08, NRW14, RFLW96, ST05, SWC07, SZP^{+}12, WD08]. managing [GMR^{+}11]. mandatory [AH97, AFHS09]. MANETs [ZMHT07]. market [CGLZ03]. Mashic [LSMR16]. Mashup [LSMR16]. masking [YLZ05]. matching [BEM^{+}13, BA12]. Mathematical [Gra92]. MC [CAFL10]. means [LG10, Lot97]. measure [KPS16]. Measures [LBF^{+}93]. Measuring [MMF15]. mechanism [CLK04, PSJ^{+}13, WS02]. mechanisms [AAC^{+}15, BDGS16]. mediated [CG06, DTX09]. mediating [JAH^{+}16]. mediation [ABFK03]. memory [GNDB16]. Merging [BSS97, CJP^{+}18, NBM95]. Message [Ano07b, BZ14, GTZ04, BPS08]. metaprogramming [LOS16]. metering [FM98]. method [Gut04b, HJT^{+}96, HBD19, RVV17, WLM01, WLM02]. methodology [CMMV07]. Methods [Mea92, CJP^{+}18]. Metric [CMMV07]. metrics [HZO^{+}13, SB18b]. micro [LWZZ15]. micro-data [LWZZ15]. microphones [GSAV18]. middle [SSM18]. Minimizing [VAGL09]. Minimum [LKAJ16]. Mining [SB18b, Cli00, LVA14, MSAV15, VC05]. misbehavior [RCBM07]. mitigate [ZZW^{+}11]. Mitigating [AWJS19, Möl05]. MITRE [LB96]. mix [JBH13]. mix-zone [JBH13]. mixtures [BFGS08, BFGS09]. mobile [AS15, ASV08, BDJP10, CMTB16, DLL^{+}18, HP00]. Model [AS92, BOS95, BJLT01, Bis95, DHRS11, DR16, JR04, Mil93b, Mot92, RS02, San92b, AHB00, BHČ04, CCD06, KL11, Low99, LHY^{+}15, Mil96, RAJ98, RB99, SB99, SVA11, TR11, ZRG08, Bis96]. Model-based [RS02]. Model-checking [DHRS11, DR16]. Modeling [BDF^{+}12, LFM^{+}02, NG08]. Modelling [GOvdR99, Tro93]. Models [NBM95, ASL96, AH00b, GL10, SRS^{+}02, TL07, YLZ05]. modern [GBL18]. Modular [CDM12, SPS^{+}14, HS04, KSS13]. monadic [HH09]. monitoring [DJLP10, JPSS16]. monotonic [ASL96]. multi [ASL96, BSSM^{+}07, CDE^{+}10, CT09, DB11, DDNP14, DRRW11, HZL^{+}18, JAH^{+}16, MS03, NSH14, RS16, Smi06, WLJW07, ZAF08, Zúq05]. multi-agent [DB11]. multi-dimensional [CT09, WLJW07]. multi-execution [DDNP14, RS16]. multi-factor [BSSM^{+}07]. multi-layer [JAH^{+}16]. multi-level [ZAF08]. multi-objective [DRRW11, HZL^{+}18]. multi-parent [ASL96]. multi-programmed [Zúq05]. multi-tenant [CDE^{+}10]. multi-threaded [MS03, NSH14, Smi06]. multiagent [HO05]. multiapplicative [SRS^{+}02]. Multicast [ZSXJ06, DMM04, DMM10, LS02, MR97]. Multilevel [AXR12, AJJ95, AJB93, JAK^{+}01, KK95, TS93, AHB00, BF99, RAJ98]. multimodal [ATI^{+}10]. Multiple [BMS95, CGM95, DFP^{+}13, KM98, SB18b, ZHZY18]. multisensor [SM05]. Multiset [DLMS04, BCLM05, CDL^{+}05]. multithreaded [DJLP10].

name [Aba98, HvdM01, SC00]. name-spaces [SC00]. Needham [War05]. needles [PS10a]. net [AH00b, ZHZY18]. NetSTAT [VK99]. network [BWJS18, BKP^{+}12, Gei13, NG08, RW97, UC07, VK99, Weh07, dC96]. network-based [VK99]. network-centric [RW97]. Networks [Vis13, ZSXJ06, AFHS09, CPPS07, Chu96, GT17, HZO^{+}13, HCM11, JBH13, KGA03, LGGO18, Lop06, Mea01, PS14, RCBM07, RH07, SPD^{+}10, TC16, TEML08, ZDM07]. Neural [Gei13]. NEXPTIME [ŢBEB08]. NEXPTIME-complete [ŢBEB08]. NFC [AS15]. no [CLD^{+}17]. no-root [CLD^{+}17]. noisy [GT17]. Non [BR05b, MG08, RWW96, ZL95, DHW04, KR03, LZ13, MB09, MSas13, RS01, SPS^{+}14, Wan06]. non-control [SPS^{+}14]. Non-delegatable [MG08]. non-disclosure [MB09]. non-interactive [LZ13]. Non-Interference [ZL95, BR05b, RWW96, DHW04, RS01]. non-malleable [MSas13]. non-repudiation [KR03, Wan06]. nondeterministic [BBG^{+}18]. Noninterference [BY95, Fol92, McL92, ABG04, Smi06, VS99, vdM15]. Norwegian [CW17, HSR08]. Notarized [GTY08]. note [BM99]. novel [SBP01]. NPATRL [MSC04]. NRL [MSC04]. number [Zúq05].

OBDD [CDF^{+}12]. obfuscated [BCC^{+}19]. obfuscation [PS14, PG09, PM18, PS10b]. Object [BOS95, TS93, BF99, DT97, HTS00]. Object-Oriented [TS93, DT97]. objective [CGKW17, DRRW11, HZL^{+}18]. objectives [BBK14]. obligations [CCBE13]. Obstruction [BBK14]. Obstruction-free [BBK14]. off [Wan06]. off-line [Wan06]. offs [DRRW11]. One [SM05, PB13]. open [BSCGS07, LAGM17, TEML08]. operating [SG96]. Operational [LBF^{+}93]. operations [Dug04, WMF^{+}17, YLZ05]. operator [MS05a, MS05b]. opinion [CYC17]. opportunities [BHSV14]. optimal [JPSS16, MCL^{+}19]. Optimality [Yah93]. optimization [DRRW11, LVA14]. optimized [HZL^{+}18]. Optimizing [JBH13]. options [BWJS18]. Oracle [SB99]. ORAM [MMRS19]. order [Coh03, Hal17]. orderings [BSS97]. Oriented [TS93, DT97, LHY^{+}15]. origin [BS04]. outsourced [BA12, CMTB16, CT09, DFP^{+}13, LCL^{+}15, MMRS19]. outsourcing [CDF^{+}11]. overflow [ZZW^{+}11]. ownership [NSMSN11].

P5 [SBS05]. packet [ZMHT07]. pairings [KM10]. Panoptis [SG02]. Paper [Ano15b]. Papers [Ano92b, Ano95b]. paradigm [DT97]. Paragon [BvDS17]. parallel [BR04]. parameter [BHSV14]. parent [ASL96]. Parity [WLJW07]. Parity-based [WLJW07]. parse [BMP18]. parsing [BMP18]. partial [ABHS09]. partitions [CFG^{+}17]. Party [Yah93, KSS13]. passive [CMS97]. passports [Sin11, LG10]. Password [BCC^{+}19, DGMT19, HJT^{+}96, MvO11]. Password-based [DGMT19]. passwords [vOT11]. Patching [BCFK15]. Paths [HDC95]. pattern [BEM^{+}13]. patterns [SR19, WDDN00]. payment [CMS97, HP00, YTLD19]. PCPOR [YY15]. Penetration [GG92]. Penetration-Resistant [GG92]. Per-session [DGMT19]. performance [BGE^{+}18, DTG16, IBS03, LŁA15, MCL^{+}19, MM96, SV15]. Performing [BT06]. persistent [TC16]. personal [MMRS19, MvO11, UTHP19]. personalization [DLL^{+}18]. pervasive [JBH13]. Petri [AH00b]. phishing [BDG^{+}10]. Phones [SSM18]. pi [BS18, DKR10]. pie [HNS^{+}14]. PILOT [BCC^{+}19]. PIN [BCC^{+}19]. PKCS#11 [CFL13, DKS10]. PKI [BVC^{+}14, BB14, NTU11, PS10a]. PKI-based [NTU11]. placement [HZL^{+}18]. plaintext [MSas13]. Planning [BDF09]. plans [CDM12]. platform [GMR^{+}11]. policies [AR12, BAVS19, BDP05, CFG^{+}17, CCBE13, HTML09, MBK^{+}15, MMHS13, NF18, RTV18, WS02]. Policy [AJS12, YGSY15, ASV08, BBI15, BW08, Bis12, CDE^{+}10, CMMV07, ES06, FF11, MB09, MP06, MT08, NRW14, ZAF08]. policy-based [CMMV07]. Policy-driven [AJS12]. Polynomial [BPWS04, HMQU09, HM10]. polynomial-time [HM10]. portscans [SHM02]. Power [San92b, ASL96, HM10, LMG^{+}18, MMF15, TL07]. Practical [BMV15, BvDS17, DCMP16, FBB^{+}18, Han11, NSMSN11, SHM02, BM12, DLL^{+}18]. practically [KSS13]. practice [DDNP14, Her09]. predictability [vOT11]. Preface [AL12, Atl11, BBW07, BCLP12, DG13, DJLS14, Foc06, Gon95, GP10, Got06, Got07, Got10, Gri11, JM92a, JM92b, JM93, JM95a, JM95b, Li12, LM92, MK93, Mil93a, Mil95a, MB12, PJ09, Sab08, Sab10, San92a, ZK06, AH00a, BR05a, BEN96, BMK97, Cup02, DM00, Foc05, Fol98, Fol99, Fri02, Gor05, Gut04a, JM96a, JM96b, JM97, JG02, LZ11, Lin99, Lin00, Mer97, Mil96, SS97, Sch03, Sch04, Syv01, Syv03, TvdR03]. Preprocessing [BW08]. preservation [DTG16]. Preserving [CWT14, BSSM^{+}07, BDG14, BLB^{+}09, DCMP16, Gut14, HSH11, HL01, LWZZ15, WMF^{+}17, ZHZY18]. prevent [HLS03]. preventing [BDF^{+}12, BBDG10, SKEG14]. Prevention [SVSM07, CRB19]. PRIME [ACK^{+}10]. principles [SS09]. Privacy [ABR13, ABM^{+}19, BSSM^{+}07, BLB^{+}09, MT08, SWH^{+}08, ZHZY18, ACDS08, ACK^{+}10, BGE^{+}18, BCL15, BCL97, BDG14, CWT14, CYC18, CDF^{+}11, DTG16, DKR09, DLYZ11, DRRW11, DCMP16, DLL^{+}18, GLS^{+}17, HSH11, HBD19, HS04, LGGO18, LYW^{+}10, LWZZ15, MMRS19, PS14, SZP^{+}12, TvdRO03, WMF^{+}17, YWW^{+}09]. privacy-aware [ACDS08]. privacy-enhanced [ACK^{+}10]. Privacy-preserving [BLB^{+}09, ZHZY18, BDG14, DCMP16, HSH11, WMF^{+}17]. Privacy-supporting [ABR13]. privacy-type [DKR09, HBD19]. Private [SRG97, AAC^{+}15, DFP^{+}13, DDL15]. proactive [MTJ^{+}19]. Probabilistic [MBK^{+}15, Shm04, VS99, ABG04, HM10, JPSS16, MTJ^{+}19, MMHS13, NS06, Smi06]. probability [KKN^{+}17]. Problem [HCH^{+}93, BCL15, BFO05, CGKW17, MSAV15, TC16]. problems [YT11]. procedure [HS05]. Process [FG95, RS01, ABG04, BCLM05, BS04, JAH^{+}16, WDDN00, YLZ05]. process-algebraic [ABG04]. processes [APRR14]. processing [AXR12, JAK^{+}01, MM96, RAJ98]. product [KSWH00]. profile [SVA11]. profiles [LWWJ02]. Programmable [HPS03]. programmed [Zúq05]. programming [BvDS17, HVL12]. programs [BBG^{+}18, BPR07, DJLP10, MS03, NSH14]. Project [ACK^{+}10]. Prometheus [CCP^{+}17]. Proof [GP12, BGSW11, BR05b, RRAZ19]. proofs [AGHP14, BU10, GL10, Hal17, MCB13, YY15]. Properties [FG95, Fol92, BJLT01, CDL06, DHRS11, DR16, DKR09, DMP03, Hal17, HBD19, NRW14, RDDM10]. property [PM18]. property-driven [PM18]. propositional [Bis12]. proprietary [DKS10]. Protecting [GLS^{+}17, BSSB06, DFP^{+}18]. Protection [AS92, Bis95, Bis96, HSR08, Mil93b, San92b, ATI^{+}10, CAFL10, DLRS01, WD08]. protections [SPS^{+}14]. Protocol [Mea92, MSC04, BHM14, BCLM05, BHČ04, CB15, CDL^{+}05, CW17, DDL15, GT17, GTZ04, Gut14, LS02, MR97, MCB13, MS05a, MS05b, NSC18, Pau01, SBS05, SBP01, War05, Yas02, ZF12, MSC04]. Protocols [SM95, Syv92, Yah93, AS15, ABK^{+}11, Bel03, Bla09, BBD^{+}05, BR05b, CPP08, Coh03, CEHM07, CDL06, DDMP05, DKR09, DGJN14, DMP03, DLMS04, FHG99, GRKL15, GJ03, GJ04, GEL98, Gut04b, HSH11, HLS03, HL01, JH09, KKN^{+}17, KSS13, KR03, KM10, KK16, Low98, Low99, Low04a, MVB10, NR11, PB13, Pau98, PQ03, PQ06, PV05, RS05, RVV17, RB99, SB18a, ŢBEB08, Wan06]. Prototype [EMO^{+}93]. Provable [GEL98]. Provably [BCM13, DJLP10, GNDB16, BMV15]. prove [DGJN14]. Proved [CB15]. provenance [AACP13]. Providing [BS04]. Proving [McL92, RB99, DMP03, FHG99]. provisioning [MP06]. pseudonym [CG06]. pseudorandom [CZ16]. Public [CMMV07, LG10, Pri06, TA92, YY15, CFG^{+}17, Lop06]. public-key [Lop06]. publication [CAB10, DGMS03, HLVA11]. Publicly [CZ16, NMP^{+}13, CGLZ03, LWZZ15]. publicly-known [LWZZ15]. publish [YSM14]. publish/subscribe [YSM14]. publishing [CDF^{+}12, DFJ^{+}15]. purpose [DGJN14]. purposes [KNTU13].

Qualified [MSZ06]. qualitative [Hal17]. quality [DRRW11, Zúq05]. quantification [VPZ16]. Quantifying [CMS09, CHM07]. Quantitative [AAP12, Hal17, YT11, ZMHT07]. quantity [Low04b]. quantum [SPD^{+}10]. queries [Bis12, BH09, DFJ^{+}16, VVB^{+}09, WLJW07]. Query [CT09, BW08, CWT14, CYC18, DFJ^{+}11, PS14].

random [Zúq05]. randomization [XC09]. randomness [BM99]. range [WLJW07]. rank [HS05]. rational [BHČ04]. RBAC [BAVS19, BGT15, LHY^{+}15, RTV18]. reader [NTU11]. real [BT06, SCB^{+}19]. real-time [BT06, SCB^{+}19]. Realm [GLP93]. Reasoning [KGA03, NF18, BHM14, BS18]. Reconstructing [CDM04]. reconstruction [HvdM03]. records [MMRS19]. Recoverable [NZS05]. recovery [CLK04, KM98]. redirection [Kre13]. Reducing [Hu92, VVB^{+}09]. reduction [MVB10]. refinement [BHM14, NF18, ZAF08]. Refining [SB18a]. reflection [WS02]. refreshments [BGSW11]. Registry [SAE^{+}05]. Regular [Ano15b]. regulating [BS02]. Rekeying [ZSXJ06]. Relating [BCLM05]. Relational [Mot92, WYSJ08]. Relations [Pau01, SP03]. relationship [RW97]. relationship-based [RW97]. release [BDF^{+}12, BS02, Cho12, HTML09]. released [YWW^{+}09]. reliability [LT17]. reliable [MR97, ZMHT07]. remapping [PSJ^{+}13]. Remote [CC17, AWJS19, MCL^{+}19]. repair [CC17]. repairing [BCM13]. repeater [SPD^{+}10]. replacement [LHM^{+}10]. replacing [KAM08]. Replicated [KK95, MM96]. replicated-architecture [MM96]. replication [SVSM07]. Report [LB96]. represent [BBI15]. reproducibility [SA16]. repudiation [KR03, Wan06]. Reputation [ST05, KNS08]. Reputation-based [ST05]. requests [VVB^{+}09]. Required [Cho12]. Requirements [AJB93, ABFK03]. Research [Ano92b, ZGD04, CLM^{+}10, Cam10, Pri06]. resiliency [CGKW17, ZV19]. resistance [KTV12]. Resistant [GG92]. Resource [Mil93b, BDJP10]. resources [JPSS16]. responding [OMSH04]. response [LFM^{+}02, SSBW12, ZGD04]. result [ACK^{+}10, Low99]. retrievability [YY15]. retrieval [DCMP16]. revisited [DGMT19, DDL15]. revocation [CAB10, HTML09, NTU11]. revoking [CMS97]. rewriting [BCLM05, CDL^{+}05, DLMS04]. RFID [ALP11, ABK^{+}11, DLYZ11, HSH11, Han11, KNTU13, NSMSN11, NTU11]. ring [LYW^{+}10]. Risk [HZL^{+}18, Mal10, SWC07, CPP08]. Risk-aware [HZL^{+}18]. RMP [VAGL09]. Robust [HJT^{+}96, MSZ06, RCBM07, ZDM07, Her09, LUV^{+}18]. Robustness [MSZ06]. Rogaway [MW04]. Role [BGS06, SB99, SM95, AJS12, AHB08, AR12, BBI15, LVA14, MSAV15, RTV18, TR11, UAV^{+}14, VAGL09, YGSY15]. Role-based [BGS06, SB99, AJS12, AR12, BBI15, TR11, VAGL09]. roles [SB18b]. rollbackability [ZGD04]. rollover [Gue09]. room [ABM^{+}19]. Root [CRB19, CLD^{+}17]. routing [PV05]. RSA [NZS05, ZDM07]. RSA-based [NZS05]. rule [LKAJ16, VC05]. run [SVSM07]. run-time [SVSM07]. runtime [HMQU09].

safely [HM13]. safety [AH00b]. sample [BDG14, Cli00]. sandboxes [PGK14]. sandboxing [LSMR16]. satisfiability [CGKW17]. SAv5 [CDWM19]. Scalable [AR12, BB14, SBS05]. scale [SA16, ZDM07]. scan [XVW^{+}06]. scenarios [Lot97, dSRCP17]. Scheduler [ZGDS13]. schedulers [HM10]. scheduling [BN07]. Schematic [AS92, San92b]. Scheme [ZSXJ06, LYW^{+}10, MSas13, NSMSN11, SZP^{+}12]. schemes [AGHP14, BM99, DTG16, Xu07]. Schroeder [War05]. Scriptless [HNS^{+}14]. scripts [DDNP14]. SDSI [Aba98, CEE^{+}01, HvdM01, JR04]. Search [MVB10, KM98, MRR^{+}18, PS14]. Search-space [MVB10]. Searchable [CGKO11, DRD11]. Second [HDC95]. SecPAL [BFG10]. Secrecy [ŢBEB08, CS13, KGA03]. secret [GT17, WMF^{+}17]. secrets [Pau01]. SECTET [AHB08]. SECTET-framework [AHB08]. Section [Ano15b, Mil99]. Secure [ABFK03, AJJ95, AJB93, BEM^{+}13, BN07, BF99, BC92, BA12, CG06, CMTB16, CPPS07, CDWM19, DDNP14, HLVA11, HVL12, KK95, PTMW10, RS16, SV03, TS93, TLY^{+}19, VC05, Yah93, ZSXJ06, AXR12, ALP11, AHB00, BHM14, BCG^{+}02, BMPR05, BMV15, BKA^{+}97, BB14, CB15, CCD06, CDF^{+}08, CDM12, DMM04, DMM10, FBB^{+}18, GNDB16, HJT^{+}96, JAK^{+}01, KSS13, LHM^{+}10, Lot97, MR97, PQ06, SR19, SC00, VIS96, ZAF08, dC96]. securely [LAGM17]. Securing [ALP11, Chu96, RH07, SST08]. Security [BSR97, CLM^{+}10, Cam10, DB11, FG95, FJ95, Gra92, Jac92, Kar00, LBF^{+}93, Mot92, SPD^{+}10, UAV^{+}14, Vis13, ZAF08, AS15, AH97, ACM04, ASV08, AFHS09, BPS08, BMP18, BBK14, BBI15, BCL15, BJLT01, BBG^{+}18, BCLM05, Bla09, BT06, BDP05, BBD^{+}05, BDGS16, BSS97, BPR07, CDE^{+}10, CGLZ03, CMMV07, CDL^{+}05, CJP^{+}18, CMPP14, CWT14, CEHM07, CCBE13, DDMP05, DKS10, DGMT19, DT97, DMP03, DLMS04, FHG99, FR06, FF11, FM98, GRKL15, GJ03, GLZ11, GL10, GEL98, GOvdR99, Gut04b, Gut14, Hal17, HH09, HLS03, HBS16, HL01, JH09, JPSS16, JAH^{+}16, KNTU13, KK16, Low98, Low99, MTJ^{+}19, Mal10, MBK^{+}15, MS03, MMHS13, MS96, MCB13, MGK^{+}17, MVB10, MGS^{+}17, MSB17, NSC18, NRW14, PB13, RS05, RVV17, RW97, RB99, RDDM10, SR19, SRS^{+}02, Sin97, Sin11]. security [SA16, SBP01, SV15, SB18a, TvdRO03, ŢBEB08, WS02, Xu07, Yas02, ZF12, dSRCP17, Ano92b, GHRS05, BR09]. Security-Enhanced [GHRS05]. security-sensitive [dSRCP17]. selection [BFGS08, BFGS09, SSBW12]. Selective [CDF^{+}11]. self [SVSM07]. self-replication [SVSM07]. semantic [AHB00, PS10b, RAJ98, FF11]. semantic-based [AHB00, RAJ98]. Semantics [JH09, PG09, Syv92, BFG10]. Semantics-based [PG09]. sensitive [BDF^{+}12, LFM^{+}02, SSBW12, dSRCP17]. sensor [CPPS07, Lop06]. sensors [KSZ02]. sentiment [CYC17]. separate [YGSY15]. Separation [Fol92, NBM95]. sequences [HFS98, QJ97]. Server [ATW97, CC17]. server-side [CC17]. Server-supported [ATW97]. Servers [CGM95, DRD11]. Service [BCL15, Mil93b, BDF09, BS02, CDM12, Mea01, Möl05, RFLW96]. services [CWT14, SVA11]. session [BCFK15, CRB19, CDF^{+}08, DGMT19]. set [BDG14, Elb08, VC05]. severity [VPZ16]. severity-based [VPZ16]. SGX [FBB^{+}18]. Shared [DRD11]. Sharing [LGGO18, DTG16, WMF^{+}17]. Sharing-habits [LGGO18]. shift [GLZ11]. Short [BFPV13, SSM18]. shuffle [LZ13]. SIA [CPPS07]. Side [KSWH00, CC17, GSAV18, KB11, MCL^{+}19, SCB^{+}19]. side-channel [GSAV18, KB11]. signature [AGHP14, LHM^{+}10]. signatures [ATW97, BFPV13, DTX09, ZDM07, NZS05]. signcryption [LYW^{+}10]. signing [NS06]. sill [HNS^{+}14]. similarity [BDG14, KPS16]. simple [CHM07]. simplifying [HL01]. simulatability [HMQU09]. simulator [SA16]. SINTRA [MM96]. SIP [GMR^{+}11]. SIP-based [GMR^{+}11]. size [Cli00]. skimming [Han11]. Skyband [CYC18]. smart [Bel03, BCG^{+}02, SRS^{+}02]. smartphones [GSAV18]. SMS [TEML08]. SMS-capable [TEML08]. SMT [AR12]. Snapshot [AJJ95]. social [LGGO18]. socio [BCL15, GBL18]. socio-technical [BCL15, GBL18]. soft [BFO05]. Software [CS05, AGHP14, WD08]. solutions [MT08]. solving [AR12]. Some [PQ03]. sound [KM10, LZ13, VIS96]. Soundness [ABHS09, BPS08, BU10, BWA10]. space [MVB10, XC09]. spaces [Aba98, CDL^{+}05, FHG99, HvdM01, KL11, SC00, SWH^{+}08]. spam [BFGS08, BFGS09, CYC17, SKEG14, GMR^{+}11]. spammer [CYC17]. spatio [TR11]. spatio-temporal [TR11]. spatiotemporal [SAsC11]. Special [BSR97, Mil99, MSB17, SV15, Vis13, CLM^{+}10, Cam10]. specific [SG02]. specification [BBFS00, CCBE13, DS97, MSC04, Sin97]. specifications [CB15]. Specifying [FJ95, YLZ05, GOvdR99]. SPIDER [GMR^{+}11]. SPIT [GMR^{+}11]. SPKI [CEE^{+}01, ES06, HvdM03, JR04]. SPKI/SDSI [CEE^{+}01, JR04]. Spread [PSJ^{+}13]. SPX [TA92]. SQL [KPS16, VVB^{+}09]. SQLiDDS [KPS16]. staged [LOS16]. standard [BCM13, BCL97]. standing [ZV19]. State [BY95, DR16, EVK02, KK16]. State-based [BY95, EVK02]. Stateful [GP12, APRR14]. Static [BBD^{+}05, JKK10, ZV19, BWA10, BBDG10, BDGS16, CHM07]. STATL [EVK02]. StatVerif [APRR14]. stealers [CCP^{+}17]. Stealing [HNS^{+}14, Kre13]. stealth [LMG^{+}18]. stealthy [SHM02]. steganography [GGS^{+}09]. step [SM05]. steps [SR19]. stock [CGLZ03]. storage [AFHS09, LT17]. storage-area [AFHS09]. Strand [FHG99, CDL^{+}05, KL11]. strategies [RCBM07]. Strategy [LT17, LWZZ15, PM18]. stream [AXR12]. Stricter [AJB93]. Strong [BCL97]. stronger [MvO11]. Structured [SC00]. study [CLD^{+}17, GP12, IBS03]. style [BPS08, GHPS13, JKK10]. Sub [CRB19]. Sub-session [CRB19]. subject [Low04a]. subscribe [YSM14]. Substring [MRR^{+}18]. sum [WLJW07]. support [MMF15]. Supported [FJ95, ATW97]. Supporting [DFP^{+}13, TS93, ABR13, Wan06]. suppression [BM12]. surveillance [TC16]. surveilled [SWH^{+}08]. survey [CDL06, NR11, Sin11]. Surviving [BWJS18]. suspicious [WDDN00]. switching [GHPS13]. Symbolic [BU16, DKR10, MS05a, MS05b, AR12, BU10, MCB13]. symmetric [CGKO11, YWW^{+}09]. Symposium [Ano92b]. synthesis [ZV19]. Synthesising [ZRG08]. System [ES06, EMO^{+}93, Tro93, ACDS08, Bis12, BKP^{+}12, DDMP05, DLRS01, HFS98, IBS03, MP06, MM96, Shm04, Sin11, SA16, UC07, VK99, VIS96]. Systematic [BMP18, KSS13]. Systems [AJB93, GLP93, GG92, WL93, ALP11, AAP12, ATI^{+}10, AH97, ACM04, BGE^{+}18, BAVS19, Bec12, BFM05, BSSB06, CMS97, DR16, DTG16, DB11, DS97, DT97, HTS00, HO05, HP00, IBS04, JHS96, KAM08, KNS08, LCL^{+}15, LJM00, LAGM17, Lot97, MS96, MG08, NTU11, PTMW10, SS10, SG96, Sin97, SV15, VPZ16, YSD13, YSM14, ZRG08, ZAF08, Zúq05]. Syverson [BHČ04].

table [DCMP16]. tables [HLVA11]. tags [HSH11, KNTU13]. taint [JKK10]. taint-style [JKK10]. Take [Bis95, Bis96]. Take-Grant [Bis95, Bis96]. tampering [BHSV14]. taps [GSAV18]. targets [CYC17]. task [AWJS19]. taxonomy [BSCGS07]. TCP [SCB^{+}19]. Technical [LB96, BCL15, GBL18]. technique [BGJ03]. techniques [BFGS08, BFGS09, BR05b, DRRW11, RB99, WMF^{+}17, XVW^{+}06]. Telephony [GMR^{+}11, SKEG14]. template [ATI^{+}10]. Temporal [BBFS00, JPSS16, MSAV15, RTV18, SB18b, TR11, UAV^{+}14]. tenant [CDE^{+}10]. tests [Gut04b]. TG [CAFL10]. TG/MC [CAFL10]. Theft [Bis95]. their [CZ16, KSS13]. Theorems [Jac92, MW04]. theoretic [KB11, SKEG14]. theories [BM12]. Theory [DDNP14, GG92, GP12, Her09, LT17, TL07]. there [GLZ11]. threaded [MS03, NSH14, Smi06]. Threat [FF11, Lot97, AWJS19, BT06, ZZW^{+}11]. threats [Mal10, YWW^{+}09]. threshold [BM99]. throughput [MR97]. ticket [HTS00]. ticket-based [HTS00]. Time [Hu92, KKN^{+}17, Tro93, BPT19, BT06, BBDG10, HM10, LWWJ02, SVSM07, SCB^{+}19, YTLD19]. time-locked [YTLD19]. Timed [CEHM07, Kre13]. Timing [Hu92, Wra92, MCL^{+}19]. TLS [KL11]. tokens [Han11]. tolerant [WLM01, WLM02]. touching [HNS^{+}14]. trace [DHRS11, DR16, RDDM10]. trace-based [DHRS11, DR16]. Traces [McL92]. Tracking [RRAZ19]. trade [DRRW11]. trade-offs [DRRW11]. traditional [CJP^{+}18]. traffic [Weh07]. Transaction [KK95, JAK^{+}01, MM96, RAJ98]. transactional [ACM15]. transactions [BGSW11, RAJ98, YTLD19]. transfer [NSMSN11]. transform [DPV09]. transformations [HL01]. Translation [GGS^{+}09, ABR13]. Translation-based [GGS^{+}09]. transparent [RS16, Wan06]. trapping [MS96]. treatment [YSM14]. Tree [HSH11, BMP18, CFG^{+}17, LŁA15]. Tree-based [HSH11]. triggers [CCF98]. TripleMon [JAH^{+}16]. Trojan [Gei13]. Trust [CLM^{+}10, Cam10, CGM95, ES06, SM95, Bec12, BVC^{+}14, CDM04, DB11, LWM03, NRW14, ST05, SZP^{+}12, WD08, ZMHT07]. trust-augmented [SZP^{+}12]. Trusted [NBM95, Gue09, SPD^{+}10]. trustees [CMS97]. TTP [Wan06]. tutorial [Möl05]. Two [AJJ95, Yah93, KSS13, Pau01, SAE^{+}05]. Two-Party [Yah93, KSS13]. Two-Snapshot [AJJ95]. Type [CFL13, Dug04, BHM14, BBDG10, DKR09, HLS03, HBD19, PQ06, VIS96, ZZW^{+}11]. Type-based [CFL13, Dug04]. Typed [PGK14, LOS16]. Typed-based [PGK14]. Types [GJ04, BHM14, BFM07, SV03]. typing [BCC^{+}19, GJ03]. typings [Smi06].

UC-soundness [BPS08]. unauthorized [HSR08]. unbounded [HBD19, PB13]. uncertain [SAsC11]. uncertainty [CAFL10]. unclassified [GM10]. unconditionally [ALP11, CCD06]. Understanding [DLL^{+}18]. Unified [Mot92, CCD06]. uniform [BS02]. unifying [MS03]. Union [BHM14]. universal [BU16]. Unleashing [Lop06]. unpatchable [BWJS18]. unspoofable [NR11]. untrusted [DRD11, GLS^{+}17, MvO11, SV03]. Unwinding [Mil95b]. update [BGSW11]. updates [BFM05]. upon [PQ03]. URA97 [SB99]. usability [MGS^{+}17, KNTU13]. used [CDL06]. User [BSCGS07, BMS95, NTU11, CWT14, DLL^{+}18, LHY^{+}15, SB99, TC16]. User-aided [NTU11]. user-oriented [LHY^{+}15]. user-role [SB99]. Using [Cli00, JPSS16, KSZ02, McL92, SCB^{+}19, TA92, WS02, WCJS97, ZZW^{+}11, AFHS09, BAVS19, BT06, CG06, CCF98, CAB10, DRRW11, FF11, GT17, GSAV18, Hal17, HFS98, HZO^{+}13, KPS16, KM10, LWWJ02, LMG^{+}18, MMHS13, MSC04, RTWH11, SG02, TvdRO03, WYSJ08, Weh07, XC09, YLZ05]. utility [DFJ^{+}15].

validation [BBD^{+}05]. value [BDF^{+}12]. variable [WDDN00]. variable-length [WDDN00]. various [XVW^{+}06]. vehicular [RH07]. Verifiable [NZS05, NMP^{+}13, ZHZY18]. Verification [APRR14, PB13, AGHP14, Bel03, Bla09, CT09, Coh03, HBD19, KR03, NSH14, NSC18, NRW14, PGK14, RVV17]. Verified [BGH^{+}13, SRS^{+}02, MSB17, ZRG08, dACD^{+}16]. verifier [DGJN14]. verify [KNTU13]. Verifying [BDJP10, BPT19, DKR09, GRKL15, GHRS05, BDF09, BB14, Pau98]. version [BCG^{+}02]. versus [DRRW11]. via [AJJ95, BMS95, CFG^{+}17, Elb08, SST08, SKEG14, ZV19]. video [QJ97, SWH^{+}08]. videos [BCC^{+}19]. View [BC92, BGSW11]. views [YWW^{+}09]. virtual [CDE^{+}10, HZL^{+}18]. visibility [CDF^{+}12]. voice [SSM18]. VoIP [SS10]. Volume [Ano95a, Ano96a, Ano97, Ano98, Ano01, Ano02, Ano03, Ano04, Ano05, Ano06, Ano07a, Ano08, Ano09, Ano10, Ano11, Ano12, Ano13, Ano14, Ano15a, Ano16, Ano17, LB96]. voting [CW17, DKR09, SZP^{+}12]. vs [WDDN00]. vulnerabilities [BHSV14, BWJS18, DS99, JKK10, RS02, ZGDS13]. Vulnerability [HCH^{+}93, BFO05, HZO^{+}13].

wall [ACM04]. Warm [MCL^{+}19]. watermarking [CS05]. wavelet [DPV09]. Web [BSR97, BHSV14, CRB19, DDNP14, GLS^{+}17, JKK10, VVB^{+}09, BS02, BVC^{+}14, HCM11, PS14, PGK14, SRG97, WCJS97]. web-based [VVB^{+}09]. WebCallerID [HCM11]. WebDAV [CAB10]. WebInject [CCP^{+}17]. WebInject-based [CCP^{+}17]. website [BBDLM14]. well [Kre13]. well-timed [Kre13]. while [DFP^{+}18]. Wide [BSR97, WCJS97]. Window [EMO^{+}93]. Windows [SAE^{+}05]. wireless [GT17, JBH13, Lop06]. wiretapping [BBS^{+}15]. within [BR04]. without [CFG^{+}17, HNS^{+}14, YGSY15]. workflow [AH97, AH00b, ACM04, CCF98, CGKW17, GOvdR99, TvdRO03]. workflows [AHB00, dSRCP17]. Working [FJ95]. World [BSR97, WCJS97]. worm [XVW^{+}06]. worms [Weh07]. wrappers [SV03]. Write [TS93]. Write-Up [TS93]. WWW [BSR97].

X.509 [BMP18, LKWB06]. XML [BFM05, DGK^{+}04]. XPath [BH09].

Yahalom [Pau01]. Yao [BPS08]. Yao-style [BPS08].

zero [BU10, DLYZ11, LZ13]. zero-knowledge [BU10, DLYZ11, LZ13]. zone [JBH13].

References

Alvim:2015:ILD

[AAC^{+}15] Mário S. Alvim, Miguel E. Andrés, Konstantinos Chatzikokolakis, Pierpaolo Degano, and Catuscia Palamidessi. On the information leakage of differentially-private mechanisms. Journal of Computer Security, 23(4):427–469, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Acar:2013:CCP

[AACP13] Umut A. Acar, Amal Ahmed, James Cheney, and Roly Perera. A core calculus for provenance. Journal of Computer Security, 21(6): 919–969, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Alvim:2012:QIF

[AAP12] Mário S. Alvim, Miguel E. Andrés, and Catuscia Palamidessi. Quantitative information flow in interactive systems. Journal of Computer Security, 20(1):3–50, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Abadi:1998:SLL

[Aba98] Martín Abadi. On SDSI’s linked local name spaces. Journal of Computer Security, 6(1–2):3–21, ???? 1998. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Altenschmidt:2003:SMR

[ABFK03] C. Altenschmidt, J. Biskup, U. Flegel, and Y. Karabulut. Secure mediation: requirements, design, and architecture. Journal of Computer Security, 11(3):365–398, ???? 2003. CODEN JC-SIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Aldini:2004:PAA

[ABG04] Alessandro Aldini, Mario Bravetti, and Roberto Gorrieri. A process-algebraic approach for the analysis of probabilistic noninterference. Journal of Computer Security, 12(2):191–245, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Adao:2009:SCF

[ABHS09] Pedro Adão, Gergei Bana, Jonathan Herzog, and Andre Scedrov. Soundness and completeness of formal encryption: The cases of key cycles and partial information leakage. Journal of Computer Security, 17(5):737–797, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Avoine:2011:FAR

[ABK^{+}11] Gildas Avoine, Muhammed Ali Bingöl, Süleyman Kardaş, Cédric Lauradoux, and Benjamin Martin. A framework for analyzing RFID distance bounding protocols. Journal of Computer Security, 19(2):289–317, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Armknecht:2019:PIR

[ABM^{+}19] Frederik Armknecht, Zinaida Benenson, Philipp Morgner, Christian Müller, and Christian Riess. Privacy implications of room climate data. Journal of Computer Security, 27(1):113–136, ???? 2019. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Aimeur:2006:BEC

[ABO06] Esma Aïmeur, Gilles Brassard, and Flavien Serge Mani Onana. Blind electronic commerce. Journal of Computer Security, 14(6): 535–559, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Arapinis:2013:PSC

[ABR13] Myrto Arapinis, Sergiu Bursuc, and Mark Ryan. Privacysupporting cloud computing by in-browser key translation. Journal of Computer Security, 21(6):847–880, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Ardagna:2008:PAA

[ACDS08] C. A. Ardagna, M. Cremonini, S. De Capitani di Vimercati, and P. Samarati. A privacy-aware access control system. Journal of Computer Security, 16(4):369–397, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Ardagna:2010:ECP

[ACK^{+}10] Claudio A. Ardagna, Jan Camenisch, Markulf Kohlweiss, Ronald Leenes, Gregory Neven, Bart Priem, Pierangela Samarati, Dieter Sommer, and Mario Verdicchio. Exploiting cryptography for privacy-enhanced access control: A result of the PRIME Project. Journal of Computer Security, 18(1):123–160, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Atluri:2004:CWS

[ACM04] Vijayalakshmi Atluri, Soon Ae Chun, and Pietro Mazzoleni. Chinese wall security for decentralized workflow management systems. Journal of Computer Security, 12(6):799–840, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

AlBouna:2015:ATD

[ACM15] Bechara Al Bouna, Chris Clifton, and Qutaibah Malluhi. Anonymizing transactional datasets. Journal of Computer Security, 23(1):89–106, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Aziz:2009:CSA

[AFHS09] Benjamin Aziz, Simon N. Foley, John Herbert, and Garret Swart. Configuring storage-area networks using mandatory security. Journal of Computer Security, 17(2):191–210, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Akinyele:2014:MGA

[AGHP14] Joseph A. Akinyele, Matthew Green, Susan Hohenberger, and Matthew Pagano. Machine-generated algorithms, proofs and software for the batch verification of digital signature schemes. Journal of Computer Security, 22(6):867–912, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Atluri:1997:EMD

[AH97] Vijayalakshmi Atluri and Wei-Kuang Huang. Enforcing mandatory and discretionary security in workflow management systems. Journal of Computer Security, 5(4):303–339, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Atluri:2000:GEP

[AH00a] Vijay Atluri and John Hale. Guest Editor’s preface. Journal of Computer Security, 8(4):241–242, ???? 2000. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Atluri:2000:PNB

[AH00b] Vijayalakshmi Atluri and Wei-Kuang Huang. A Petri net based safety analysis of workflow authorization models. Journal of Computer Security, 8(2–3):209–240, ???? 2000. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Atluri:2000:SBE

[AHB00] Vijayalakshmi Atluri, Wei-Kuang Huang, and Elisa Bertino. A semantic-based execution model for multilevel secure workflows. Journal of Computer Security, 8(1):3–41, ???? 2000. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Alam:2008:CBR

[AHB08] Muhammad Alam, Michael Hafner, and Ruth Breu. Constraint based role based access control in the SECTET-framework. Journal of Computer Security, 16(2):223–260, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Atluri:1993:ASC

[AJB93] Vijayalakshmi Atluri, Sushil Jajodia, and Elisa Bertino. Achieving stricter correctness requirements in multilevel secure database management systems. Journal of Computer Security, 2(4):311– 351, ???? 1993. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Ammann:1995:CCS

[AJJ95] Paul Ammann, Frank Jaeckle, and Sushil Jajodia. Concurrency control in a secure multilevel database via a two-snapshot algorithm. Journal of Computer Security, 3(2–3):87–113, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Ahn:2012:PDR

[AJS12] Gail-Joon Ahn, Jing Jin, and Mohamed Shehab. Policy-driven role-based access management for ad-hoc collaboration. Journal of Computer Security, 20(2–3):223–257, ???? 2012. CODEN JC-SIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Armando:2012:P

[AL12] Alessandro Armando and Gavin Lowe. Preface. Journal of Computer Security, 20(1):1, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Alomair:2011:SLC

[ALP11] Basel Alomair, Loukas Lazos, and Radha Poovendran. Securing low-cost RFID systems: An unconditionally secure approach. Journal of Computer Security, 19(2):229–257, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:1992:AI

[Ano92a] Anonymous. Author index. Journal of Computer Security, 1(3– 4):413, ???? 1992. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:1992:CPE

[Ano92b] Anonymous. Call for papers: European Symposium on Research in Computer Security. Journal of Computer Security, 1(1):131, ???? 1992. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:1993:AI

[Ano93] Anonymous. Author index. Journal of Computer Security, 2(4): 353–354, ???? 1993. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:1995:AIV

[Ano95a] Anonymous. Author index volume 3. Journal of Computer Security, 3(4):325–326, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:1995:CP

[Ano95b] Anonymous. Call for papers. Journal of Computer Security, 3(4): 323–324, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:1996:AIV

[Ano96a] Anonymous. Author index volume 4 (1996). Journal of Computer Security, 4(4):361–362, ???? 1996. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:1996:IA

[Ano96b] Anonymous. Information for authors. Journal of Computer Security, 4(1):113–119, ???? 1996. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:1997:AIV

[Ano97] Anonymous. Author index volume 5 (1997). Journal of Computer Security, 5(4):383–384, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:1998:AIV

[Ano98] Anonymous. Author index volume 6 (1998). Journal of Computer Security, 6(4):287, ???? 1998. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:2001:AIV

[Ano01] Anonymous. Author index volume 9 (2001). Journal of Computer Security, 9(4):339–340, ???? 2001. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:2002:AIV

[Ano02] Anonymous. Author index volume 10 (2002). Journal of Computer Security, 10(4):433–434, ???? 2002. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:2003:AIV

[Ano03] Anonymous. Author index volume 11 (2003). Journal of Computer Security, 11(4):723–725, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:2004:AIV

[Ano04] Anonymous. Author index volume 12 (2004). Journal of Computer Security, 12(6):933–935, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:2005:AIV

[Ano05] Anonymous. Author index volume 13 (2005). Journal of Computer Security, 13(6):905–907, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:2006:AIV

[Ano06] Anonymous. Author index volume 14 (2006). Journal of Computer Security, 14(6):625–626, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:2007:AIV

[Ano07a] Anonymous. Author index volume 15 (2007). Journal of Computer Security, 15(6):717–719, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:2007:MGE

[Ano07b] Anonymous. Message from the Guest Editors. Journal of Computer Security, 15(1):1–2, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:2008:AIV

[Ano08] Anonymous. Author index volume 16 (2008). Journal of Computer Security, 16(6):791–793, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:2009:AIV

[Ano09] Anonymous. Author index volume 17 (2009). Journal of Computer Security, 17(6):969–971, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:2010:AIV

[Ano10] Anonymous. Author index volume 18 (2010). Journal of Computer Security, 18(6):1301–1305, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:2011:AIV

[Ano11] Anonymous. Author index volume 19 (2011). Journal of Computer Security, 19(6):1173–1176, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:2012:AIV

[Ano12] Anonymous. Author index volume 20 (2012). Journal of Computer Security, 20(6):765–767, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:2013:AIV

[Ano13] Anonymous. Author index volume 21 (2013). Journal of Computer Security, 21(6):1009–1012, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:2014:AIV

[Ano14] Anonymous. Author index volume 22 (2014). Journal of Computer Security, 22(6):1051–1054, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:2015:AIV

[Ano15a] Anonymous. Author index volume 23 (2015). Journal of Computer Security, 23(6):789–791, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:2015:RPS

[Ano15b] Anonymous. Regular paper section. Journal of Computer Security, 23(5):639, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:2016:AIV

[Ano16] Anonymous. Author index volume 24 (2016). Journal of Computer Security, 24(6):839–841, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Anonymous:2017:AIV

[Ano17] Anonymous. Author index volume 25 (2017). Journal of Computer Security, 25(6):631–633, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Arapinis:2014:SVS

[APRR14] Myrto Arapinis, Joshua Phillips, Eike Ritter, and Mark D. Ryan. StatVerif: Verification of stateful processes. Journal of Computer Security, 22(5):743–821, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Armando:2012:SAS

[AR12] Alessandro Armando and Silvio Ranise. Scalable automated symbolic analysis of administrative role-based access control policies by SMT solving. Journal of Computer Security, 20(4):309–352, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Ammann:1992:ESP

[AS92] Paul E. Ammann and Ravi S. Sandhu. The extended schematic protection model. Journal of Computer Security, 1(3–4):335–383, ???? 1992. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Alshehri:2015:FFS

[AS15] Ali Alshehri and Steve Schneider. A formal framework for security analysis of NFC mobile coupon protocols. Journal of Computer Security, 23(6):685–707, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Ammann:1996:EPM

[ASL96] Paul Ammann, Ravi S. Sandhu, and Richard Lipton. The expressive power of multi-parent creation in monotonic access control models. Journal of Computer Security, 4(2–3):149–165, ???? 1996. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Atluri:2008:ESP

[ASV08] Vijayalakshmi Atluri, Heechang Shin, and Jaideep Vaidya. Efficient security policy enforcement for the mobile environment. Journal of Computer Security, 16(4):439–475, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Argyropoulos:2010:BTP

[ATI^{+}10] Savvas Argyropoulos, Dimitrios Tzovaras, Dimosthenis Ioannidis, Yannis Damousis, Michael G. Strintzis, Martin Braun, and Serge Boverie. Biometric template protection in multimodal authentication systems based on error correcting codes. Journal of Computer Security, 18(1):161–185, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Atluri:2011:P

[Atl11] Vijay Atluri. Preface. Journal of Computer Security, 19(3):365, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Asokan:1997:SSS

[ATW97] N. Asokan, G. Tsudik, and M. Waidner. Server-supported signatures. Journal of Computer Security, 5(1):91–108, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Alhebaishi:2019:MIT

[AWJS19] Nawaf Alhebaishi, Lingyu Wang, Sushil Jajodia, and Anoop Singhal. Mitigating the insider threat of remote administrators in clouds through maintenance task assignments. Journal of Computer Security, 27(4):427–458, ???? 2019. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Adaikkalavan:2012:MSD

[AXR12] Raman Adaikkalavan, Xing Xie, and Indrakshi Ray. Multilevel secure data stream processing: Architecture and implementation. Journal of Computer Security, 20(5):547–581, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Blanton:2012:SOC

[BA12] Marina Blanton and Mehrdad Aliasgari. Secure outsourced computation of iris matching. Journal of Computer Security, 20(2–3): 259–305, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Batra:2019:DAP

[BAVS19] Gunjan Batra, Vijayalakshmi Atluri, Jaideep Vaidya, and Shamik Sural. Deploying ABAC policies using RBAC systems. Journal of Computer Security, 27(4):483–506, ???? 2019. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Buchmann:2014:TMS

[BB14] Nicolas Buchmann and Harald Baier. Towards a more secure and scalable verifying PKI of eMRTD. Journal of Computer Security, 22(6):1025–1049, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bodei:2005:SVS

[BBD^{+}05] Chiara Bodei, Mikael Buchholtz, Pierpaolo Degano, Flemming Nielson, and Hanne Riis Nielson. Static validation of security protocols. Journal of Computer Security, 13(3):347–390, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bodei:2010:DPT

[BBDG10] Chiara Bodei, Linda Brodo, Pierpaolo Degano, and Han Gao. Detecting and preventing type flaws at static time. Journal of Computer Security, 18(2):229–264, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bansal:2014:DCA

[BBDLM14] Chetan Bansal, Karthikeyan Bhargavan, Antoine Delignat-Lavaud, and Sergio Maffeis. Discovering concrete attacks on website authorization by formal analysis. Journal of Computer Security, 22(4):601–657, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bertino:2000:LBA

[BBFR00] Elisa Bertino, Francesco Buccafurri, Elena Ferrari, and Pasquale Rullo. A logic-based approach for enforcing access control. Journal of Computer Security, 8(2–3):109–139, ???? 2000. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bertino:2000:TAB

[BBFS00] Elisa Bertino, Piero Andrea Bonatti, Elena Ferrari, and Maria Luisa Sapino. Temporal authorization bases: From specification to integration. Journal of Computer Security, 8(4):309–353, ???? 2000. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bischof:2018:LDS

[BBG^{+}18] Simon Bischof, Joachim Breitner, Jürgen Graf, Martin Hecker, Martin Mohr, and Gregor Snelting. Low-deterministic security for low-nondeterministic programs. Journal of Computer Security, 26(3):335–366, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Belim:2015:AGR

[BBI15] Sergey Belim, Nadezda Bogachenko, and Evgeniy Ilushechkin. An analysis of graphs that represent a role-based security policy hierarchy. Journal of Computer Security, 23(5):641–657, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Basin:2014:OFA

[BBK14] David Basin, Samuel J. Burri, and Günter Karjoth. Obstructionfree authorization enforcement: Aligning security and business objectives. Journal of Computer Security, 22(5):661–698, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bates:2015:AWK

[BBS^{+}15] Adam Bates, Kevin R. B. Butler, Micah Sherr, Clay Shields, Patrick Traynor, and Dan Wallach. Accountable wiretapping — or — I know they can hear you now. Journal of Computer Security, 23(2):167–195, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Backes:2007:P

[BBW07] Michael Backes, David Basin, and Michael Waidner. Preface. Journal of Computer Security, 15(6):561, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bieber:1992:LVS

[BC92] Pierre Bieber and Frédéric Cuppens. A logical view of secure dependencies. Journal of Computer Security, 1(1):99–129, ???? 1992. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Balagani:2019:PPP

[BCC^{+}19] Kiran Balagani, Matteo Cardaioli, Mauro Conti, Paolo Gasti, Martin Georgiev, Tristan Gurtler, Daniele Lain, Charissa Miller, Kendall Molas, Nikita Samarin, Eugen Saraci, Gene Tsudik, and Lynn Wu. PILOT: Password and PIN information leakage from obfuscated typing videos. Journal of Computer Security, 27(4): 405–425, ???? 2019. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bugliesi:2015:CPB

[BCFK15] Michele Bugliesi, Stefano Calzavara, Riccardo Focardi, and Wilayat Khan. CookiExt: Patching the browser against session hijacking attacks. Journal of Computer Security, 23(4):509–537, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bieber:2002:CSI

[BCG^{+}02] P. Bieber, J. Cazin, P. Girard, J.-L. Lanet, V. Wiels, and G. Zanon. Checking secure interactions of smart card applets: extended version. Journal of Computer Security, 10(4):369–398, ???? 2002. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bergadano:1997:SAP

[BCL97] F. Bergadano, B. Crispo, and M. Lomas. Strong authentication and privacy with standard browsers. Journal of Computer Security, 5(3):191–212, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bella:2015:SSP

[BCL15] Giampaolo Bella, Paul Curzon, and Gabriele Lenzini. Service security and privacy as a socio-technical problem. Journal of Computer Security, 23(5):563–585, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bistarelli:2005:RMR

[BCLM05] Stefano Bistarelli, Iliano Cervesato, Gabriele Lenzini, and Fabio Martinelli. Relating multiset rewriting and process algebras for security protocol analysis. Journal of Computer Security, 13(1): 3–47, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Barthe:2012:P

[BCLP12] Gilles Barthe, Jorge Cuellar, Javier Lopez, and Alexander Pretschner. Preface. Journal of Computer Security, 20(4):307–308, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Basin:2013:PRI

[BCM13] David Basin, Cas Cremers, and Simon Meier. Provably repairing the ISO/IEC 9798 standard for entity authentication. Journal of Computer Security, 21(6):817–846, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bartoletti:2009:PVS

[BDF09] Massimo Bartoletti, Pierpaolo Degano, and Gian Luigi Ferrari. Planning and verifying service composition. Journal of Computer Security, 17(5):799–837, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bezzi:2012:MPI

[BDF^{+}12] Michele Bezzi, Sabrina De Capitani di Vimercati, Sara Foresti, Giovanni Livraga, Pierangela Samarati, and Roberto Sassi. Modeling and preventing inferences from sensitive value distributions in data release. Journal of Computer Security, 20(4):393–436, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bergholz:2010:NFA

[BDG^{+}10] André Bergholz, Jan De Beer, Sebastian Glahn, Marie-Francine Moens, Gerhard Paaß, and Siehyun Strobel. New filtering approaches for phishing email. Journal of Computer Security, 18 (1):7–35, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Blundo:2014:EEP

[BDG14] Carlo Blundo, Emiliano De Cristofaro, and Paolo Gasti. EsPRESSO: Efficient privacy-preserving evaluation of sample set similarity. Journal of Computer Security, 22(3):355–381, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bodei:2016:CAS

[BDGS16] Chiara Bodei, Pierpaolo Degano, Letterio Galletta, and Francesco Salvatori. Context-aware security: Linguistic mechanisms and static analysis. Journal of Computer Security, 24(4):427–477, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Besson:2010:VRA

[BDJP10] Frédéric Besson, Guillaume Dufay, Thomas Jensen, and David Pichardie. Verifying resource access control on mobile interactive devices. Journal of Computer Security, 18(6):971–998, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bodei:2005:CSP

[BDP05] Chiara Bodei, Pierpaolo Degano, and Corrado Priami. Checking security policies through an enhanced Control Flow Analysis. Journal of Computer Security, 13(1):49–85, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Becker:2012:IFT

[Bec12] Moritz Y. Becker. Information flow in trust management systems. Journal of Computer Security, 20(6):677–708, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bella:2003:IVS

[Bel03] Giampaolo Bella. Inductive verification of smart card protocols. Journal of Computer Security, 11(1):87–132, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Baron:2013:SPM

[BEM^{+}13] Joshua Baron, Karim El Defrawy, Kirill Minkovich, Rafail Ostrovsky, and Eric Tressler. 5PM: Secure pattern matching. Journal of Computer Security, 21(5):601–625, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bertino:1996:GEP

[BEN96] Elisa Bertino, Gérard Eizenberg, and Roger M. Needham. Guesteditors’ preface. Journal of Computer Security, 4(1):1–2, ???? 1996. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bertino:1999:SOD

[BF99] Elisa Bertino and Elena Ferrari. Secure object deletion and garbage collection in multilevel object bases. Journal of Computer Security, 7(4):257–285, ???? 1999. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Becker:2010:SDS

[BFG10] Moritz Y. Becker, Cédric Fournet, and Andrew D. Gordon. Sec-PAL: Design and semantics of a decentralized authorization language. Journal of Computer Security, 18(4):619–665, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Belsis:2008:AEF

[BFGS08] Petros Belsis, Kostas Fragos, Stefanos Gritzalis, and Christos Skourlas. Applying effective feature selection techniques with hierarchical mixtures of experts for spam classification. Journal of Computer Security, 16(6):761–790, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Belsis:2009:AEF

[BFGS09] Petros Belsis, Kostas Fragos, Stefanos Gritzalis, and Christos Skourlas. Applying effective feature selection techniques with hierarchical mixtures of experts for spam classification. Journal of Computer Security, 17(3):239–268, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bertino:2005:ACU

[BFM05] E. Bertino, E. Ferrari, and G. Mella. An approach to cooperative updates of XML documents in distributed systems. Journal of Computer Security, 13(2):191–242, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bugliesi:2007:DTA

[BFM07] Michele Bugliesi, Riccardo Focardi, and Matteo Maffei. Dynamic types for authentication. Journal of Computer Security, 15(6): 563–617, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bistarelli:2005:SCB

[BFO05] Stefano Bistarelli, Simon N. Foley, and Barry O’Sullivan. A soft constraint-based approach to the cascade vulnerability problem. Journal of Computer Security, 13(5):699–720, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Blazy:2013:SBS

[BFPV13] Olivier Blazy, Georg Fuchsbauer, David Pointcheval, and Damien Vergnaud. Short blind signatures. Journal of Computer Security, 21(5):627–661, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Balagani:2018:IAC

[BGE^{+}18] Kiran S. Balagani, Paolo Gasti, Aaron Elliott, Azriel Richardson, and Mike O’Neal. The impact of application context on privacy and performance of keystroke authentication systems. Journal of Computer Security, 26(4):543–556, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Barthe:2013:VIH

[BGH^{+}13] Gilles Barthe, Benjamin Grégoire, Sylvain Heraud, Federico Olmedo, and Santiago Zanella-Béguelin. Verified indifferentiable hashing into elliptic curves. Journal of Computer Security, 21(6): 881–917, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Barbara:2003:CBC

[BGJ03] Daniel Barbará, Rajni Goel, and Sushil Jajodia. A checksumbased corruption detection technique. Journal of Computer Security, 11(3):315–329, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Braghin:2006:RBA

[BGS06] Chiara Braghin, Daniele Gorla, and Vladimiro Sassone. Role-based access control for a distributed calculus. Journal of Computer Security, 14(2):113–155, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Biskup:2011:IPV

[BGSW11] Joachim Biskup, Christian Gogolin, Jens Seiler, and Torben Weibert. Inference-proof view update transactions with forwarded refreshments. Journal of Computer Security, 19(3):487–529, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bonatti:2015:EDR

[BGT15] Piero Bonatti, Clemente Galdi, and Davide Torres. Event-driven RBAC. Journal of Computer Security, 23(6):709–757, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bottcher:2009:IDA

[BH09] Stefan Böttcher and Rita Hartel. Information disclosure by answers to XPath queries. Journal of Computer Security, 17(1): 69–99, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Buttyan:2004:FMR

[BHČ04] Levente Buttyán, Jean-Pierre Hubaux, and Srdjan Čapkun. A formal model of rational exchange and its application to the analysis of Syverson’s protocol. Journal of Computer Security, 12(3–4): 551–587, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Backes:2014:UIR

[BHM14] Michael Backes, Cătălin Hriţcu, and Matteo Maffei. Union, intersection and refinement types and reasoning about type disjointness for secure protocol implementations. Journal of Computer Security, 22(2):301–353, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bisht:2014:ADP

[BHSV14] Prithvi Bisht, Timothy Hinrichs, Nazari Skrupsky, and V. N. Venkatakrishnan. Automated detection of parameter tampering opportunities and vulnerabilities in web applications. Journal of Computer Security, 22(3):415–465, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bishop:1995:TIT

[Bis95] Matt Bishop. Theft of information in the take-Grant protection model. Journal of Computer Security, 3(4):283–308, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bishop:1996:CIF

[Bis96] Matt Bishop. Conspiracy and information flow in the Take-Grant Protection Model. Journal of Computer Security, 4(4):331–359, ???? 1996. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Biskup:2012:DPA

[Bis12] Joachim Biskup. Dynamic policy adaptation for inference control of queries to a propositional information system. Journal of Computer Security, 20(5):509–546, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Besson:2001:MCS

[BJLT01] Frédéric Besson, Thomas Jensen, Daniel Le Métayer, and Tommy Thorn. Model checking security properties of control flow graphs. Journal of Computer Security, 9(3):217–250, ???? 2001. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bryce:1997:CEI

[BKA^{+}97] Ciarán Bryce, Winfried Kühnhauser, Rémy Amouroux, Mauricio López, and Harry Rudnik. CWASAR: a European infrastructure for secure electronic commerce. Journal of Computer Security, 5(3):225–235, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bowen:2012:SGI

[BKP^{+}12] Brian M. Bowen, Vasileios P. Kemerlis, Pratap Prabhu, Angelos D. Keromytis, and Salvatore J. Stolfo. A system for generating and injecting indistinguishable network decoys. Journal of Computer Security, 20(2–3):199–221, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Blanchet:2009:AVC

[Bla09] Bruno Blanchet. Automatic verification of correspondences for security protocols. Journal of Computer Security, 17(4):363–434, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Byun:2009:PPI

[BLB^{+}09] Ji-Won Byun, Tiancheng Li, Elisa Bertino, Ninghui Li, and Yonglak Sohn. Privacy-preserving incremental data dissemination. Journal of Computer Security, 17(1):43–68, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Buldas:2002:ECA

[BLL02] Ahto Buldas, Peeter Laud, and Helger Lipmaa. Eliminating counterevidence with applications to accountable certificate management. Journal of Computer Security, 10(3):273–296, ???? 2002. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Blundo:1999:NRD

[BM99] Carlo Blundo and Barbara Masucci. A note on the randomness in dynamic threshold schemes. Journal of Computer Security, 7 (1):73–85, ???? 1999. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bielova:2012:IES

[BM12] Nataliia Bielova and Fabio Massacci. Iterative enforcement by suppression: Towards practical enforcement theories. Journal of Computer Security, 20(1):51–79, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Baldwin:2010:AFI

[BMBS10] Adrian Baldwin, Marco Casassa Mont, Yolanta Beres, and Simon Shiu. Assurance for federated identity management. Journal of Computer Security, 18(4):541–572, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bertino:1997:GEP

[BMK97] Elisa Bertino, Emilio Montolivo, and Helmut Kurth. Guest editors’ preface. Journal of Computer Security, 5(1):1–2, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Barenghi:2018:SPX

[BMP18] Alessandro Barenghi, Nicholas Mainardi, and Gerardo Pelosi. Systematic parsing of X.509: Eradicating security issues with a parse tree. Journal of Computer Security, 26(6):817–849, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bossi:2005:IFS

[BMPR05] Annalisa Bossi, Damiano Macedonio, Carla Piazza, and Sabina Rossi. Information flow in secure contexts. Journal of Computer Security, 13(3):391–422, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Blundo:1995:MKD

[BMS95] C. Blundo, Luiz A. Frota Mattos, and D. R. Stinson. Multiple key distribution maintaining user anonymity via broadcast channels. Journal of Computer Security, 3(4):309–322, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Boureanu:2015:PPS

[BMV15] Ioana Boureanu, Aikaterini Mitrokotsa, and Serge Vaudenay. Practical and provably secure distance-bounding. Journal of Computer Security, 23(2):229–257, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Barthe:2007:SIF

[BN07] Gilles Barthe and Leonor Prensa Nieto. Secure information flow for a concurrent language with scheduling. Journal of Computer Security, 15(6):647–689, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bertino:1995:EAM

[BOS95] Elisa Bertino, Fabio Origgi, and Pierangela Samarati. An extended authorization model for object databases. Journal of Computer Security, 3(2–3):169–206, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bossi:2007:CIF

[BPR07] Annalisa Bossi, Carla Piazza, and Sabina Rossi. Compositional information flow security for concurrent programs. Journal of Computer Security, 15(3):373–416, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Backes:2008:KDM

[BPS08] Michael Backes, Birgit Pfitzmann, and Andre Scedrov. Keydependent message security under active attacks — BRSIM/UCsoundness of Dolev–Yao-style encryption with key cycles. Journal of Computer Security, 16(5):497–530, ???? 2008. CODEN JC-SIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Blazy:2019:VCT

[BPT19] Sandrine Blazy, David Pichardie, and Alix Trieu. Verifying constant-time implementations by abstract interpretation. Journal of Computer Security, 27(1):137–163, ???? 2019. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Backes:2004:PL

[BPWS04] Michael Backes, Birgit Pfitzmann, Michael Waidner, and Michael Steiner. Polynomial liveness. Journal of Computer Security, 12 (3–4):589–617, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Broadfoot:2004:EAW

[BR04] P. J. Broadfoot and A. W. Roscoe. Embedding agents within the intruder to detect parallel attacks. Journal of Computer Security, 12(3–4):379–408, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bella:2005:GEP

[BR05a] Giampaolo Bella and Peter Ryan. Guest Editors’ preface. Journal of Computer Security, 13(5):697, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bugliesi:2005:NIP

[BR05b] Michele Bugliesi and Sabina Rossi. Non-interference proof techniques for the analysis of cryptographic protocols. Journal of Computer Security, 13(1):87–113, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bella:2009:JCS

[BR09] Giampaolo Bella and Peter Y. A. Ryan. Journal of Computer Security. Journal of Computer Security, 17(3):237, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Brose:2002:MAC

[Bro02] Gerald Brose. Manageable access control for CORBA. Journal of Computer Security, 10(4):301–337, ???? 2002. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bonatti:2002:UFR

[BS02] Piero A. Bonatti and Pierangela Samarati. A uniform framework for regulating service access and information release on the Web.

Journal of Computer Security, 10(3):241–271, ???? 2002. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Buchholz:2004:PPO

[BS04] Florian P. Buchholz and Clay Shields. Providing process origin information to aid in computer forensic investigations. Journal of Computer Security, 12(5):753–776, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Blanchet:2018:ARE

[BS18] Bruno Blanchet and Ben Smyth. Automated reasoning for equivalences in the applied pi calculus with barriers. Journal of Computer Security, 26(3):367–422, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bhargav-Spantzel:2007:UCT

[BSCGS07] Abhilasha Bhargav-Spantzel, Jan Camenisch, Thomas Gross, and Dieter Sommer. User centricity: A taxonomy and open issues. Journal of Computer Security, 15(5):493–527, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bertino:1997:SIS

[BSR97] Elisa Bertino, Pierangela Samarati, and Gian Paolo Rossi. Special issue on security in the World Wide Web (WWW). Journal of Computer Security, 5(3):189–190, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bonatti:1997:MHS

[BSS97] P. A. Bonatti, M. L. Sapino, and V. S. Subrahmanian. Merging heterogeneous security orderings. Journal of Computer Security, 5(1):3–29, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bhargav-Spantzel:2006:EPD

[BSSB06] Abhilasha Bhargav-Spantzel, Anna C. Squicciarini, and Elisa Bertino. Establishing and protecting digital identity in federation systems. Journal of Computer Security, 14(3):269–300, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bhargav-Spantzel:2007:PPM

[BSSM^{+}07] Abhilasha Bhargav-Spantzel, Anna C. Squicciarini, Shimon Modi, Matthew Young, Elisa Bertino, and Stephen J. Elliott. Privacy preserving multi-factor authentication with biometrics. Journal of Computer Security, 15(5):529–560, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Blyth:2006:PRT

[BT06] Andrew Blyth and Paula Thomas. Performing real-time threat assessment of security incidents using data fusion of IDS logs. Journal of Computer Security, 14(6):513–534, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Backes:2010:CSS

[BU10] Michael Backes and Dominique Unruh. Computational soundness of symbolic zero-knowledge proofs. Journal of Computer Security, 18(6):1077–1155, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bohl:2016:SUC

[BU16] Florian Böhl and Dominique Unruh. Symbolic universal composability. Journal of Computer Security, 24(1):1–38, ???? 2016. CO-DEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Braun:2014:CTM

[BVC^{+}14] Johannes Braun, Florian Volk, Jiska Classen, Johannes Buchmann, and Max Mühlhäuser. CA trust management for the Web PKI. Journal of Computer Security, 22(6):913–959, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Broberg:2017:PPP

[BvDS17] Niklas Broberg, Bart van Delft, and David Sands. Paragon — practical programming with information flow control. Journal of Computer Security, 25(4–5):323–365, ???? 2017. CODEN JC-SIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Biskup:2008:PCQ

[BW08] Joachim Biskup and Lena Wiese. Preprocessing for controlled query evaluation with availability policy. Journal of Computer Security, 16(4):477–494, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Baudet:2010:GAC

[BWA10] Mathieu Baudet, Bogdan Warinschi, and Martín Abadi. Guessing attacks and the computational soundness of static equivalence.

Journal of Computer Security, 18(5):909–968, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Borbor:2018:SUV

[BWJS18] Daniel Borbor, Lingyu Wang, Sushil Jajodia, and Anoop Singhal. Surviving unpatchable vulnerabilities through heterogeneous network hardening options. Journal of Computer Security, 26(6): 761–789, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bevier:1995:SBA

[BY95] William R. Bevier and William D. Young. A state-based approach to noninterference. Journal of Computer Security, 3(1):55–70, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Backes:2014:MGE

[BZ14] Michael Backes and Steve Zdancewic. Message from the Guest Editors. Journal of Computer Security, 22(5):659–660, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Chadwick:2010:ICR

[CAB10] David W. Chadwick, Sean Antony, and Rune Bjerk. Instant certificate revocation and publication using WebDAV. Journal of Computer Security, 18(3):475–496, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Conrad:2010:AUT

[CAFL10] James R. Conrad, Jim Alves-Foss, and Sauchi Stephen Lee. Analyzing uncertainty in TG protection graphs with TG/MC. Journal of Computer Security, 18(5):667–699, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Campolargo:2010:JSI

[Cam10] Mário Campolargo. JCS special issue on EU-funded ICT research on trust and security: Foreword. Journal of Computer Security, 18(1):??, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Cade:2015:PGI

[CB15] David Cadé and Bruno Blanchet. Proved generation of implementations from computationally secure protocol specifications. Journal of Computer Security, 23(3):331–402, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Chen:2017:RDI

[CC17] Bo Chen and Reza Curtmola. Remote data integrity checking with server-side repair. Journal of Computer Security, 25(6):537–584, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Cuppens:2013:FSM

[CCBE13] Frédéric Cuppens, Nora Cuppens-Boulahia, and Yehia Elrakaiby. Formal specification and management of security policies with collective group obligations. Journal of Computer Security, 21(1): 149–190, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Cimato:2006:UMU

[CCD06] Stelvio Cimato, Antonella Cresti, and Paolo D’Arco. A unified model for unconditionally secure key distribution. Journal of Computer Security, 14(1):45–64, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Casati:1998:EWA

[CCF98] Fabio Casati, Silvana Castano, and Maria Grazia Fugini. Enforcing workflow authorization constraints using triggers. Journal of Computer Security, 6(4):257–285, ???? 1998. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Continella:2017:PAW

[CCP^{+}17] Andrea Continella, Michele Carminati, Mario Polino, Andrea Lanzi, Stefano Zanero, and Federico Maggi. Prometheus: Analyzing WebInject-based information stealers. Journal of Computer Security, 25(2):117–137, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Cabuk:2010:TAS

[CDE^{+}10] Serdar Cabuk, Chris I. Dalton, Konrad Eriksson, Dirk Kuhlmann, HariGovind V. Ramasamy, Gianluca Ramunno, Ahmad-Reza Sadeghi, Matthias Schunter, and Christian Stüble. Towards automated security policy enforcement in multi-tenant virtual data centers. Journal of Computer Security, 18(1):89–121, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Castano:1997:ADG

[CDF97] S. Castano, S. De Capitani di Vimercati, and M. G. Fugini. Automated derivation of global authorizations for database federations. Journal of Computer Security, 5(4):271–301, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Corin:2008:SCS

[CDF^{+}08] Ricardo Corin, Pierre-Malo Deniélou, Cédric Fournet, Karthikeyan Bhargavan, and James Leifer. A secure compiler for session abstractions. Journal of Computer Security, 16(5):573–636, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Ciriani:2011:SDO

[CDF^{+}11] Valentina Ciriani, Sabrina De Capitani di Vimercati, Sara Foresti, Sushil Jajodia, Stefano Paraboschi, and Pierangela Samarati. Selective data outsourcing for enforcing privacy. Journal of Computer Security, 19(3):531–566, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Ciriani:2012:OAE

[CDF^{+}12] Valentina Ciriani, Sabrina De Capitani di Vimercati, Sara Foresti, Giovanni Livraga, and Pierangela Samarati. An OBDD approach to enforce confidentiality and visibility constraints in data publishing. Journal of Computer Security, 20(5):463–508, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Cervesato:2005:CBS

[CDL^{+}05] Iliano Cervesato, Nancy A. Durgin, Patrick D. Lincoln, John C. Mitchell, and Andre Scedrov. A comparison between strand spaces and multiset rewriting for security protocol analysis. Journal of Computer Security, 13(2):265–316, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Cortier:2006:SAP

[CDL06] Véronique Cortier, Stéphanie Delaune, and Pascal Lafourcade. A survey of algebraic properties used in cryptographic protocols. Journal of Computer Security, 14(1):1–43, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Chander:2004:RTM

[CDM04] Ajay Chander, Drew Dean, and John C. Mitchell. Reconstructing trust management. Journal of Computer Security, 12(1):131–164, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Costa:2012:MPS

[CDM12] Gabriele Costa, Pierpaolo Degano, and Fabio Martinelli. Modular plans for secure service composition. Journal of Computer Security, 20(1):81–117, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Cremers:2019:SAG

[CDWM19] Cas Cremers, Martin Dehnel-Wild, and Kevin Milner. Secure authentication in the grid: a formal analysis of DNP3 SAv5. Journal of Computer Security, 27(2):203–232, ???? 2019. CODEN JC-SIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Clarke:2001:CCD

[CEE^{+}01] Dwaine Clarke, Jean-Emile Elien, Carl Ellison, Matt Fredette, Alexander Morcos, and Ronald L. Rivest. Certificate chain discovery in SPKI/SDSI. Journal of Computer Security, 9(4):285–322, ???? 2001. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Corin:2007:TAS

[CEHM07] R. Corin, S. Etalle, P. H. Hartel, and A. Mader. Timed analysis of security protocols. Journal of Computer Security, 15(6):619–645, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Crampton:2017:CEI

[CFG^{+}17] Jason Crampton, Naomi Farley, Gregory Gutin, Mark Jones, and Bertram Poettering. Cryptographic enforcement of information flow policies without public information via tree partitions. Journal of Computer Security, 25(6):511–535, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Centenaro:2013:TBA

[CFL13] Matteo Centenaro, Riccardo Focardi, and Flaminia L. Luccio. Type-based analysis of key management in PKCS#11 cryptographic devices. Journal of Computer Security, 21(6):971–1007, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Candebat:2006:SPM

[CG06] Thibault Candebat and David Gray. Secure pseudonym management using mediated identity-based encryption. Journal of Computer Security, 14(3):249–267, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Curtmola:2011:SSE

[CGKO11] Reza Curtmola, Juan Garay, Seny Kamara, and Rafail Ostrovsky. Searchable symmetric encryption: Improved definitions and efficient constructions. Journal of Computer Security, 19(5):895–934, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Crampton:2017:BOW

[CGKW17] Jason Crampton, Gregory Gutin, Daniel Karapetyan, and Rémi Watrigant. The bi-objective workflow satisfiability problem and workflow resiliency. Journal of Computer Security, 25(1):83–115, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Campbell:2003:ECP

[CGLZ03] Katherine Campbell, Lawrence A. Gordon, Martin P. Loeb, and Lei Zhou. The economic cost of publicly announced information security breaches: empirical evidence from the stock market. Journal of Computer Security, 11(3):431–448, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Chen:1995:DTA

[CGM95] Liqun Chen, Dieter Gollmann, and Christopher J. Mitchell. Distributing trust amongst multiple authentication servers. Journal of Computer Security, 3(4):255–267, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Clark:2007:SAQ

[CHM07] David Clark, Sebastian Hunt, and Pasquale Malacaria. A static analysis for quantifying information flow in a simple imperative language. Journal of Computer Security, 15(3):321–371, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Chong:2012:RIR

[Cho12] Stephen Chong. Required information release. Journal of Computer Security, 20(6):637–676, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Chuang:1996:SAN

[Chu96] Shaw-Cheng Chuang. Securing ATM networks. Journal of Computer Security, 4(4):289–329, ???? 1996. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Chakraborty:2018:HAD

[CJP^{+}18] Tanmoy Chakraborty, Sushil Jajodia, Noseong Park, Andrea Pugliese, Edoardo Serra, and V. S. Subrahmanian. Hybrid adversarial defense: Merging honeypots and traditional security methods. Journal of Computer Security, 26(5):615–645, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Cheng:2017:SFN

[CLD^{+}17] Yao Cheng, Yingjiu Li, Robert Deng, Lingyun Ying, and Wei He. A study on a feasible no-root approach on Android. Journal of Computer Security, 25(3):231–253, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Clifton:2000:USS

[Cli00] Chris Clifton. Using sample size to limit exposure to data mining. Journal of Computer Security, 8(4):281–307, ???? 2000. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Cho:2004:GKR

[CLK04] Taenam Cho, Sang-Ho Lee, and Won Kim. A group key recovery mechanism based on logical key hierarchy. Journal of Computer Security, 12(5):711–736, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Camenisch:2010:JSI

[CLM^{+}10] Jan Camenisch, Javier Lopez, Fabio Massacci, Massimo Ciscato, and Thomas Skordas. JCS special issue on EU-funded ICT research on trust and security. Journal of Computer Security, 18 (1):1–5, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Casola:2007:PBM

[CMMV07] Valentina Casola, Antonino Mazzeo, Nicola Mazzocca, and Valeria Vittorini. A policy-based methodology for security evaluation: A security metric for public key infrastructures. Journal of Computer Security, 15(2):197–229, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Chatzikokolakis:2014:FAS

[CMPP14] Konstantinos Chatzikokolakis, Sebastian Alexander Mödersheim, Catuscia Palamidessi, and Jun Pang. Foundational aspects of security. Journal of Computer Security, 22(2):201–202, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Camenisch:1997:DPS

[CMS97] Jan Camenisch, Ueli Maurer, and Markus Stadler. Digital payment systems with passive anonymity-revoking trustees. Journal of Computer Security, 5(1):69–89, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Clarkson:2009:QIF

[CMS09] Michael R. Clarkson, Andrew C. Myers, and Fred B. Schneider. Quantifying information flow with beliefs. Journal of Computer Security, 17(5):655–701, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Carter:2016:SOG

[CMTB16] Henry Carter, Benjamin Mood, Patrick Traynor, and Kevin Butler. Secure outsourced garbled circuit evaluation for mobile devices. Journal of Computer Security, 24(2):137–180, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Cohen:2003:FOV

[Coh03] Ernie Cohen. First-order verification of cryptographic protocols. Journal of Computer Security, 11(2):189–216, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Chatzikokolakis:2008:BRI

[CPP08] Konstantinos Chatzikokolakis, Catuscia Palamidessi, and Prakash Panangaden. On the Bayes risk in information-hiding protocols. Journal of Computer Security, 16(5):531–571, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Chan:2007:SSI

[CPPS07] Haowen Chan, Adrian Perrig, Bartosz Przydatek, and Dawn Song. SIA: Secure information aggregation in sensor networks. Journal of Computer Security, 15(1):69–102, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Calzavara:2019:SSH

[CRB19] Stefano Calzavara, Alvise Rabitti, and Michele Bugliesi. Subsession hijacking on the web: Root causes and prevention. Journal of Computer Security, 27(2):233–257, ???? 2019. CODEN JC-SIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Collberg:2005:SWF

[CS05] Christian Collberg and Tapas Ranjan Sahoo. Software watermarking in the frequency domain: Implementation, analysis, and attacks. Journal of Computer Security, 13(5):721–755, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Clarkson:2010:H

[CS10] Michael R. Clarkson and Fred B. Schneider. Hyperproperties. Journal of Computer Security, 18(6):1157–1210, ???? 2010. CO-DEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Cortier:2013:AFH

[CS13] Véronique Cortier and Ben Smyth. Attacking and fixing Helios: An analysis of ballot secrecy. Journal of Computer Security, 21(1): 89–148, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Cheng:2009:QAV

[CT09] Weiwei Cheng and Kian-Lee Tan. Query assurance verification for outsourced multi-dimensional databases. Journal of Computer Security, 17(1):101–126, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Cuppens:2002:GEP

[Cup02] Frédéric Cuppens. Guest Editor’s preface. Journal of Computer Security, 10(4):299–300, ???? 2002. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Cortier:2017:FAN

[CW17] Véronique Cortier and Cyrille Wiedling. A formal analysis of the Norwegian E-voting protocol. Journal of Computer Security, 25 (1):21–57, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Chen:2014:PUQ

[CWT14] Yen-Chung Chen, Yu-Sung Wu, and Wen-Guey Tzeng. Preserving user query privacy in cloud-based security services. Journal of Computer Security, 22(6):997–1024, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Choo:2017:DOS

[CYC17] Euijin Choo, Ting Yu, and Min Chi. Detecting opinion spammer groups and spam targets through community discovery and sentiment analysis. Journal of Computer Security, 25(3):283–318, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Chen:2018:KSQ

[CYC18] Ling Chen, Ting Yu, and Rada Chirkova. k-Skyband query answering with differential privacy. Journal of Computer Security, 26(5):647–676, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Chen:2016:PEP

[CZ16] Yu Chen and Zongyang Zhang. Publicly evaluable pseudorandom functions and their applications. Journal of Computer Security, 24(2):289–320, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

deAmorim:2016:VIF

[dACD^{+}16] Arthur Azevedo de Amorim, Nathan Collins, André DeHon, Delphine Demange, Cătălin Hriţcu, David Pichardie, Benjamin C. Pierce, Randy Pollack, and Andrew Tolmach. A verified information-flow architecture. Journal of Computer Security, 24 (6):689–734, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Damiani:2008:GE

[Dam08] Ernesto Damiani. Guest editorial. Journal of Computer Security, 16(4):367–368, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

DeAngelis:2011:SAT

[DB11] David DeAngelis and K. Suzanne Barber. Security applications of trust in multi-agent systems. Journal of Computer Security, 19 (1):57–99, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

dAusbourg:1996:CCD

[dC96] Bruno d’Ausbourg and Christel Calas. Controlling causal dependencies over a secure network. Journal of Computer Security, 4 (1):3–25, ???? 1996. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

DiCrescenzo:2016:PPP

[DCMP16] Giovanni Di Crescenzo, Debra L. Cook, Allen McIntosh, and Euthimios Panagos. Practical and privacy-preserving information retrieval from a database table. Journal of Computer Security, 24(4): 479–506, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Dreier:2015:BFP

[DDL15] Jannik Dreier, Jean-Guillaume Dumas, and Pascal Lafourcade. Brandt’s fully private auction protocol revisited. Journal of Computer Security, 23(5):587–610, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Datta:2005:DSC

[DDMP05] Anupam Datta, Ante Derek, John C. Mitchell, and Dusko Pavlovic. A derivation system and compositional logic for security protocols. Journal of Computer Security, 13(3):423–482, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

DeGroef:2014:SME

[DDNP14] Willem De Groef, Dominique Devriese, Nick Nikiforakis, and Frank Piessens. Secure multi-execution of web scripts: Theory and practice. Journal of Computer Security, 22(4):469–509, ????

2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

DeCapitanidiVimercati:2011:AED

[DFJ^{+}11] Sabrina De Capitani di Vimercati, Sara Foresti, Sushil Jajodia, Stefano Paraboschi, and Pierangela Samarati. Authorization enforcement in distributed query evaluation. Journal of Computer Security, 19(4):751–794, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

DeCapitanidiVimercati:2015:LAI

[DFJ^{+}15] Sabrina De Capitani di Vimercati, Sara Foresti, Sushil Jajodia, Giovanni Livraga, Stefano Paraboschi, and Pierangela Samarati. Loose associations to increase utility in data publishing. Journal of Computer Security, 23(1):59–88, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

DeCapitanidiVimercati:2016:EIC

[DFJ^{+}16] Sabrina De Capitani di Vimercati, Sara Foresti, Sushil Jajodia, Stefano Paraboschi, and Pierangela Samarati. Efficient integrity checks for join queries in the cloud. Journal of Computer Security, 24(3):347–378, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

DeCapitanidiVimercati:2013:SCM

[DFP^{+}13] Sabrina De Capitani di Vimercati, Sara Foresti, Stefano Paraboschi, Gerardo Pelosi, and Pierangela Samarati. Supporting concurrency and multiple indexes in private access to outsourced data. Journal of Computer Security, 21(3):425–461, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

DeCapitanidiVimercati:2018:EAW

[DFP^{+}18] Sabrina De Capitani di Vimercati, Sara Foresti, Stefano Paraboschi, Gerardo Pelosi, and Pierangela Samarati. Enforcing authorizations while protecting access confidentiality. Journal of Computer Security, 26(2):143–175, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Degano:2013:P

[DG13] Pierpaolo Degano and Joshua D. Guttman. Preface. Journal of Computer Security, 21(6):779–780, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Dupressoir:2014:GGP

[DGJN14] François Dupressoir, Andrew D. Gordon, Jan Jürjens, and David A. Naumann. Guiding a general-purpose C verifier to prove cryptographic protocols. Journal of Computer Security, 22(5):823– 866, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Devanbu:2004:FAX

[DGK^{+}04] Premkumar Devanbu, Michael Gertz, April Kwong, Charles Martel, Glen Nuckolls, and Stuart G. Stubblebine. Flexible authentication of XML documents. Journal of Computer Security, 12(6): 841–864, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Devanbu:2003:ADP

[DGMS03] Premkumar Devanbu, Michael Gertz, Charles Martel, and Stuart G. Stubblebine. Authentic data publication over the Internet. Journal of Computer Security, 11(3):291–314, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Demay:2019:PSS

[DGMT19] Grégory Demay, Peter Gaži, Ueli Maurer, and Björn Tackmann. Per-session security: Password-based cryptography revisited. Journal of Computer Security, 27(1):75–111, ???? 2019. CO-DEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

DSouza:2011:MCT

[DHRS11] Deepak D’Souza, Raveendra Holla, K. R. Raghavendra, and Barbara Sprick. Model-checking trace-based information flow properties. Journal of Computer Security, 19(1):101–138, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

DiPierro:2004:ANI

[DHW04] Alessandra Di Pierro, Chris Hankin, and Herbert Wiklicky. Approximate non-interference. Journal of Computer Security, 12(1): 37–81, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Dam:2010:PCI

[DJLP10] Mads Dam, Bart Jacobs, Andreas Lundblad, and Frank Piessens. Provably correct inline monitoring for multithreaded Java-like programs. Journal of Computer Security, 18(1):37–59, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Desmet:2014:P

[DJLS14] Lieven Desmet, Martin Johns, Benjamin Livshits, and Andrei Sabelfeld. Preface. Journal of Computer Security, 22(4):467–468, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Delaune:2009:VPT

[DKR09] Stéphanie Delaune, Steve Kremer, and Mark Ryan. Verifying privacy-type properties of electronic voting protocols. Journal of Computer Security, 17(4):435–487, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Delaune:2010:SBA

[DKR10] Stéphanie Delaune, Steve Kremer, and Mark D. Ryan. Symbolic bisimulation for the applied pi calculus. Journal of Computer Security, 18(2):317–377, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Delaune:2010:FSA

[DKS10] Stéphanie Delaune, Steve Kremer, and Graham Steel. Formal security analysis of PKCS#11 and proprietary extensions. Journal of Computer Security, 18(6):1211–1245, ???? 2010. CODEN JC-SIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Diao:2018:AMU

[DLL^{+}18] Wenrui Diao, Rui Liu, Xiangyu Liu, Zhe Zhou, Zhou Li, and Kehuan Zhang. Accessing mobile user’s privacy based on IME personalization: Understanding and practical attacks. Journal of Computer Security, 26(3):283–309, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Durgin:2004:MRC

[DLMS04] Nancy Durgin, Patrick Lincoln, John Mitchell, and Andre Scedrov. Multiset rewriting and the complexity of bounded security protocols. Journal of Computer Security, 12(2):247–311, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

DeCapitanidiVimercati:2001:GIP

[DLRS01] Sabrina De Capitani di Vimercati, Patrick Lincoln, Livio Ricciulli, and Pierangela Samarati. Global infrastructure protection system.

Journal of Computer Security, 9(4):251–283, ???? 2001. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Deng:2011:ZKB

[DLYZ11] Robert H. Deng, Yingjiu Li, Moti Yung, and Yunlei Zhao. A zeroknowledge based framework for RFID privacy. Journal of Computer Security, 19(6):1109–1146, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Deswarte:2000:GEP

[DM00] Yves Deswarte and Catherine Meadows. Guest editors’ preface. Journal of Computer Security, 8(2–3):87, ???? 2000. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

DiPietro:2004:KMH

[DMM04] Roberto Di Pietro, Luigi V. Mancini, and Alessandro Mei. Key management for high bandwidth secure multicast. Journal of Computer Security, 12(5):693–709, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

DiPietro:2010:HKS

[DMM10] Roberto Di Pietro, Luigi V. Mancini, and Alessandro Mei. Hierarchies of keys in secure multicast communications. Journal of Computer Security, 18(5):839–860, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Durgin:2003:CLP

[DMP03] Nancy Durgin, John Mitchell, and Dusko Pavlovic. A compositional logic for proving security properties of protocols. Journal of Computer Security, 11(4):677–721, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Dimitrakakis:2015:ELA

[DMV15] Christos Dimitrakakis, Aikaterini Mitrokotsa, and Serge Vaudenay. Expected loss analysis for authentication in constrained channels. Journal of Computer Security, 23(3):309–329, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Dainotti:2009:CAA

[DPV09] Alberto Dainotti, Antonio Pescapé, and Giorgio Ventre. A cascade architecture for DoS attacks detection based on the wavelet transform. Journal of Computer Security, 17(6):945–968, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

DSouza:2016:MCT

[DR16] Deepak D’Souza and K. R. Raghavendra. Model-checking tracebased information flow properties for infinite-state systems. Journal of Computer Security, 24(5):617–643, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Dong:2011:SSE

[DRD11] Changyu Dong, Giovanni Russello, and Naranker Dulay. Shared and searchable encrypted data for untrusted servers. Journal of Computer Security, 19(3):367–397, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Dewri:2011:EPV

[DRRW11] Rinku Dewri, Indrajit Ray, Indrakshi Ray, and Darrell Whitley. Exploring privacy versus data quality trade-offs in anonymization techniques using multi-objective optimization. Journal of Computer Security, 19(5):935–974, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

DeCapitanidiVimercati:1997:ASE

[DS97] Sabrina De Capitani di Vimercati and Pierangela Samarati. Authorization specification and enforcement in federated database systems. Journal of Computer Security, 5(2):155–188, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Daniels:1999:IHA

[DS99] Thomas E. Daniels and Eugene H. Spafford. Identification of host audit data to detect attacks on low-level IP vulnerabilities. Journal of Computer Security, 7(1):3–35, ???? 1999. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

dosSantos:2017:AFE

[dSRCP17] Daniel Ricardo dos Santos, Silvio Ranise, Luca Compagna, and Serena Elisa Ponta. Automatically finding execution scenarios to deploy security-sensitive workflows. Journal of Computer Security, 25(3):255–282, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Demurjian:1997:TDP

[DT97] S. A. Demurjian Sr. and T. C. Ting. Towards a definitive paradigm for security in object-oriented systems and applications. Journal of Computer Security, 5(4):341–382, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Davidson:2016:CSS

[DTG16] Michal Davidson, Tamir Tassa, and Ehud Gudes. Content sharing schemes in DRM systems with enhanced performance and privacy preservation. Journal of Computer Security, 24(6):667–688, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Ding:2009:LFM

[DTX09] Xuhua Ding, Gene Tsudik, and Shouhuai Xu. Leak-free mediated group signatures. Journal of Computer Security, 17(4):489–514, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Duggan:2004:TBC

[Dug04] Dominic Duggan. Type-based cryptographic operations. Journal of Computer Security, 12(3–4):485–550, ???? 2004. CODEN JC-SIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Elbirt:2008:AAI

[Elb08] A. J. Elbirt. Accelerated AES implementations via generalized instruction set extensions. Journal of Computer Security, 16(3): 265–288, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Epstein:1993:HAW

[EMO^{+}93] Jeremy Epstein, John McHugh, Hilarie Orman, Rita Pascale, Ann Marmor-Squires, Bonnie Danner, Charles R. Martin, Martha Branstad, Glenn Benson, and Doug Rothnie. A high assurance window system prototype. Journal of Computer Security, 2(2–3): 159–190, ???? 1993. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Eamani:2006:LBP

[ES06] Arun K. Eamani and A. Prasad Sistla. Language based policy analysis in a SPKI trust management system. Journal of Computer Security, 14(4):327–357, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Eckmann:2002:SAL

[EVK02] Steven T. Eckmann, Giovanni Vigna, and Richard A. Kemmerer. STATL: An attack language for state-based intrusion detection. Journal of Computer Security, 10(1–2):71–103, ???? 2002. CO-DEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Fuhry:2018:HPS

[FBB^{+}18] Benny Fuhry, Raad Bahmani, Ferdinand Brasser, Florian Hahn, Florian Kerschbaum, and Ahmad-Reza Sadeghi. HardIDX: Practical and secure index with SGX in a malicious environment. Journal of Computer Security, 26(5):677–706, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Foley:2011:MSP

[FF11] Simon N. Foley and William M. Fitzgerald. Management of security policy configuration using a Semantic Threat Graph approach. Journal of Computer Security, 19(3):567–605, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Focardi:1995:CSP

[FG95] Riccardo Focardi and Roberto Gorrieri. A classification of security properties for process algebras. Journal of Computer Security, 3 (1):5–33, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Fabrega:1999:SSP

[FHG99] F. Javier Thayer Fábrega, Jonathan C. Herzog, and Joshua D. Guttman. Strand spaces: proving security protocols correct. Journal of Computer Security, 7(2–3):191–230, ???? 1999. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Foley:1995:SSC

[FJ95] Simon N. Foley and Jeremy L. Jacob. Specifying security for computer supported collaborative working. Journal of Computer Security, 3(4):233–253, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Franklin:1998:AML

[FM98] Matthew K. Franklin and Dahlia Malkhi. Auditable metering with lightweight security. Journal of Computer Security, 6(4):237–255, ???? 1998. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Focardi:2005:GEP

[Foc05] Riccardo Focardi. Guest Editor’s preface. Journal of Computer Security, 13(3):345, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Focardi:2006:P

[Foc06] Riccardo Focardi. Preface. Journal of Computer Security, 14(2): 111, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Focardi:2010:E

[Foc10] Riccardo Focardi. Editorial. Journal of Computer Security, 18 (6):969, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Foley:1992:ASN

[Fol92] Simon N. Foley. Aggregation and separation as noninterference properties. Journal of Computer Security, 1(2):159–188, ???? 1992. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Foley:1998:GEP

[Fol98] Simon N. Foley. Guest editors’ preface. Journal of Computer Security, 6(1–2):1, ???? 1998. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Foley:1999:GEP

[Fol99] Simon N. Foley. Guest editor’s preface. Journal of Computer Security, 7(2–3):87, ???? 1999. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Focardi:2006:IFS

[FR06] Riccardo Focardi and Sabina Rossi. Information flow security in dynamic contexts. Journal of Computer Security, 14(1):65–110, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Frincke:2002:GEP

[Fri02] Deborah Frincke. Guest Editor’s preface. Journal of Computer Security, 10(1–2):1–3, ???? 2002. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Giustolisi:2018:ICM

[GBL18] Rosario Giustolisi, Giampaolo Bella, and Gabriele Lenzini. Invalid certificates in modern browsers: a socio-technical analysis. Journal of Computer Security, 26(4):509–541, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Geigel:2013:NNT

[Gei13] Arturo Geigel. Neural network Trojan. Journal of Computer Security, 21(2):191–232, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Gray:1998:PSC

[GEL98] James W. Gray III, Kin Fai Epsilon Ip, and King-Shan Lui. Provable security for cryptographic protocols — exact analysis and engineering applications. Journal of Computer Security, 6(1–2): 23–52, ???? 1998. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Gupta:1992:TTP

[GG92] Sarbari Gupta and Virgil D. Gligor. Towards a theory of penetration-resistant systems and its applications. Journal of Computer Security, 1(2):133–158, ???? 1992. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Grothoff:2009:TBS

[GGS^{+}09] Christian Grothoff, Krista Grothoff, Ryan Stutsman, Ludmila Alkhutova, and Mikhail Atallah. Translation-based steganography. Journal of Computer Security, 17(3):269–303, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Gentry:2013:FSB

[GHPS13] Craig Gentry, Shai Halevi, Chris Peikert, and Nigel P. Smart. Field switching in BGV-style homomorphic encryption. Journal of Computer Security, 21(5):663–684, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Guttman:2005:VIF

[GHRS05] Joshua D. Guttman, Amy L. Herzog, John D. Ramsdell, and Clement W. Skorupka. Verifying information flow goals in Security-Enhanced Linux. Journal of Computer Security, 13(1): 115–134, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Gordon:2003:ATS

[GJ03] Andrew D. Gordon and Alan Jeffrey. Authenticity by typing for security protocols. Journal of Computer Security, 11(4):451–519, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Gordon:2004:TEA

[GJ04] Andrew D. Gordon and Alan Jeffrey. Types and effects for asymmetric cryptographic protocols. Journal of Computer Security, 12 (3–4):435–483, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Goubault-Larrecq:2010:FMF

[GL10] Jean Goubault-Larrecq. Finite models for formal security proofs. Journal of Computer Security, 18(6):1247–1299, ???? 2010. CO-DEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Gligor:1993:IRA

[GLP93] Virgil D. Gligor, Shyh-Wei Luan, and Joseph N. Pato. On interrealm authentication in large distributed systems. Journal of Computer Security, 2(2–3):137–157, ???? 1993. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Giffin:2017:HPD

[GLS^{+}17] Daniel Giffin, Amit Levy, Deian Stefan, David Terei, David Mazières, John Mitchell, and Alejandro Russo. Hails: Protecting data privacy in untrusted web applications. Journal of Computer Security, 25(4–5):427–461, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Gordon:2011:IIS

[GLZ11] Lawrence A. Gordon, Martin P. Loeb, and Lei Zhou. The impact of information security breaches: Has there been a downward shift in costs? Journal of Computer Security, 19(1):33–56, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Giacobazzi:2010:ACU

[GM10] Roberto Giacobazzi and Isabella Mastroeni. Adjoining classified and unclassified information by abstract interpretation. Journal of Computer Security, 18(5):751–797, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Gritzalis:2011:SPM

[GMR^{+}11] Dimitris Gritzalis, Giannis Marias, Yacine Rebahi, Yannis Soupionis, and Sven Ehlert. SPIDER: A platform for managing SIP-based Spam over Internet Telephony (SPIT). Journal of Computer Security, 19(5):835–867, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Guanciale:2016:PSM

[GNDB16] Roberto Guanciale, Hamed Nemati, Mads Dam, and Christoph Baumann. Provably secure memory isolation for Linux on ARM. Journal of Computer Security, 24(6):793–837, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Gong:1995:GEP

[Gon95] Li Gong. Guest Editor’s preface. Journal of Computer Security, 3(1):3, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Gorrieri:2005:GEP

[Gor05] Roberto Gorrieri. Guest Editor’s preface. Journal of Computer Security, 13(1):1–2, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Goto:2006:P

[Got06] Atsuhiro Goto. Preface. Journal of Computer Security, 14(3):247, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Goto:2007:P

[Got07] Atsuhiro Goto. Preface. Journal of Computer Security, 15(5):491, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Goto:2010:P

[Got10] Atsuhiro Goto. Preface. Journal of Computer Security, 18(4):497, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Gudes:1999:MSI

[GOvdR99] Ehud Gudes, Martin S. Olivier, and Reind P. van de Riet. Modelling, specifying and implementing workflow security in Cyberspace. Journal of Computer Security, 7(4):287–315, ???? 1999. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Gorla:2010:P

[GP10] Daniele Gorla and Catuscia Palamidessi. Preface. Journal of Computer Security, 18(2):189, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Garg:2012:SAL

[GP12] Deepak Garg and Frank Pfenning. Stateful authorization logic — proof theory and a case study. Journal of Computer Security, 20(4):353–391, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Gray:1992:TMF

[Gra92] James W. Gray III. Toward a mathematical foundation for information flow security. Journal of Computer Security, 1(3–4): 255–294, ???? 1992. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Gritzalis:2011:GEP

[Gri11] Dimitris Gritzalis. Guest Editor’s preface. Journal of Computer Security, 19(6):1027–1028, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Gibson-Robinson:2015:VLS

[GRKL15] Thomas Gibson-Robinson, Allaa Kamil, and Gavin Lowe. Verifying layered security protocols. Journal of Computer Security, 23(3):259–307, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Gupta:2018:SCA

[GSAV18] Haritabh Gupta, Shamik Sural, Vijayalakshmi Atluri, and Jaideep Vaidya. A side-channel attack on smartphones: Deciphering key taps using built-in microphones. Journal of Computer Security, 26(2):255–281, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Guan:2017:SKE

[GT17] Albert Guan and Wen-Guey Tzeng. A secret key establishment protocol for wireless networks using noisy channels. Journal of Computer Security, 25(2):139–151, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Goodrich:2008:NFI

[GTY08] Michael T. Goodrich, Roberto Tamassia, and Danfeng (Daphne) Yao. Notarized federated ID management and authentication. Journal of Computer Security, 16(4):399–418, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Guttman:2004:FAP

[GTZ04] Joshua D. Guttman, F. Javier Thayer, and Lenore D. Zuck. The faithfulness of abstract protocol analysis: Message authentication. Journal of Computer Security, 12(6):865–891, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Guette:2009:ATK

[Gue09] Gilles Guette. Automating trusted key rollover in DNSSEC. Journal of Computer Security, 17(6):839–854, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Guttman:2004:GEP

[Gut04a] Joshua Guttman. Guest Editor’s preface. Journal of Computer Security, 12(1):1, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Guttman:2004:ATD

[Gut04b] Joshua D. Guttman. Authentication tests and disjoint encryption: A design method for security protocols. Journal of Computer Security, 12(3–4):409–433, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Guttman:2009:I

[Gut09] Joshua D. Guttman. Introduction. Journal of Computer Security, 17(5):515, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Guttman:2014:EPP

[Gut14] Joshua D. Guttman. Establishing and preserving protocol security goals. Journal of Computer Security, 22(2):203–267, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Halpern:2017:QQP

[Hal17] Joseph Y. Halpern. From qualitative to quantitative proofs of security properties using first-order conditional logic. Journal of Computer Security, 25(1):1–19, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Hancke:2011:PES

[Han11] Gerhard P. Hancke. Practical eavesdropping and skimming attacks on high-frequency RFID tokens. Journal of Computer Security, 19(2):259–288, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Hirschi:2019:MUV

[HBD19] Lucca Hirschi, David Baelde, and Stéphanie Delaune. A method for unbounded verification of privacy-type properties. Journal of Computer Security, 27(3):277–342, ???? 2019. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Hedin:2016:IFS

[HBS16] Daniel Hedin, Luciano Bello, and Andrei Sabelfeld. Informationflow security for JavaScript and its APIs. Journal of Computer Security, 24(2):181–234, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Horton:1993:CVP

[HCH^{+}93] J. D. Horton, R. H. Cooper, W. F. Hyslop, B. G. Nickerson, O. K. Ward, Robert Harland, Elton Ashby, and W. M. Stewart. The cascade vulnerability problem. Journal of Computer Security, 2(4): 279–290, ???? 1993. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Hsu:2011:WLC

[HCM11] Francis Hsu, Hao Chen, and Sridhar Machiraju. WebCallerID: Leveraging cellular networks for Web authentication. Journal of Computer Security, 19(5):869–893, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Hinke:1995:FAD

[HDC95] Thomas H. Hinke, Harry S. Delugach, and Asha Chandrasekhar. A fast algorithm for detecting second paths in database inference analysis. Journal of Computer Security, 3(2–3):147–168, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Herzberg:2009:FPT

[Her09] Amir Herzberg. Folklore, practice and theory of robust combiners. Journal of Computer Security, 17(2):159–189, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Hofmeyr:1998:IDU

[HFS98] Steven A. Hofmeyr, Stephanie Forrest, and Anil Somayaji. Intrusion detection using sequences of system calls. Journal of Computer Security, 6(3):151–180, ???? 1998. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Harrison:2009:AIF

[HH09] William L. Harrison and James Hook. Achieving information flow security through monadic control of effects. Journal of Computer Security, 17(5):599–653, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Hauser:1996:RSP

[HJT^{+}96] Ralf Hauser, Philippe Janson, Gene Tsudik, Els Van Herreweghen, and Refik Molva. Robust and secure password and key change method. Journal of Computer Security, 4(1):97–111, ???? 1996. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Hallgren:2018:AB

[HKOS18] Per Hallgren, Ravi Kishore, Martín Ochoa, and Andrei Sabelfeld. Assuring BetterTimes. Journal of Computer Security, 26(4):557– 587, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Hui:2001:FPS

[HL01] Mei Lin Hui and Gavin Lowe. Fault-preserving simplifying transformations for security protocols. Journal of Computer Security, 9 (1–2):3–46, ???? 2001. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Heather:2003:HPT

[HLS03] James Heather, Gavin Lowe, and Steve Schneider. How to prevent type flaw attacks on security protocols. Journal of Computer Security, 11(2):217–244, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

He:2011:SCP

[HLVA11] Xiaoyun He, Haibing Lu, Jaideep Vaidya, and Nabil Adam. Secure construction and publication of contingency tables from distributed data. Journal of Computer Security, 19(3):453–484, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Hamadou:2010:CPS

[HM10] Sardaouna Hamadou and John Mullins. Calibrating the power of schedulers for probabilistic polynomial-time calculus. Journal of Computer Security, 18(2):265–316, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Herzberg:2013:FJL

[HM13] Amir Herzberg and Ronen Margulies. Forcing Johnny to login safely. Journal of Computer Security, 21(3):393–424, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Hofheinz:2009:PRS

[HMQU09] Dennis Hofheinz, Jörn Müller-Quade, and Dominique Unruh. Polynomial runtime in simulatability definitions. Journal of Computer Security, 17(5):703–735, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Heiderich:2014:SAS

[HNS^{+}14] M. Heiderich, M. Niemietz, F. Schuster, T. Holz, and J. Schwenk. Scriptless attacks: Stealing more pie without touching the sill. Journal of Computer Security, 22(4):567–599, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Halpern:2005:AIH

[HO05] Joseph Y. Halpern and Kevin R. O’Neill. Anonymity and information hiding in multiagent systems. Journal of Computer Security, 13(3):483–514, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Horn:2000:APF

[HP00] Günther Horn and Bart Preneel. Authentication and payment in future mobile systems. Journal of Computer Security, 8(2–3): 183–207, ???? 2000. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Hale:2003:PAC

[HPS03] John Hale, Mauricio Papa, and Sujeet Shenoi. Programmable access control. Journal of Computer Security, 11(3):331–351, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Hughes:2004:IHA

[HS04] Dominic Hughes and Vitaly Shmatikov. Information hiding, anonymity and privacy: a modular approach. Journal of Computer Security, 12(1):3–36, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Heather:2005:DPE

[HS05] James Heather and Steve Schneider. A decision procedure for the existence of a rank function. Journal of Computer Security, 13(2): 317–344, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Halevi:2011:TBH

[HSH11] Tzipora Halevi, Nitesh Saxena, and Shai Halevi. Tree-based HB protocols for privacy-preserving authentication of RFID tags. Journal of Computer Security, 19(2):343–363, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Hagen:2008:PAU

[HSR08] Janne Merete Hagen, Tormod Kalberg Sivertsen, and Chunming Rong. Protection against unauthorized access and computer crime in Norwegian enterprises. Journal of Computer Security, 16(3): 341–366, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Hu:2009:CRR

[HTML09] Nan Hu, Giri K. Tayi, Chengyu Ma, and Yingjiu Li. Certificate revocation release policies. Journal of Computer Security, 17(2): 127–157, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Hale:2000:TBA

[HTS00] John Hale, Jody Threet, and Sujeet Shenoi. A ticket-based access control architecture for object systems. Journal of Computer Security, 8(1):43–65, ???? 2000. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Hu:1992:RTC

[Hu92] Wei-Ming Hu. Reducing timing channels with fuzzy time. Journal of Computer Security, 1(3–4):233–254, ???? 1992. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Halpern:2001:LSL

[HvdM01] Joseph Y. Halpern and Ron van der Meyden. A logic for SDSI’s linked local name spaces. Journal of Computer Security, 9(1–2): 105–142, ???? 2001. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Halpern:2003:LRS

[HvdM03] Joseph Y. Halpern and Ron van der Meyden. A logical reconstruction of SPKI. Journal of Computer Security, 11(4):581–613, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Hong:2012:SED

[HVL12] Yuan Hong, Jaideep Vaidya, and Haibing Lu. Secure and efficient distributed linear programming. Journal of Computer Security, 20(5):583–634, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Han:2018:RAM

[HZL^{+}18] Jin Han, Wangyu Zang, Li Liu, Songqing Chen, and Meng Yu. Risk-aware multi-objective optimized virtual machine placement in the cloud. Journal of Computer Security, 26(5):707–730, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Homer:2013:AVM

[HZO^{+}13] John Homer, Su Zhang, Xinming Ou, David Schmidt, Yanhui Du, S. Raj Rajagopalan, and Anoop Singhal. Aggregating vulnerability metrics in enterprise networks using attack graphs. Journal of Computer Security, 21(4):561–597, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Iheagwara:2003:CEE

[IBS03] Charles Iheagwara, Andrew Blyth, and Mukesh Singhal. A comparative experimental evaluation study of intrusion detection system performance in a gigabit environment. Journal of Computer Security, 11(1):1–33, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Iheagwara:2004:CEM

[IBS04] Charles Iheagwara, Andrew Blyth, and Mukesh Singhal. Cost effective management frameworks for intrusion detection systems. Journal of Computer Security, 12(5):777–798, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Jacob:1992:BTA

[Jac92] Jeremy Jacob. Basic theorems about security. Journal of Computer Security, 1(3–4):385–411, ???? 1992. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Jing:2016:TML

[JAH^{+}16] Yiming Jing, Gail-Joon Ahn, Hongxin Hu, Haehyun Cho, and Ziming Zhao. TripleMon: A multi-layer security framework for mediating inter-process communication on Android. Journal of Computer Security, 24(4):405–426, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Jajodia:2001:MST

[JAK^{+}01] Sushil Jajodia, Vijayalakshmi Atluri, Thomas F. Keefe, Catherine D. McCollum, and Ravi Mukkamala. Multilevel secure transaction processing. Journal of Computer Security, 9(3):165–195, ???? 2001. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Jadliwala:2013:OMZ

[JBH13] Murtuza Jadliwala, Igor Bilogrevic, and Jean-Pierre Hubaux. Optimizing mix-zone coverage in pervasive wireless networks. Journal of Computer Security, 21(3):317–346, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Jajodia:2002:GEP

[JG02] Sushil Jajodia and Dimitris Gritzalis. Guest Editors’ preface. Journal of Computer Security, 10(3):211, ???? 2002. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Jacobs:2009:SLS

[JH09] Bart Jacobs and Ichiro Hasuo. Semantics and logic for security protocols. Journal of Computer Security, 17(6):909–944, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Jiwa:1996:BAD

[JHS96] Azad Jiwa, Thomas Hardjono, and Jennifer Seberry. Beacons for authentication in distributed systems. Journal of Computer Security, 4(1):81–96, ???? 1996. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Jovanovic:2010:SAD

[JKK10] Nenad Jovanovic, Christopher Kruegel, and Engin Kirda. Static analysis for detecting taint-style vulnerabilities in web applications. Journal of Computer Security, 18(5):861–907, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Jajodia:1992:EPa

[JM92a] Sushil Jajodia and Jonathan Millen. Editors’ preface. Journal of Computer Security, 1(1):1–3, ???? 1992. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Jajodia:1992:EPb

[JM92b] Sushil Jajodia and Jonathan Millen. Editors’ preface. Journal of Computer Security, 1(3–4):215, ???? 1992. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Jajodia:1993:EPa

[JM93] Sushil Jajodia and Jonathan Millen. Editors’ preface. Journal of Computer Security, 2(2–3):85, ???? 1993. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Jajodia:1995:EPa

[JM95a] Sushil Jajodia and Jonathan Millen. Editors’ preface. Journal of Computer Security, 3(1):1, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Jajodia:1995:EPc

[JM95b] Sushil Jajodia and Jonathan Millen. Editors’ preface. Journal of Computer Security, 3(4):231, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Jajodia:1996:EPa

[JM96a] Sushil Jajodia and Jonathan Millen. Editors’ preface. Journal of Computer Security, 4(2–3):121, ???? 1996. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Jajodia:1996:EPb

[JM96b] Sushil Jajodia and Jonathan Millen. Editors’ preface. Journal of Computer Security, 4(4):265, ???? 1996. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Jajodia:1997:EP

[JM97] Sushil Jajodia and Jonathan Millen. Editors’ preface. Journal of Computer Security, 5(2):109, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Jajodia:2010:E

[JM10] Sushil Jajodia and Jon Millen. Editorial. Journal of Computer Security, 18(2):187, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Jajodia:2016:UTP

[JPSS16] Sushil Jajodia, Noseong Park, Edoardo Serra, and V. S. Subrahmanian. Using temporal probabilistic logic for optimal monitoring of security events with limited resources. Journal of Computer Security, 24(6):735–791, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Jha:2004:MCS

[JR04] S. Jha and T. Reps. Model checking SPKI/SDSI. Journal of Computer Security, 12(3–4):317–353, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Kayem:2008:RCK

[KAM08] Anne V. D. M. Kayem, Selim G. Akl, and Patrick Martin. On replacing cryptographic keys in hierarchical key management systems. Journal of Computer Security, 16(3):289–309, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Karjoth:2000:ACS

[Kar00] Günter Karjoth. Authorization in CORBA security. Journal of Computer Security, 8(2–3):89–108, ???? 2000. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Kopf:2011:ADI

[KB11] Boris Köpf and David Basin. Automatically deriving informationtheoretic bounds for adaptive side-channel attacks. Journal of Computer Security, 19(1):1–31, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Kakkar:2003:RAS

[KGA03] Pankaj Kakkar, Carl A. Gunter, and Martín Abadi. Reasoning about secrecy for active networks. Journal of Computer Security, 11(2):245–287, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Kang:1995:TMM

[KK95] I. E. Kang and T. F. Keefe. Transaction management for multilevel secure replicated databases. Journal of Computer Security, 3(2–3): 115–145, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Kremer:2016:AAS

[KK16] Steve Kremer and Robert Künnemann. Automated analysis of security protocols with global state. Journal of Computer Security, 24(5):583–616, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Kanovich:2017:TCC

[KKN^{+}17] Max Kanovich, Tajana Ban Kirigin, Vivek Nigam, Andre Scedrov, and Carolyn Talcott. Time, computational complexity, and probability in the analysis of distance-bounding protocols. Journal of Computer Security, 25(6):585–630, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Kamil:2011:ATS

[KL11] Allaa Kamil and Gavin Lowe. Analysing TLS in the strand spaces model. Journal of Computer Security, 19(5):975–1025, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Knudsen:1998:SMD

[KM98] Lars R. Knudsen and Keith M. Martin. In search of multiple domain key recovery. Journal of Computer Security, 6(4):219–235, ???? 1998. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Kremer:2010:CSA

[KM10] Steve Kremer and Laurent Mazaré. Computationally sound analysis of protocols using bilinear pairings. Journal of Computer Security, 18(6):999–1033, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Krukow:2008:LFH

[KNS08] Karl Krukow, Mogens Nielsen, and Vladimiro Sassone. A logical framework for history-based access control and reputation systems. Journal of Computer Security, 16(1):63–101, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Kobsa:2013:CJV

[KNTU13] Alfred Kobsa, Rishab Nithyanand, Gene Tsudik, and Ersin Uzun. Can Jannie verify? Usability of display-equipped RFID tags for security purposes. Journal of Computer Security, 21(3):347–370, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Kar:2016:SSI

[KPS16] Debabrata Kar, Suvasini Panigrahi, and Srikanth Sundararajan. SQLiDDS: SQL injection detection using document similarity measure. Journal of Computer Security, 24(4):507–539, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Kremer:2003:GBV

[KR03] Steve Kremer and Jean-François Raskin. A game-based verification of non-repudiation and fair exchange protocols. Journal of Computer Security, 11(3):399–429, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Kreitz:2013:FSW

[Kre13] Gunnar Kreitz. Flow stealing: A well-timed redirection attack. Journal of Computer Security, 21(3):371–391, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Kolesnikov:2013:SAP

[KSS13] Vladimir Kolesnikov, Ahmad-Reza Sadeghi, and Thomas Schneider. A systematic approach to practically efficient general twoparty secure function evaluation protocols and their modular design. Journal of Computer Security, 21(2):283–315, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Kelsey:2000:SCC

[KSWH00] John Kelsey, Bruce Schneier, David Wagner, and Chris Hall. Side channel cryptanalysis of product ciphers. Journal of Computer Security, 8(2–3):141–158, ???? 2000. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Kerschbaum:2002:UIS

[KSZ02] Florian Kerschbaum, Eugene H. Spafford, and Diego Zamboni. Using internal sensors and embedded detectors for intrusion detection. Journal of Computer Security, 10(1–2):23–70, ???? 2002. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Kusters:2012:GBD

[KTV12] Ralf Küsters, Tomasz Truderung, and Andreas Vogt. A gamebased definition of coercion resistance and its applications. Journal of Computer Security, 20(6):709–764, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Liu:2017:FBO

[LAGM17] Jed Liu, Owen Arden, Michael D. George, and Andrew C. Myers. Fabric: Building open distributed systems securely by construction. Journal of Computer Security, 25(4–5):367–426, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

LaPadula:1996:F

[LaP96] Leonard J. LaPadula. Foreword. Journal of Computer Security, 4(2–3):233–238, ???? 1996. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

LaPadula:1996:MTR

[LB96] Leonard J. LaPadula and D. Elliott Bell. MITRE technical report 2547, volume II. Journal of Computer Security, 4(2–3):239–263, ???? 1996. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Littlewood:1993:TOM

[LBF^{+}93] Bev Littlewood, Sarah Brocklehurst, Norman Fenton, Peter Mellor, Stella Page, David Wright, John Dobson, John McDermid, and Dieter Gollmann. Towards operational measures of computer security. Journal of Computer Security, 2(2–3):211–229, ???? 1993. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Li:2015:NAC

[LCL^{+}15] Jin Li, Xiaofeng Chen, Jingwei Li, Chunfu Jia, Jianfeng Ma, and Wenjing Lou. New access control systems based on outsourced attribute-based encryption. Journal of Computer Security, 23(6): 659–683, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Lee:2002:TCS

[LFM^{+}02] Wenke Lee, Wei Fan, Matthew Miller, Salvatore J. Stolfo, and Erez Zadok. Toward cost-sensitive modeling for intrusion detection and response. Journal of Computer Security, 10(1–2):5–22, ???? 2002. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Lekkas:2010:PMT

[LG10] Dimitrios Lekkas and Dimitris Gritzalis. e-Passports as a means towards a Globally Interoperable Public Key Infrastructure. Journal of Computer Security, 18(3):379–396, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Levy:2018:SHB

[LGGO18] Silvie Levy, Ehud Gudes, and Nurit Gal-Oz. Sharing-habits based privacy control in social networks. Journal of Computer Security, 26(2):177–205, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Li:2010:CCB

[LHM^{+}10] Jiguo Li, Xinyi Huang, Yi Mu, Willy Susilo, and Qianhong Wu. Constructions of certificate-based signature secure against key replacement attacks. Journal of Computer Security, 18(3):421–449, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Lu:2015:TUO

[LHY^{+}15] Haibing Lu, Yuan Hong, Yanjiang Yang, Lian Duan, and Nazia Badar. Towards user-oriented RBAC model. Journal of Computer Security, 23(1):107–129, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Li:2012:GEP

[Li12] Yingjiu Li. Guest Editor’s preface. Journal of Computer Security, 20(5):461–462, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Lin:1999:GEP

[Lin99] T. Y. Lin. Guest editor’s preface. Journal of Computer Security, 7(4):255, ???? 1999. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Lin:2000:GEP

[Lin00] T. Y. Lin. Guest editor’s preface. Journal of Computer Security, 8(1):1, ???? 2000. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Liu:2000:ICI

[LJM00] Peng Liu, Sushil Jajodia, and Catherine D. McCollum. Intrusion confinement by isolation in information systems. Journal of Computer Security, 8(4):243–279, ???? 2000. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Le:2016:MCR

[LKAJ16] Meixing Le, Krishna Kant, Malek Athamnah, and Sushil Jajodia. Minimum cost rule enforcement for cooperative database access. Journal of Computer Security, 24(3):379–403, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Lippert:2006:LCM

[LKWB06] M. Lippert, V. Karatsiolis, A. Wiesmaier, and J. Buchmann. Lifecycle management of X.509 certificates based on LDAP directories. Journal of Computer Security, 14(5):419–439, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Lowden:2015:DPA

[LŁA15] Jason Lowden, Marcin Łukowiak, and Sonia Lopez Alarcon. Design and performance analysis of efficient Keccak tree hashing on GPU architectures. Journal of Computer Security, 23(5):541–562, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Lunt:1992:GEP

[LM92] Teresa F. Lunt and John McLean. Guest Editors’ preface. Journal of Computer Security, 1(3–4):217–218, ???? 1992. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Luckett:2018:ISM

[LMG^{+}18] Patrick Luckett, J. Todd McDonald, William B. Glisson, Ryan Benton, Joel Dawson, and Blair A. Doyle. Identifying stealth malware using CPU power consumption and learning algorithms. Journal of Computer Security, 26(5):589–613, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Lopez:2006:UPK

[Lop06] Javier Lopez. Unleashing public-key cryptography in wireless sensor networks. Journal of Computer Security, 14(5):469–482, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Lester:2016:IFA

[LOS16] Martin Lester, Luke Ong, and Max Schäfer. Information flow analysis for a dynamically typed language with staged metaprogramming. Journal of Computer Security, 24(5):541–582, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Lotz:1997:TSM

[Lot97] Volkmar Lotz. Threat scenarios as a means to formally develop secure systems. Journal of Computer Security, 5(1):31–67, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Lowe:1998:CCA

[Low98] Gavin Lowe. Casper: A compiler for the analysis of security protocols. Journal of Computer Security, 6(1–2):53–84, ???? 1998. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Lowe:1999:TCR

[Low99] Gavin Lowe. Towards a completeness result for model checking of security protocols. Journal of Computer Security, 7(2–3):89–146, ???? 1999. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Lowe:2004:APS

[Low04a] Gavin Lowe. Analysing protocols subject to guessing attacks. Journal of Computer Security, 12(1):83–97, ???? 2004. CODEN JC-SIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Lowe:2004:DIF

[Low04b] Gavin Lowe. Defining information flow quantity. Journal of Computer Security, 12(3–4):619–653, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Levine:2002:HMB

[LS02] Brian Neil Levine and Clay Shields. Hordes: a multicast based protocol for anonymity. Journal of Computer Security, 10(3):213– 240, ???? 2002. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Luo:2016:MCM

[LSMR16] Zhengqin Luo, José Fragoso Santos, Ana Almeida Matos, and Tamara Rezk. Mashic compiler: Mashup sandboxing based on inter-frame communication. Journal of Computer Security, 24(1): 91–136, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Lin:2017:SAC

[LT17] Chung-Yi Lin and Wen-Guey Tzeng. Strategy analysis for cloud storage reliability management based on game theory. Journal of Computer Security, 25(2):153–171, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Lorenzi:2018:TDR

[LUV^{+}18] David Lorenzi, Emre Uzun, Jaideep Vaidya, Shamik Sural, and Vijayalakshmi Atluri. Towards designing robust CAPTCHAs. Journal of Computer Security, 26(6):731–760, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Lu:2014:OFR

[LVA14] Haibing Lu, Jaideep Vaidya, and Vijayalakshmi Atluri. An optimization framework for role mining. Journal of Computer Security, 22(1):1–31, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Li:2003:DCC

[LWM03] Ninghui Li, William H. Winsborough, and John C. Mitchell. Distributed credential chain discovery in trust management. Journal of Computer Security, 11(1):35–86, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Li:2002:EPA

[LWWJ02] Yingjiu Li, Ningning Wu, X. Sean Wang, and Sushil Jajodia. Enhancing profiles for anomaly detection using time granularities. Journal of Computer Security, 10(1–2):137–157, ???? 2002. CO-DEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Liu:2015:JSD

[LWZZ15] Wen Ming Liu, Lingyu Wang, Lei Zhang, and Shunzhi Zhu. kjump: A strategy to design publicly-known algorithms for privacy preserving micro-data disclosure. Journal of Computer Security, 23(2):131–165, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Li:2010:ESS

[LYW^{+}10] Chung Ki Li, Guomin Yang, Duncan S. Wong, Xiaotie Deng, and Sherman S. M. Chow. An efficient signcryption scheme with key privacy and its extension to ring signcryption. Journal of Computer Security, 18(3):451–473, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Li:2011:GEP

[LZ11] Yingjiu Li and Jianying Zhou. Guest Editors’ preface. Journal of Computer Security, 19(2):227–228, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Lipmaa:2013:MEC

[LZ13] Helger Lipmaa and Bingsheng Zhang. A more efficient computationally sound non-interactive zero-knowledge shuffle argument. Journal of Computer Security, 21(5):685–719, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Malacaria:2010:RAS

[Mal10] Pasquale Malacaria. Risk assessment of security threats for looping constructs. Journal of Computer Security, 18(2):191–228, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Matsumoto:1998:HCC

[Mat98] Tsutomu Matsumoto. Human–computer cryptography: An attempt. Journal of Computer Security, 6(3):129–149, ???? 1998. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Matos:2009:DND

[MB09] Ana Almeida Matos and Gérard Boudol. On declassification and the non-disclosure policy. Journal of Computer Security, 17(5): 549–597, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Myers:2012:P

[MB12] Andrew Myers and Michael Backes. Preface. Journal of Computer Security, 20(6):635–636, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Mallios:2015:PCE

[MBK^{+}15] Yannis Mallios, Lujo Bauer, Dilsun Kaynar, Fabio Martinelli, and Charles Morisset. Probabilistic cost enforcement of security policies. Journal of Computer Security, 23(6):759–787, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Meier:2013:ECM

[MCB13] Simon Meier, Cas Cremers, and David Basin. Efficient construction of machine-checked symbolic protocol security proofs. Journal of Computer Security, 21(1):41–87, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

McLean:1992:PNF

[McL92] John McLean. Proving noninterference and functional correctness using traces. Journal of Computer Security, 1(1):37–57, ???? 1992. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Ma:2019:TOP

[MCL^{+}19] Ziqiang Ma, Quanwei Cai, Jingqiang Lin, Bo Luo, and Jiwu Jing. Towards the optimal performance of integrating Warm and Delay against remote cache timing side channels on block ciphers. Journal of Computer Security, 27(5):547–580, ???? 2019. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Meadows:1992:AFM

[Mea92] Catherine Meadows. Applying formal methods to the analysis of a key management protocol. Journal of Computer Security, 1(1): 5–35, ???? 1992. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Meadows:2001:CBF

[Mea01] Catherine Meadows. A cost-based framework for analysis of denial of service in networks. Journal of Computer Security, 9(1–2):143– 164, ???? 2001. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Merritt:1997:GEP

[Mer97] Michael Merritt. Guest editor’s preface. Journal of Computer Security, 5(2):111–112, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Murray:2008:NDA

[MG08] Toby Murray and Duncan Grove. Non-delegatable authorities in capability systems. Journal of Computer Security, 16(6):743–759, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Migault:2017:DEI

[MGK^{+}17] Daniel Migault, Tobias Guggemos, Sylvain Killian, Maryline Laurent, Guy Pujolle, and Jean Philippe Wary. Diet-ESP: IP layer security for IoT. Journal of Computer Security, 25(2):173–203, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Mohamed:2017:SUD

[MGS^{+}17] Manar Mohamed, Song Gao, Niharika Sachdeva, Nitesh Saxena, Chengcui Zhang, Ponnurangam Kumaraguru, and Paul C. Van Oorschot. On the security and usability of dynamic cognitive game CAPTCHAs. Journal of Computer Security, 25(3):205–230, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Millen:1993:EPb

[Mil93a] Jonathan Millen. Editor’s preface. Journal of Computer Security, 2(4):277, ???? 1993. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Millen:1993:RAM

[Mil93b] Jonathan K. Millen. A resource allocation model for denial of service protection. Journal of Computer Security, 2(2–3):89–106, ???? 1993. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Millen:1995:EPb

[Mil95a] Jonathan K. Millen. Editor’s preface. Journal of Computer Security, 3(2–3):85, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Millen:1995:UFC

[Mil95b] Jonathan K. Millen. Unwinding forward correctability. Journal of Computer Security, 3(1):35–54, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Millen:1996:EPB

[Mil96] Jonathan Millen. Editor’s preface to the Bell–LaPadula model. Journal of Computer Security, 4(2–3):229–231, ???? 1996. CO-DEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Millen:1999:SSI

[Mil99] Jonathan Millen. Special section on intrusion detection. Journal of Computer Security, 7(1):1, ???? 1999. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

McLean:1993:GEP

[MK93] John McLean and Richard Kemmerer. Guest Editors’ preface. Journal of Computer Security, 2(2–3):87–88, ???? 1993. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

McDermott:1996:APC

[MM96] John McDermott and Ravi Mukkamala. Analytic performance comparison of transaction processing algorithms for the SINTRA replicated-architecture database system. Journal of Computer Security, 4(2–3):189–228, ???? 1996. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Merlo:2015:MEP

[MMF15] Alessio Merlo, Mauro Migliardi, and Paolo Fontanelli. Measuring and estimating power consumption in Android to support energybased intrusion detection. Journal of Computer Security, 23(5): 611–637, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Mardziel:2013:DEK

[MMHS13] Piotr Mardziel, Stephen Magill, Michael Hicks, and Mudhakar Srivatsa. Dynamic enforcement of knowledge-based security policies using probabilistic abstract interpretation. Journal of Computer Security, 21(4):463–532, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Maffei:2019:GOP

[MMRS19] Matteo Maffei, Giulio Malavolta, Manuel Reinert, and Dominique Schröder. Group ORAM for privacy and access control in outsourced personal records. Journal of Computer Security, 27(1): 1–47, ???? 2019. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Molsa:2005:MDS

[Möl05] Jarmo Mölsä. Mitigating denial of service attacks: A tutorial. Journal of Computer Security, 13(6):807–837, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Motro:1992:UMS

[Mot92] Amihai Motro. A unified model for security and integrity in relational databases. Journal of Computer Security, 1(2):189–213, ???? 1992. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

McDaniel:2006:EPA

[MP06] Patrick McDaniel and Atul Prakash. Enforcing provisioning and authorization policy in the Antigone system. Journal of Computer Security, 14(6):483–511, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Malkhi:1997:HTS

[MR97] Dahlia Malkhi and Michael Reiter. A high-throughput secure reliable multicast protocol. Journal of Computer Security, 5(2):113– 127, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Moataz:2018:SSE

[MRR^{+}18] Tarik Moataz, Indrajit Ray, Indrakshi Ray, Abdullatif Shikfa, Frédéric Cuppens, and Nora Cuppens. Substring search over encrypted data. Journal of Computer Security, 26(1):1–30, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Maurer:1996:CSB

[MS96] Ueli M. Maurer and Pierre E. Schmid. A calculus for security boots trapping in distributed systems. Journal of Computer Security, 4 (1):55–80, ???? 1996. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Mantel:2003:UAS

[MS03] Heiko Mantel and Andrei Sabelfeld. A unifying approach to the security of distributed and multi-threaded programs. Journal of Computer Security, 11(4):615–676, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Millen:2005:SPAa

[MS05a] Jonathan Millen and Vitaly Shmatikov. Symbolic protocol analysis with an Abelian group operator or Diffie–Hellman exponentiation. Journal of Computer Security, 13(3):515–564, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Millen:2005:SPAb

[MS05b] Jonathan Millen and Vitaly Shmatikov. Symbolic protocol analysis with an Abelian group operator or Diffie–Hellman exponentiation. Journal of Computer Security, 13(4):695, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Myers:2013:BBC

[MSas13] Steven Myers, Mona Sergi, and abhi shelat. Black-box construction of a more than non-malleable CCA1 encryption scheme from plaintext awareness. Journal of Computer Security, 21(5):721–748, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Mitra:2015:GTR

[MSAV15] Barsha Mitra, Shamik Sural, Vijayalakshmi Atluri, and Jaideep Vaidya. The generalized temporal role mining problem. Journal of Computer Security, 23(1):31–58, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Murray:2017:SIV

[MSB17] Toby Murray, Andrei Sabelfeld, and Lujo Bauer. Special issue on verified information flow security. Journal of Computer Security, 25(4–5):319–321, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Meadows:2004:FSA

[MSC04] Catherine Meadows, Paul Syverson, and Iliano Cervesato. Formal specification and analysis of the Group Domain of Interpretation Protocol using NPATRL and the NRL Protocol Analyzer. Journal of Computer Security, 12(6):893–931, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Myers:2006:ERD

[MSZ06] Andrew C. Myers, Andrei Sabelfeld, and Steve Zdancewic. Enforcing robust declassification and qualified robustness. Journal of Computer Security, 14(2):157–196, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Mont:2008:PPE

[MT08] Marco Casassa Mont and Robert Thyne. Privacy policy enforcement in enterprises with identity management solutions. Journal of Computer Security, 16(2):133–163, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Majumdar:2019:LPD

[MTJ^{+}19] Suryadipta Majumdar, Azadeh Tabiban, Yosr Jarraya, Momen Oqaily, Amir Alimohammadifar, Makan Pourzandi, Lingyu Wang, and Mourad Debbabi. Learning probabilistic dependencies among events for proactive security auditing in clouds. Journal of Computer Security, 27(2):165–202, ???? 2019. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Modersheim:2010:CDS

[MVB10] Sebastian Mödersheim, Luca Viganò, and David Basin. Constraint differentiation: Search-space reduction for the constraint-based analysis of security protocols. Journal of Computer Security, 18(4): 575–618, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Mannan:2011:LPD

[MvO11] Mohammad Mannan and P. C. van Oorschot. Leveraging personal devices for stronger password authentication from untrusted computers. Journal of Computer Security, 19(4):703–750, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Micciancio:2004:CTA

[MW04] Daniele Micciancio and Bogdan Warinschi. Completeness theorems for the Abadi–Rogaway language of encrypted expressions. Journal of Computer Security, 12(1):99–129, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Notargiacomo:1995:MMI

[NBM95] LouAnna Notargiacomo, Barbara T. Blaustein, and Catherine D. McCollum. Merging models: Integrity, dynamic separation of duty and trusted data management. Journal of Computer Security, 3 (2–3):207–230, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Neville:2018:RAF

[NF18] Ultan Neville and Simon N. Foley. Reasoning about firewall policies through refinement and composition. Journal of Computer Security, 26(2):207–254, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Nayyar:2008:AAM

[NG08] Harshit Nayyar and Ali A. Ghorbani. Approximate autoregressive modeling for network attack detection. Journal of Computer Security, 16(2):165–197, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Nieto:2013:PVC

[NMP^{+}13] Juan González Nieto, Mark Manulis, Bertram Poettering, Jothi Rangasamy, and Douglas Stebila. Publicly verifiable ciphertexts. Journal of Computer Security, 21(5):749–778, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Nguyen:2011:APB

[NR11] L. H. Nguyen and A. W. Roscoe. Authentication protocols based on low-bandwidth unspoofable channels: A comparative survey. Journal of Computer Security, 19(1):139–201, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Niu:2014:FVS

[NRW14] Jianwei Niu, Mark Reith, and William H. Winsborough. Formal verification of security properties in trust management policy. Journal of Computer Security, 22(1):69–153, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Norman:2006:APC

[NS06] Gethin Norman and Vitaly Shmatikov. Analysis of probabilistic contract signing. Journal of Computer Security, 14(6):561–589, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Nguyen:2018:ASP

[NSC18] Binh Thanh Nguyen, Christoph Sprenger, and Cas Cremers. Abstractions for security protocol verification. Journal of Computer Security, 26(4):459–508, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Ngo:2014:EVC

[NSH14] Tri Minh Ngo, Mariëlle Stoelinga, and Marieke Huisman. Effective verification of confidentiality for multi-threaded programs. Journal of Computer Security, 22(2):269–300, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Ng:2011:PRO

[NSMSN11] Ching Yu Ng, Willy Susilo, Yi Mu, and Rei Safavi-Naini. Practical RFID ownership transfer scheme. Journal of Computer Security, 19(2):319–341, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Nithyanand:2011:UAR

[NTU11] Rishab Nithyanand, Gene Tsudik, and Ersin Uzun. User-aided reader revocation in PKI-based RFID systems. Journal of Computer Security, 19(6):1147–1172, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Nenadic:2005:RBV

[NZS05] Aleksandra Nenadić, Ning Zhang, and Qi Shi. RSA-based Verifiable and Recoverable Encryption of Signatures and its application in certified e-mail delivery. Journal of Computer Security, 13(5): 757–777, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Ourston:2004:CIA

[OMSH04] Dirk Ourston, Sara Matzner, William Stump, and Bryan Hopkins. Coordinated Internet attacks: responding to attack complexity. Journal of Computer Security, 12(2):165–190, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Paulson:1998:IAV

[Pau98] Lawrence C. Paulson. The inductive approach to verifying cryptographic protocols. Journal of Computer Security, 6(1–2):85–128, ???? 1998. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Paulson:2001:RBS

[Pau01] Lawrence C. Paulson. Relations between secrets: two formal analyses of the Yahalom protocol. Journal of Computer Security, 9(3): 197–216, ???? 2001. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Paiola:2013:VSP

[PB13] Miriam Paiola and Bruno Blanchet. Verification of security protocols with lists: From length one to unbounded length. Journal of Computer Security, 21(6):781–816, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Preda:2009:SBC

[PG09] Mila Dalla Preda and Roberto Giacobazzi. Semantics-based code obfuscation by abstract interpretation. Journal of Computer Security, 17(6):855–908, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Politz:2014:TBV

[PGK14] Joe Gibbs Politz, Arjun Guha, and Shriram Krishnamurthi. Typed-based verification of Web sandboxes. Journal of Computer Security, 22(4):511–565, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Petkovic:2009:P

[PJ09] Milan Petković and Willem Jonker. Preface. Journal of Computer Security, 17(1):1–3, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Preda:2018:CPD

[PM18] Mila Dalla Preda and Isabella Mastroeni. Characterizing a property-driven obfuscation strategy. Journal of Computer Security, 26(1):31–69, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Pereira:2003:SAU

[PQ03] Olivier Pereira and Jean-Jacques Quisquater. Some attacks upon authenticated group key agreement protocols. Journal of Computer Security, 11(4):555–580, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Pereira:2006:IBS

[PQ06] Olivier Pereira and Jean-Jacques Quisquater. On the impossibility of building secure Cliques-type authenticated group key agreement protocols. Journal of Computer Security, 14(2):197–246, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Price:2006:PKI

[Pri06] Geraint Price. Public key infrastructures: A research agenda. Journal of Computer Security, 14(5):391–417, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Pala:2010:FPN

[PS10a] Massimiliano Pala and Sean W. Smith. Finding the PKI needles in the Internet haystack. Journal of Computer Security, 18(3): 397–420, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Pucella:2010:IOS

[PS10b] Riccardo Pucella and Fred B. Schneider. Independence from obfuscation: A semantic framework for diversity. Journal of Computer Security, 18(5):701–749, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Peddinti:2014:WSQ

[PS14] Sai Teja Peddinti and Nitesh Saxena. Web search query privacy: Evaluating query obfuscation and anonymizing networks. Journal of Computer Security, 22(1):155–199, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Phatak:2013:SIN

[PSJ^{+}13] Dhananjay Phatak, Alan T. Sherman, Nikhil Joshi, Bhushan Sonawane, Vivek G. Relan, and Amol Dawalbhakta. Spread identity: A new dynamic address remapping mechanism for anonymity and DDoS defense. Journal of Computer Security, 21(2):233–281, ????

2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Pirretti:2010:SAB

[PTMW10] Matthew Pirretti, Patrick Traynor, Patrick McDaniel, and Brent Waters. Secure attribute-based systems. Journal of Computer Security, 18(5):799–837, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Pescape:2005:EAA

[PV05] Antonio Pescapè and Giorgio Ventre. Experimental analysis of attacks against intradomain routing protocols. Journal of Computer Security, 13(6):877–903, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Quisquater:1997:ASS

[QJ97] Jean-Jacques Quisquater and Marc Joye. Authentication of sequences with the SL_{2} hash function: application to video sequences. Journal of Computer Security, 5(3):213–223, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Ray:1998:SBT

[RAJ98] Indrakshi Ray, Paul Ammann, and Sushil Jajodia. A semanticbased transaction processing model for multilevel transactions. Journal of Computer Security, 6(3):181–217, ???? 1998. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Roscoe:1999:PSP

[RB99] A. W. Roscoe and P. J. Broadfoot. Proving security protocols with model checkers by data independence techniques. Journal of Computer Security, 7(2–3):147–190, ???? 1999. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Radosavac:2007:DIM

[RCBM07] S. Radosavac, Alvaro A. Cárdenas, John S. Baras, and George V. Moustakides. Detecting IEEE 802.11 MAC layer misbehavior in ad hoc networks: Robust strategies against individual and colluding attackers. Journal of Computer Security, 15(1):103–128, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Roy:2010:ITP

[RDDM10] Arnab Roy, Anupam Datta, Ante Derek, and John C. Mitchell. Inductive trace properties for computational security. Journal of Computer Security, 18(6):1035–1073, ???? 2010. CODEN JC-SIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Reiter:1996:KMS

[RFLW96] Michael K. Reiter, Matthew K. Franklin, John B. Lacy, and Rebecca N. Wright. The Ω key management service. Journal of Computer Security, 4(4):267–287, ???? 1996. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Raya:2007:SVA

[RH07] Maxim Raya and Jean-Pierre Hubaux. Securing vehicular ad hoc networks. Journal of Computer Security, 15(1):39–68, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Rubio:2019:TAI

[RRAZ19] Juan E. Rubio, Rodrigo Roman, Cristina Alcaraz, and Yan Zhang. Tracking APTs in industrial ecosystems: a proof of concept. Journal of Computer Security, 27(5):521–546, ???? 2019. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Ryan:2001:PAN

[RS01] P. Y. A. Ryan and S. A. Schneider. Process algebra and noninterference. Journal of Computer Security, 9(1–2):75–103, ???? 2001. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Ramakrishnan:2002:MBA

[RS02] C. R. Ramakrishnan and R. Sekar. Model-based analysis of configuration vulnerabilities. Journal of Computer Security, 10(1–2): 189–209, ???? 2002. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Ramanujam:2005:DCE

[RS05] R. Ramanujam and S. P. Suresh. Decidability of context-explicit security protocols. Journal of Computer Security, 13(1):135–165, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Rafnsson:2016:SME

[RS16] Willard Rafnsson and Andrei Sabelfeld. Secure multi-execution: Fine-grained, declassification-aware, and transparent. Journal of Computer Security, 24(1):39–90, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Ranise:2018:AEA

[RTV18] Silvio Ranise, Anh Truong, and Luca Viganò. Automated and efficient analysis of administrative temporal RBAC policies with role hierarchies. Journal of Computer Security, 26(4):423–458, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Rieck:2011:AAM

[RTWH11] Konrad Rieck, Philipp Trinius, Carsten Willems, and Thorsten Holz. Automatic analysis of malware behavior using machine learning. Journal of Computer Security, 19(4):639–668, ???? 2011. CO-DEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Rocchetto:2017:IBM

[RVV17] Marco Rocchetto, Luca Viganò, and Marco Volpe. An interpolation-based method for the verification of security protocols. Journal of Computer Security, 25(6):463–510, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Roscheisen:1997:NCD

[RW97] Martin Röscheisen and Terry Winograd. A network-centric design for relationship-based security and access control. Journal of Computer Security, 5(3):249–254, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Roscoe:1996:NIT

[RWW96] A. W. Roscoe, J. C. P. Woodcock, and L. Wulf. Non-interference through determinism. Journal of Computer Security, 4(1):27–53, ???? 1996. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Sonchack:2016:ELS

[SA16] John Sonchack and Adam J. Aviv. Exploring large scale security system reproducibility with the LESS simulator. Journal of Computer Security, 24(5):645–665, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Sabelfeld:2008:P

[Sab08] Andrei Sabelfeld. Preface. Journal of Computer Security, 16(5): 495, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Sabelfeld:2010:P

[Sab10] Andrei Sabelfeld. Preface. Journal of Computer Security, 18(6): 1075, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Stolfo:2005:CET

[SAE^{+}05] Salvatore J. Stolfo, Frank Apap, Eleazar Eskin, Katherine Heller, Shlomo Hershkop, Andrew Honig, and Krysta Svore. A comparative evaluation of two algorithms for Windows Registry Anomaly Detection. Journal of Computer Security, 13(4):659–693, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Sandhu:1992:GEP

[San92a] Ravi Sandhu. Guest Editor’s preface. Journal of Computer Security, 1(2):131–132, ???? 1992. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Sandhu:1992:EPS

[San92b] Ravi S. Sandhu. Expressive power of the schematic protection model. Journal of Computer Security, 1(1):59–98, ???? 1992. CO-DEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Shin:2011:EES

[SAsC11] Heechang Shin, Vijayalakshmi Atluri, and June suh Cho. Efficiently enforcing spatiotemporal access control under uncertain location information. Journal of Computer Security, 19(3):607–637, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Sandhu:1999:RBA

[SB99] Ravi Sandhu and Venkata Bhamidipati. Role-based administration of user-role assignment: The URA97 model and its Oracle implementation. Journal of Computer Security, 7(4):317–342, ???? 1999. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Sprenger:2018:RSP

[SB18a] Christoph Sprenger and David Basin. Refining security protocols. Journal of Computer Security, 26(1):71–120, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Stoller:2018:MHT

[SB18b] Scott D. Stoller and Thang Bui. Mining hierarchical temporal roles with multiple metrics. Journal of Computer Security, 26(1): 121–142, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Song:2001:ANA

[SBP01] Dawn Xiaodong Song, Sergey Berezin, and Adrian Perrig. Athena: a novel approach to efficient automatic security protocol analysis. Journal of Computer Security, 9(1–2):47–74, ???? 2001. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Sherwood:2005:PPS

[SBS05] Rob Sherwood, Bobby Bhattacharjee, and Aravind Srinivasan. P5: A protocol for scalable anonymous communication. Journal of Computer Security, 13(6):839–876, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Spalka:2000:SNS

[SC00] Adrian Spalka and Armin B. Cremers. Structured name-spaces in secure databases. Journal of Computer Security, 8(1):67–86, ???? 2000. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Stergiopoulos:2019:USC

[SCB^{+}19] George Stergiopoulos, Georgia Chronopoulou, Evangelos Bitsikas, Nikolaos Tsalis, and Dimitris Gritzalis. Using side channel TCP features for real-time detection of malware connections. Journal of Computer Security, 27(5):507–520, ???? 2019. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Schneider:2003:GEP

[Sch03] Steve Schneider. Guest Editor’s preface. Journal of Computer Security, 11(4):449–450, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Schneider:2004:GEP

[Sch04] Steve Schneider. Guest Editor’s preface. Journal of Computer Security, 12(3–4):313–315, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Shieh:1996:DIL

[SG96] Shiuh-Pyng Shieh and Virgil D. Gligor. Detecting illicit leakage of information in operating systems. Journal of Computer Security, 4(2–3):123–148, ???? 1996. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic). URL http://www.cl.cam.ac.uk/

~fapp2/steganography/bibliography/061242.html.

Spinellis:2002:PID

[SG02] Diomidis Spinellis and Dimitris Gritzalis. Panoptis: Intrusion detection using a domain-specific language. Journal of Computer Security, 10(1–2):159–176, ???? 2002. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Staniford:2002:PAD

[SHM02] Stuart Staniford, James A. Hoagland, and Joseph M. McAlerney. Practical automated detection of stealthy portscans. Journal of Computer Security, 10(1–2):105–136, ???? 2002. CODEN JC-SIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Shmatikov:2004:PAA

[Shm04] Vitaly Shmatikov. Probabilistic analysis of an anonymity system. Journal of Computer Security, 12(3–4):355–377, ???? 2004. CO-DEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Sinclair:1997:ASS

[Sin97] Jane Sinclair. Action systems for security specification. Journal of Computer Security, 5(2):129–154, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Sinha:2011:SSS

[Sin11] Anshuman Sinha. A survey of system security in contactless electronic passports. Journal of Computer Security, 19(1):203–226, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Soupionis:2014:GTA

[SKEG14] Yannis Soupionis, Remous-Aris Koutsiamanis, Pavlos Efraimidis, and Dimitris Gritzalis. A game-theoretic analysis of preventing spam over Internet Telephony via audio CAPTCHA-based authentication. Journal of Computer Security, 22(3):383–413, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Simmons:1995:RTI

[SM95] G. J. Simmons and Catherine Meadows. The role of trust in information integrity protocols. Journal of Computer Security, 3(1): 71–84, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Siaterlis:2005:OSA

[SM05] Christos Siaterlis and Vasilis Maglaris. One step ahead to multisensor data fusion for DDoS detection. Journal of Computer Security, 13(5):779–806, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Smith:2006:ITP

[Smi06] Geoffrey Smith. Improved typings for probabilistic noninterference in a multi-threaded language. Journal of Computer Security, 14(6): 591–623, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Sumii:2003:LRE

[SP03] Eijiro Sumii and Benjamin C. Pierce. Logical relations for encryption. Journal of Computer Security, 11(4):521–554, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Salvail:2010:STR

[SPD^{+}10] Louis Salvail, Momtchil Peev, Eleni Diamanti, Romain Alléaume, Norbert Lütkenhaus, and Thomas Länger. Security of trusted repeater quantum key distribution networks. Journal of Computer Security, 18(1):61–87, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Schlesinger:2014:MPA

[SPS^{+}14] Cole Schlesinger, Karthik Pattabiraman, Nikhil Swamy, David Walker, and Benjamin Zorn. Modular protections against noncontrol data attacks. Journal of Computer Security, 22(5):699–742, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Salva:2019:CAS

[SR19] Sébastien Salva and Loukmen Regainia. A catalogue associating security patterns and attack steps to design secure applications. Journal of Computer Security, 27(1):49–74, ???? 2019. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Syverson:1997:PWB

[SRG97] Paul F. Syverson, Michael G. Reed, and David M. Goldschlag. Private Web browsing. Journal of Computer Security, 5(3):237– 248, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Schellhorn:2002:VFS

[SRS^{+}02] Gerhard Schellhorn, Wolfgang Reif, Axel Schairer, Paul Karger, Vernon Austel, and David Toll. Verified formal security models for multiapplicative smart cards. Journal of Computer Security, 10(4):339–367, ???? 2002. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Samarati:1997:GEP

[SS97] Pierangela Samarati and Ravi Sandhu. Guest editors’ preface. Journal of Computer Security, 5(4):269–270, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Sabelfeld:2009:DDP

[SS09] Andrei Sabelfeld and David Sands. Declassification: Dimensions and principles. Journal of Computer Security, 17(5):517–548, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Saklikar:2010:IFV

[SS10] Samir Saklikar and Subir Saha. Identity federation for VoIP systems. Journal of Computer Security, 18(4):499–540, ???? 2010. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Stakhanova:2012:TCS

[SSBW12] Natalia Stakhanova, Chris Strasburg, Samik Basu, and Johnny S. Wong. Towards cost-sensitive assessment of intrusion response selection. Journal of Computer Security, 20(2–3):169–198, ????

2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Shirvanian:2018:SVI

[SSM18] Maliheh Shirvanian, Nitesh Saxena, and Dibya Mukhopadhyay. Short voice imitation man-in-the-middle attacks on Crypto Phones: Defeating humans and machines. Journal of Computer Security, 26(3):311–333, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Shroff:2008:SIF

[SST08] Paritosh Shroff, Scott F. Smith, and Mark Thober. Securing information flow via dynamic capture of dependencies. Journal of Computer Security, 16(5):637–688, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Shmatikov:2005:RBT

[ST05] Vitaly Shmatikov and Carolyn Talcott. Reputation-based trust management. Journal of Computer Security, 13(1):167–190, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Shen:2015:DIC

[ST15] Shiuan-Tzuo Shen and Wen-Guey Tzeng. Delegated integrity check for hierarchical cloud data. Journal of Computer Security, 23(4):471–508, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Sewell:2003:SCU

[SV03] Peter Sewell and Jan Vitek. Secure composition of untrusted code: box π, wrappers, and causality types. Journal of Computer Security, 11(2):135–187, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Spalazzi:2015:SIS

[SV15] Luca Spalazzi and Luca Viganò. Special issue on security and high performance computing systems. Journal of Computer Security, 23(5):539–540, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Shin:2011:PAM

[SVA11] Heechang Shin, Jaideep Vaidya, and Vijayalakshmi Atluri. A profile anonymization model for location-based services. Journal of Computer Security, 19(5):795–833, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Skormin:2007:PIA

[SVSM07] V. Skormin, A. Volynkin, D. Summerville, and J. Moronski. Prevention of information attacks by run-time detection of selfreplication in computer codes. Journal of Computer Security, 15(2):273–302, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Skalka:2007:RMD

[SWC07] Christian Skalka, X. Sean Wang, and Peter Chapin. Risk management for distributed authorization. Journal of Computer Security, 15(4):447–489, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Spindler:2008:PVS

[SWH^{+}08] Torsten Spindler, Christoph Wartmann, Ludger Hovestadt, Daniel Roth, Luc Van Gool, and Andreas Steffen. Privacy in video surveilled spaces. Journal of Computer Security, 16(2):199–222, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Syverson:1992:KBS

[Syv92] Paul F. Syverson. Knowledge, belief, and semantics in the analysis of cryptographic protocols. Journal of Computer Security, 1(3–4): 317–334, ???? 1992. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Syverson:2001:GEP

[Syv01] Paul F. Syverson. Guest Editor’s preface. Journal of Computer Security, 9(1–2):1–2, ???? 2001. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Syverson:2003:GEP

[Syv03] Paul F. Syverson. Guest Editor’s preface. Journal of Computer Security, 11(2):133, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Sun:2012:TAV

[SZP^{+}12] Yanjie Sun, Chenyi Zhang, Jun Pang, Baptiste Alcalde, and Sjouke Mauw. A trust-augmented voting scheme for collaborative privacy management. Journal of Computer Security, 20(4):437–459, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Tardo:1992:SGA

[TA92] Joseph J. Tardo and Kannan Alagappan. SPX: Global authentication using public key certificates. Journal of Computer Security, 1(3–4):295–316, ???? 1992. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Tiplea:2008:SBS

[ŢBEB08] Ferucio L. Ţiplea, Cătălin V. Bîrjoveanu, Constantin Enea, and Ioana Boureanu. Secrecy for bounded security protocols with freshness check is NEXPTIME-complete. Journal of Computer Security, 16(6):689–712, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Taglienti:2016:UAP

[TC16] Claudio Taglienti and James Cannady. The user attribution problem and the challenge of persistent surveillance of user activity in complex networks. Journal of Computer Security, 24(2):235–288, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Traynor:2008:EOF

[TEML08] Patrick Traynor, William Enck, Patrick McDaniel, and Thomas La Porta. Exploiting open functionality in SMS-capable cellular networks. Journal of Computer Security, 16(6):713–742, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Tripunitara:2007:TCE

[TL07] Mahesh V. Tripunitara and Ninghui Li. A theory for comparing the expressive power of access control models. Journal of Computer Security, 15(2):231–272, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Tian:2019:DSD

[TLY^{+}19] Yangguang Tian, Yingjiu Li, Guomin Yang, Willy Susilo, Yi Mu, Hui Cui, and Yinghui Zhang. DABKE: Secure deniable attributebased key exchange framework. Journal of Computer Security, 27(2):259–275, ???? 2019. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Toahchoodee:2011:FAS

[TR11] Manachai Toahchoodee and Indrakshi Ray. On the formalization and analysis of a spatio-temporal role-based access control model. Journal of Computer Security, 19(3):399–452, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Trostle:1993:MFT

[Tro93] Jonathan T. Trostle. Modelling a fuzzy time system. Journal of Computer Security, 2(4):291–309, ???? 1993. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Thomas:1993:KAM

[TS93] Roshan K. Thomas and Ravi S. Sandhu. A kernelized architecture for multilevel secure object-oriented databases supporting writeup. Journal of Computer Security, 2(2–3):231–275, ???? 1993. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Thuraisingham:2003:GEP

[TvdR03] Bhavani Thuraisingham and Reind van de Riet. Guest Editors’ preface. Journal of Computer Security, 11(3):289, ???? 2003. CO-DEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Teepe:2003:WAS

[TvdRO03] Wouter Teepe, Reind van de Riet, and Martin Olivier. WorkFlow analyzed for security and privacy in using databases. Journal of Computer Security, 11(3):353–363, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Uzun:2014:SAT

[UAV^{+}14] Emre Uzun, Vijayalakshmi Atluri, Jaideep Vaidya, Shamik Sural, Anna Lisa Ferrara, Gennaro Parlato, and P. Madhusudan. Security analysis for temporal role based access control. Journal of Computer Security, 22(6):961–996, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Uribe:2007:AAF

[UC07] Tomás E. Uribe and Steven Cheung. Automatic analysis of firewall and network intrusion detection system configurations. Journal of Computer Security, 15(6):691–715, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Urban:2019:ALP

[UTHP19] Tobias Urban, Dennis Tatang, Thorsten Holz, and Norbert Pohlmann. Analyzing leakage of personal information by malware. Journal of Computer Security, 27(4):459–481, ???? 2019. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Vaidya:2009:ERM

[VAGL09] Jaideep Vaidya, Vijayalakshmi Atluri, Qi Guo, and Haibing Lu. Edge-RMP: Minimizing administrative assignments for role-based access control. Journal of Computer Security, 17(2):211–235, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Vaidya:2005:SSI

[VC05] Jaideep Vaidya and Chris Clifton. Secure set intersection cardinality with application to association rule mining. Journal of Computer Security, 13(4):593–622, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

vanderMeyden:2015:WII

[vdM15] Ron van der Meyden. What, indeed, is intransitive noninterference? Journal of Computer Security, 23(2):197–228, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Vaidya:2011:I

[VG11] Jaideep Vaidya and Ehud Gudes. Introduction. Journal of Computer Security, 19(3):485–486, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Volpano:1996:STS

[VIS96] Dennis Volpano, Cynthia Irvine, and Geoffrey Smith. A sound type system for secure flow analysis. Journal of Computer Security, 4 (2–3):167–187, ???? 1996. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Visconti:2013:SIA

[Vis13] Ivan Visconti. Special issue: Advances in security for communication networks. Journal of Computer Security, 21(5):599–600, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Vigna:1999:NNB

[VK99] Giovanni Vigna and Richard A. Kemmerer. NetSTAT: A networkbased intrusion detection system. Journal of Computer Security, 7 (1):37–71, ???? 1999. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

vanOorschot:2011:EPC

[vOT11] P. C. van Oorschot and Julie Thorpe. Exploiting predictability in click-based graphical passwords. Journal of Computer Security, 19(4):669–702, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Vavilis:2016:SBQ

[VPZ16] Sokratis Vavilis, Milan Petković, and Nicola Zannone. A severitybased quantification of data leakages in database systems. Journal of Computer Security, 24(3):321–345, ???? 2016. CODEN JC-SIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Volpano:1999:PNC

[VS99] Dennis Volpano and Geoffrey Smith. Probabilistic noninterference in a concurrent language. Journal of Computer Security, 7(2–3): 231–253, ???? 1999. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Vigna:2009:REA

[VVB^{+}09] Giovanni Vigna, Fredrik Valeur, Davide Balzarotti, William Robertson, Christopher Kruegel, and Engin Kirda. Reducing errors in the anomaly-based detection of web-based attacks through the combined analysis of web requests and SQL queries. Journal of Computer Security, 17(3):305–329, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Wang:2006:GNR

[Wan06] Guilin Wang. Generic non-repudiation protocols supporting transparent off-line TTP. Journal of Computer Security, 14(5):441–467, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Warinschi:2005:CAN

[War05] Bogdan Warinschi. A computational analysis of the Needham– Schroeder–(Lowe) protocol. Journal of Computer Security, 13(3):

565–591, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Winslett:1997:UDC

[WCJS97] M. Winslett, N. Ching, V. Jones, and I. Slepchin. Using digital credentials on the World Wide Web. Journal of Computer Security, 5(3):255–267, ???? 1997. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Wang:2008:CBH

[WD08] Lifu Wang and Partha Dasgupta. Coprocessor-based hierarchical trust management for software integrity and digital identity protection. Journal of Computer Security, 16(3):311–339, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Wespi:2000:FVV

[WDDN00] Andreas Wespi, Hervé Debar, Marc Dacier, and Mehdi Nassehi. Fixed- vs. variable-length patterns for detecting suspicious process behavior. Journal of Computer Security, 8(2–3):159–181, ???? 2000. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Wehner:2007:AWN

[Weh07] Stephanie Wehner. Analyzing worms and network traffic using compression. Journal of Computer Security, 15(3):303–320, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Woo:1993:ADS

[WL93] Thomas Y. C. Woo and Simon S. Lam. Authorization in distributed systems: A new approach. Journal of Computer Security, 2(2–3):107–136, ???? 1993. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Wang:2007:PBI

[WLJW07] Lingyu Wang, Yingjiu Li, Sushil Jajodia, and Duminda Wijesekera. Parity-based inference control for multi-dimensional range sum queries. Journal of Computer Security, 15(4):417–445, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Wright:2001:DGM

[WLM01] Rebecca N. Wright, Patrick D. Lincoln, and Jonathan K. Millen. Depender graphs: A method of fault-tolerant certificate distribution. Journal of Computer Security, 9(4):323–338, ???? 2001. CO-DEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Wright:2002:DGM

[WLM02] Rebecca N. Wright, Patrick D. Lincoln, and Jonathan K. Millen. Depender graphs: A method of fault-tolerant certificate distribution. Journal of Computer Security, 10(3):297, ???? 2002. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Wuller:2017:DPP

[WMF^{+}17] Stefan Wüller, Daniel Mayer, Fabian Förg, Samuel Schüppen, Benjamin Assadsolimani, Ulrike Meyer, and Susanne Wetzel. Designing privacy-preserving interval operations based on homomorphic encryption and secret sharing techniques. Journal of Computer Security, 25(1):59–81, ???? 2017. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Wray:1992:ACT

[Wra92] John C. Wray. An analysis of covert timing channels. Journal of Computer Security, 1(3–4):219–232, ???? 1992. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Welch:2002:URM

[WS02] Ian Welch and Robert J. Stroud. Using reflection as a mechanism for enforcing security policies on compiled code. Journal of Computer Security, 10(4):399–432, ???? 2002. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Wang:2004:CBI

[WWJ04] Lingyu Wang, Duminda Wijesekera, and Sushil Jajodia. Cardinality-based inference control in data cubes. Journal of Computer Security, 12(5):655–692, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Wang:2008:IIA

[WYSJ08] Lingyu Wang, Chao Yao, Anoop Singhal, and Sushil Jajodia. Implementing interactive analysis of attack graphs using relational databases. Journal of Computer Security, 16(4):419–437, ????

2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Xu:2009:ASL

[XC09] Haizhi Xu and Steve J. Chapin. Address-space layout randomization using code islands. Journal of Computer Security, 17(3): 331–362, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Xu:2007:SGC

[Xu07] Shouhuai Xu. On the security of group communication schemes. Journal of Computer Security, 15(1):129–169, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Xia:2006:EWD

[XVW^{+}06] Jianhong Xia, Sarma Vangala, Jiang Wu, Lixin Gao, and Kevin Kwiat. Effective worm detection for various scan techniques. Journal of Computer Security, 14(4):359–387, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Yahalom:1993:OAT

[Yah93] Raphael Yahalom. Optimality of asynchronous two-party secure data-exchange protocols. Journal of Computer Security, 2(2–3): 191–209, ???? 1993. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Yasinsac:2002:ESP

[Yas02] Alec Yasinsac. An environment for security protocol intrusion detection. Journal of Computer Security, 10(1–2):177–188, ???? 2002. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Yee:2008:GEIa

[YGH08] George O. M. Yee, Ali A. Ghorbani, and Patrick C. K. Hung. Guest Editors’ introduction. Journal of Computer Security, 16(2): 103–106, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Yang:2015:PAA

[YGSY15] Ping Yang, Mikhail I. Gofman, Scott D. Stoller, and Zijiang Yang. Policy analysis for administrative role based access control without separate administration. Journal of Computer Security, 23(1):1– 29, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Yu:2005:SUI

[YLZ05] Meng Yu, Peng Liu, and Wanyu Zang. Specifying and using intrusion masking models to process distributed operations. Journal of Computer Security, 13(4):623–658, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Yee:2008:GEIb

[YRY08] George O. M. Yee, Chunming Rong, and Laurence T. Yang. Guest Editors’ introduction. Journal of Computer Security, 16(3):261– 264, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Yarmand:2013:BBA

[YSD13] Mohammad H. Yarmand, Kamran Sartipi, and Douglas G. Down. Behavior-based access control for distributed healthcare systems. Journal of Computer Security, 21(1):1–39, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Yuen:2014:TCT

[YSM14] Tsz Hon Yuen, Willy Susilo, and Yi Mu. Towards a cryptographic treatment of publish/subscribe systems. Journal of Computer Security, 22(1):33–67, ???? 2014. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Yasuoka:2011:BPQ

[YT11] Hirotoshi Yasuoka and Tachio Terauchi. On bounding problems of quantitative information flow. Journal of Computer Security, 19(6):1029–1082, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Yu:2019:CAF

[YTLD19] Xingjie Yu, Michael Shiwen Thang, Yingjiu Li, and Robert Huijie Deng. Collusion attacks and fair time-locked deposits for fastpayment transactions in Bitcoin. Journal of Computer Security, 27(3):375–403, ???? 2019. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Yao:2009:EPT

[YWW^{+}09] Chao Yao, Lingyu Wang, X. Sean Wang, Claudio Bettini, and Sushil Jajodia. Evaluating privacy threats in released database views by symmetric indistinguishability. Journal of Computer Security, 17(1):5–42, ???? 2009. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Yuan:2015:PPC

[YY15] Jiawei Yuan and Shucheng Yu. PCPOR: Public and constant-cost proofs of retrievability in cloud1. Journal of Computer Security, 23(3):403–425, ???? 2015. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Zhou:2008:SPR

[ZAF08] Jie Zhou and Jim Alves-Foss. Security policy refinement and enforcement for the design of multi-level secure systems. Journal of Computer Security, 16(2):107–131, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Zanin:2007:RRD

[ZDM07] Giorgio Zanin, Roberto Di Pietro, and Luigi V. Mancini. Robust RSA distributed signatures for large-scale long-lived ad hoc networks. Journal of Computer Security, 15(1):171–196, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Zhou:2012:FAS

[ZF12] Hongbin Zhou and Simon N. Foley. Fast automatic security protocol generation. Journal of Computer Security, 20(2–3):119–167, ???? 2012. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Zhang:2004:RAR

[ZGD04] Jian Zhang, Jian Gong, and Yong Ding. Research on automated rollbackability of intrusion response. Journal of Computer Security, 12(5):737–751, ???? 2004. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Zhou:2013:SVC

[ZGDS13] Fangfei Zhou, Manish Goel, Peter Desnoyers, and Ravi Sundaram. Scheduler vulnerabilities and coordinated attacks in cloud computing. Journal of Computer Security, 21(4):533–559, ???? 2013. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Zhang:2018:PPV

[ZHZY18] Jun Zhang, Meiqi He, Gongxian Zeng, and Siu-Ming Yiu. Privacypreserving verifiable elastic net among multiple institutions in the cloud. Journal of Computer Security, 26(6):791–815, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Zhou:2006:P

[ZK06] Jianying Zhou and Meng-Chow Kang. Preface. Journal of Computer Security, 14(5):389–390, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Zakinthinos:1995:CNI

[ZL95] A. Zakinthinos and E. S. Lee. The composability of noninterference. Journal of Computer Security, 3(4):269–281, ???? 1995. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Zouridaki:2007:HQT

[ZMHT07] Charikleia Zouridaki, Brian L. Mark, Marek Hejmo, and Roshan K. Thomas. Hermes: A quantitative trust establishment framework for reliable data packet delivery in MANETs. Journal of Computer Security, 15(1):3–38, ???? 2007. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Zhang:2008:SVA

[ZRG08] Nan Zhang, Mark Ryan, and Dimitar P. Guelev. Synthesising verified access control systems through model checking. Journal of Computer Security, 16(1):1–61, ???? 2008. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Zhu:2006:GEG

[ZSXJ06] Sencun Zhu, Sanjeev Setia, Shouhuai Xu, and Sushil Jajodia. GKMPAN: An efficient group rekeying scheme for secure multicast in ad-hoc networks. Journal of Computer Security, 14(4): 301–325, ???? 2006. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Zuquete:2005:EHQ

[Zúq05] André Zúquete. An efficient high quality random number generator for multi-programmed systems. Journal of Computer Security, 13(2):243–263, ???? 2005. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic). URL http://dl.acm.org/

citation.cfm?id=1077819.1077821.

Zavatteri:2019:LMS

[ZV19] Matteo Zavatteri and Luca Viganò. Last man standing: Static, decremental and dynamic resiliency via controller synthesis. Journal of Computer Security, 27(3):343–373, ???? 2019. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Zhang:2011:UTA

[ZZW^{+}11] Chao Zhang, Wei Zou, Tielei Wang, Yu Chen, and Tao Wei. Using type analysis in compiler to mitigate integer-overflow-to-buffer-overflow threat. Journal of Computer Security, 19(6):1083–1107, ???? 2011. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).